### v1

# Multi-Objective min-max Online Convex Optimization

Abstract
In online convex optimization (OCO), a single loss function sequence is revealed over a time horizon of $T$, and an online algorithm has to choose its action at time $t$, before the loss function at time $t$ is revealed. The goal of the online algorithm is to incur minimal penalty (called regret) compared to a static optimal action made by an optimal offline algorithm knowing all functions of the sequence in advance.
In this paper, we broaden the horizon of OCO, and consider multi-objective OCO, where there are $K$ distinct loss function sequences, and an algorithm has to choose its action at time $t$, before the $K$ loss functions at time $t$ are revealed. To capture the tradeoff between tracking the $K$ different sequences, we consider the min-max regret, where the benchmark (optimal offline algorithm) takes a static action across all time slots that minimizes the maximum of the total loss (summed across time slots) incurred by each of the $K$ sequences. An online algorithm is allowed to change its action across time slots, and its min-max regret is defined as the difference between its min-max cost and that of the benchmark. The min-max regret is a stringent performance measure and an algorithm with small regret needs to ‘track’ all loss function sequences closely at all times.
We consider this min-max regret in the i.i.d. input setting where all loss functions are i.i.d. generated from an unknown distribution. For the i.i.d. model we propose a simple algorithm that combines the well-known Hedge and online gradient descent (OGD) and show via a remarkably simple proof that its expected min-max regret is $O(\sqrt{T\log K})$.
1 Introduction
Online convex optimization (OCO) Hazan (2019) is a basic learning problem that models a wide variety of applications such as spam filtering Hazan (2019), capacity provisioning in cloud networks Lin et al. (2012) etc., where at time $t$, an online algorithm $\mathcal{A}$ has to choose an action $x_{t}\in\mathcal{X}$, after which the adversary reveals the convex objective/loss function $f_{t}$ on which $x_{t}$ is evaluated. As a benchmark, the offline optimal algorithm is considered, that knows the entire set of $f_{t}$, $t=1,\dots,T$, however, can only choose a single action across all times. To quantify the performance of online algorithms, the metric of static regret $R^{\text{s}ingle}_{T}$ is studied, that is defined as follows, where $x_{t}$’s are the actions taken by $\mathcal{A}$,
$R^{\text{s}ingle}_{\mathcal{A}}(T)=\sup_{f_{t}}\left\{\sum_{t=1}^{T}f_{t}(x_{t})-\min_{x}\sum_{t=1}^{T}f_{t}(x)\right\}.$ (1)
In a long line of work, comprehensively summarized in Hazan (2019), tight guarantees (both upper and lower bounds) have been obtained for the OCO problem, and $R^{\text{s}ingle}(T)=\inf_{\mathcal{A}}R^{\text{s}ingle}_{\mathcal{A}}(T)=\Theta(\sqrt{T})$. In particular, online gradient descent (OGD), mirror descent, and follow-the-regularized-leader (FTRL) Hazan (2019) all achieve the optimal static regret performance for OCO.
Many modern real-world decision problems, however, involve multiple conflicting performance metrics, motivating the multi-objective OCO framework—an extension of OCO that aims to achieve Pareto-efficient trade-offs among competing objectives. However, Pareto optimality is too challenging to work with in the online setting, both in terms of defining the right benchmark as well as the required analytical treatment. One such attempt was made at this recently in Jiang et al. (2023), however, it used a weak regret metric, which remained independent of the number of objectives. We provide more details of this in the Related Work section.
Thus, in this paper, we consider the robust alternative in the multi-objective case: the $\it min-max$} OCO, where the objective of an algorithm is to minimize the maximum of the total loss (summed across time slots) incurred by each of the multiple objective functions. A compelling example for considering the min-max approach is fair machine learning, where performance is typically evaluated across multiple demographic groups, which naturally induces a multi-objective optimization problem. While Pareto efficiency identifies trade-offs between groups, it may still permit highly unequal solutions, as a classifier can be Pareto optimal even if it performs very poorly on a disadvantaged group. To address this limitation, fairness is often modeled via a min–max objective: maximize the minimum accuracy (or minimize the maximum error) across groups Chen et al. (2020); Chouldechova and Roth (2020)).
The min-max approach is also well suited for modern networked systems—such as data centers, wireless edge networks, and shared cloud infrastructures—where online decisions must balance performance across multiple service classes, users, or flows whose objectives (throughput, latency, energy) often conflict. The min–max regret criterion provides a principled way to ensure worst-case fairness by minimizing the largest cumulative loss among competing entities, which is crucial in fair resource allocation Sinha et al. (2023), fair allocation in clouds Hao et al. (2016), fair assignment of reusable resources Liu and Hajiesmaili (2025) congestion-controlled transport Kouvelis and Yu (1997); Bertsimas et al. (2011), and dynamic resource provisioning in data centers Urgaonkar et al. (2010). Unlike average-regret formulations that can conceal systematic bias, minimizing min–max regret explicitly guards against persistent under-allocation to any group or flow—an increasingly important property in multi-tenant and QoS-sensitive networks.
The formal min-max OCO problem that we consider is as follows. We consider that there are two convex loss function sequences $f_{t},g_{t},$ that are revealed after $\mathcal{A}$ chooses action $x_{t}$ at time $t$ for $t=1,\dots,T$. The min-max cost of $\mathcal{A}$ is
$C_{\mathcal{A}}=\max\left\{\sum_{t=1}^{T}f_{t}(x_{t}),\sum_{t=1}^{T}g_{t}(x_{t})\right\}.$ (2)
Two loss function sequences are considered for simplicity of exposition since they capture the basic difficulty. Extension to $K>2$ sequences is straightforward and detailed in Remark 7. As a benchmark, we consider the static offline optimal algorithm $\mathsf{OPT}$ that knows $f_{t},g_{t}$ for $t=1,\dots,T$ at time $t=1$ itself but can only choose a single action $x^{\star}$ across all times such that
$x^{\star}=\arg\min_{x\in\mathcal{X}}\max\left\{\sum_{t=1}^{T}f_{t}(x),\sum_{t=1}^{T}g_{t}(x)\right\},$ (3)
and the optimal cost is
$C_{\mathsf{OPT}}=\max\left\{\sum_{t=1}^{T}f_{t}(x^{\star}),\sum_{t=1}^{T}g_{t}(x^{\star})\right\}.$ (4)
Consequently, we define the min-max static regret of $\mathcal{A}$ with actions $x_{t}$ as
$R_{\mathcal{A}}(T)=\sup_{f_{t},g_{t}}\{C_{\mathcal{A}}-C_{\mathsf{OPT}}\}.$ (5)
We refer to (5) as the min-max OCO problem.
Remark 1 (On the regret definition).
The chosen benchmark in (5) is deliberately stringent. If we considered a weaker benchmark, e.g., minimizing $\min_{x\in X}\sum_{t=1}^{T}\max\{f_{t}(x),g_{t}(x)\}$), then applying standard OGD algorithm on the surrogate function $h_{t}(x)=\max\{f_{t}(x),g_{t}(x)\}$ would yield regret $O(\sqrt{T})$ independent of the number of objectives, underestimating the intrinsic hardness of balancing multiple sequences. In particular:
  • Our definition forces the algorithm to track both sequences globally and simultaneously. This means the online algorithm cannot focus on one objective at the expense of the other, nor can it switch actions adaptively across time as the per-slot benchmark allows.
  • Our benchmark ensures robustness across objectives: the comparator $x^{\star}$ is the best fixed action that minimizes the worst cumulative cost.
  • This definition aligns with fairness and robustness considerations (e.g., in machine learning across demographic groups), where good performance must be guaranteed for all objectives, not merely on average or in expectation.
Thus, while the regret metric in (5) is harder to minimize, it is an appropriate notion when robustness across multiple objectives is essential.
1.1 Related Work
Because of its widespread applications, multi-objective OCO has been an object of immense interest, however, with limited success. For example, Mahdavi et al. (2013) wanted to study the multi-objective OCO and stated the difficulty as “A fundamental difference between single- and multi-objective optimization is that for the latter it is not obvious how to evaluate the optimization quality”. Further, it said “Since it is impossible to simultaneously minimize multiple loss functions and in order to avoid complications caused by handling more than one objective, we choose one function as the objective and try to bound other objectives by appropriate thresholds. This work gave rise to the now popular problem of constrained online convex optimization (COCO), where the problem is transformed into a constrained optimization problem with one objective and one or more constraint functions, that is described next.
1.1.1 COCO
In COCO, on every round $t,$ an online algorithm first chooses an admissible action $x_{t}\in\mathcal{X}\subset{\mathbb{R}}^{d},$ and then the adversary chooses a convex loss/cost function $f_{t}:\mathcal{X}\to\mathbb{R}$ and a constraint function of the form $g_{t}(x)\leq 0,$ where $g_{t}:\mathcal{X}\to\mathbb{R}$ is a convex function. Let $\mathcal{X}^{\star}$ be the feasible set consisting of all admissible actions that satisfy all constraints $g_{t}(x)\leq 0,t\in[T]$. A standard assumption is made that $\mathcal{X}^{\star}$ is not empty (called the feasibility assumption). Since $g_{t}$’s are revealed after the action $x_{t}$ is chosen, an online algorithm $\mathcal{A}$ need not take feasible actions on each round, and in addition to the static regret defined as
$\textrm{R}^{\text{COCO}}_{\mathcal{A}}(T)\equiv\sup_{\{f_{t}\}_{t=1}^{T}}\left\{\sum_{t=1}^{T}f_{t}(x_{t})-\inf_{x\in\mathcal{X}^{\star}}\sum_{t=1}^{T}f_{t}(x^{\star})\right\},$ (6)
an additional metric of interest is the total cumulative constraint violation (CCV) defined as
$\textrm{CCV}_{\mathcal{A}}(T)\equiv\sum_{t=1}^{T}\max(g_{t}(x_{t}),0).$
The goal with COCO is to design an online algorithm to simultaneously achieve a small regret (6) with respect to any admissible benchmark $x^{\star}\in\mathcal{X}^{\star}$ and a small CCV.
1.1.2 Prior Work on COCO
(A) i.i.d. loss and constraints: COCO with independent and identically distributed (i.i.d.) loss functions $f_{t}$’s and constraint functions $g_{t}$’s was studied in Mahdavi et al. (2013) while its extension to stationary and ergodic processes can be found in Uziel and El-Yaniv (2017), where both derived an algorithm with regret $O(\sqrt{T})$ and CCV $O(T^{3/4})$.
(B) Time-invariant constraints: COCO with time-invariant constraints, i.e., $g_{t}=g,\forall\ t$ (Yuan and Lamperski, 2018; Jenatton et al., 2016; Mahdavi et al., 2012; Yi et al., 2021) has been considered extensively, where functions $g$ are assumed to be known to the algorithm a priori. The best known results in this case give static regret of $O(T^{\max\{c,1-c\}})$ and CCV of $O(T^{(1-c)/2})$ for any $c\in(0,1)$. The algorithm is allowed to take actions that are infeasible at any time to avoid the costly projection step of the vanilla projected OGD algorithm and the main objective was to design an efficient algorithm with a small regret and CCV while avoiding the explicit projection step.
(C) Time-varying constraints: The more difficult question is solving COCO problem when the constraint functions, i.e., $g_{t}$’s, change arbitrarily with time $t$. One popular algorithm for solving COCO considered a Lagrangian function optimization that is updated using the primal and dual variables (Yu et al., 2017; Sun et al., 2017; Yi et al., 2023). Alternatively, (Neely and Yu, 2017) and (Liakopoulos et al., 2019) used the drift-plus-penalty (DPP) framework (Neely, 2010) to solve the COCO, but which needed additional assumption, e.g. the Slater’s condition in (Neely and Yu, 2017) and with weaker form of the feasibility assumption in (Neely and Yu, 2017).
(Guo et al., 2022) obtained bounds similar to (Neely and Yu, 2017) but without assuming Slater’s condition. Very recently, the state of the art guarantees on simultaneous bounds on regret $O(\sqrt{T})$ and CCV $O(\sqrt{T}\log T)$ for COCO were derived in (Sinha and Vaze, 2024) with a very simple algorithm that combines the loss function at time $t$ and the CCV accrued till time $t$ in a single loss function, and then executes the online gradient descent (OGD) algorithm on the single loss function with an adaptive step-size. Most recently (Sinha and Vaze, 2025) showed lower CCV is possible with increased regret compared to (Sinha and Vaze, 2024) while (Vaze and Sinha, 2025) obtained instance dependent CCV bounds which can be $O(1)$ for specific cases while achieving regret of $O(\sqrt{T})$, respectively. The COCO problem has been considered in the dynamic setting as well (Chen and Giannakis, 2018; Cao and Liu, 2018; Vaze, 2022; Liu et al., 2022) where the benchmark $x^{\star}$ in (6) is replaced by $x_{t}^{\star}$ ($x_{t}^{\star}=\arg\min_{x}f_{t}(x)$) that is also allowed to change its actions over time.
Remark 2.
Even though the min-max OCO and the COCO problem appear related, but unfortunately results for one do not give results for the other in either direction.
1.2 Global Cost Functions in Experts Setting
A special case of the min-max OCO problem (5) was considered in Even-Dar et al. (2009), where there are $K$ experts, with expert $i$’s loss at time $t$, being $\ell_{t}(i)$. If expert $i$ is selected by an algorithm $\mathcal{A}$ with probability $p_{t}(i)$ at time $t$, then the overall expected loss for expert $i$ is $L_{i}=\sum_{t=1}^{T}p_{t}(i)\ell_{t}(i)$ where $T$ is the time horizon, and the cost of $\mathcal{A}$ is the global cost Even-Dar et al. (2009)
$C_{\mathcal{A}}=\max_{i=1,\dots,K}\{L_{i}\}.$ (7)
The regret of $\mathcal{A}$ is defined as usual with respect to the static optimal offline algorithm that is required to choose a fixed distribution across all time slots while knowing all losses in advance.
The name global refers to the fact that the objective function (7) captures the end-to-end cost function rather than a summation of the instantaneous losses. For this formulation, which we refer to as GlobalExpertsProblem, Even-Dar et al. (2009) proposed an algorithm called MULTI and showed its regret to be $O(\sqrt{T}\log(K))$. GlobalExpertsProblem differs with our min-max OCO formulation, in two fundamental ways, i) the objective inside the $\max$ is a linear function of the actions $p_{t}(i)$, and ii) the action $p_{t}(i)$ taken at time $t$ affects each expert’s loss but only in an average sense since $\sum_{i=1}^{K}p_{t}(i)=1$ while with min-max OCO the action $x_{t}$ chosen at time $t$ affects both $\sum_{t=1}^{T}f_{t}(x_{t})$ and $\sum_{t=1}^{T}g_{t}(x_{t})$ in full. Even though this appears to only be a small difference, it changes everything. As a consequence the result on GlobalExpertsProblem does not apply for min-max OCO even when $f_{t},g_{t}$ are linear. Extensions of Even-Dar et al. (2009) with additive global functions can be found in Azar et al. (2014); Kesselheim and Singla (2020) while for non-additive ones in Rakhlin et al. (2011).
1.3 Pareto Optimal Multi-Objective OCO
More recently, Jiang et al. (2023) attempted to study the multi-objective OCO from the Pareto optimality point of view. Towards this end, it used the concept of Pareto-suboptimality-gap (PSG), for $k=1,\dots,K$ loss functions $f_{t}^{k}$, and defined the regret to be [^1] We write the equivalent simpler form given on Page 21 Jiang et al. (2023) compared to the original Definition 4 made on Page 5 of Jiang et al. (2023) that is mathematically more complicated.
$\sup_{x^{\star}\in\mathcal{P}^{\star}}\min_{k=1,\dots,K}\sum_{t=1}^{T}(f_{t}^{k}(x_{t})-f_{t}^{k}(x^{\star})).$ (8)
where $\mathcal{P}^{\star}$ is the set of points on the Pareto frontier for the $K$-dimensional vector $\sum_{t=1}^{T}f_{t}^{k}]_{k=1}^{K}$. Jiang et al. (2023) proposed a mirror-descent inspired algorithm and showed that its regret (8) is $O(\sqrt{T})$ which notably is independent of $K$. The minimization over the $K$ possible loss functions in (8) made the regret (8) not depend on $K$, which highlights the weakness of the metric. Thus, even though Jiang et al. (2023) made a valid attempt at modelling the general multi-objective OCO, the weakness of the metric limited its applicability.
1.4 Per-slot $\it{\it min-max}$} multi-objective optimization
Compared to our ‘global’ min-max objective (4), in Lee et al. (2022), a per-slot $\it min-max$} multi-objective OCO was studied, where in particular, at each round $t=1,\dots,T$: i) an online algorithm chooses $x_{t}\in\mathcal{X}$ (convex compact set), ii) the adversary chooses $y_{t}\in\mathcal{Y}$ (convex compact set), and iii) the online algorithm suffers a vector loss:
$\ell_{t}(x_{t},y_{t})=\big(\ell_{t,1}(x_{t},y_{t}),\dots,\ell_{t,d}(x_{t},y_{t})\big)\in[-1,1]^{d}.$
The goal of an online algorithm is to solve $\min_{\{x_{t}\}}\max_{j\in[d]}\sum_{t=1}^{T}\ell_{t,j}(x_{t},y_{t})$, knowing $\ell_{t}(.,.)$, but not $y_{t}$, and its performance is compared against a per-round local minimax value
$w_{L}^{t}:=\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\max_{j\in[d]}\ell_{t,j}(x,y),$
where the cumulative benchmark is:
$W_{L}^{T}:=\sum_{t=1}^{T}w_{L}^{t},\quad\text{and regret}\quad\mathrm{R}^{d}_{\mathcal{A}}(T):=\max_{j\in[d]}\sum_{t=1}^{T}\ell_{t,j}(x_{t},y_{t})-W_{L}^{T}.$ (9)
A Hedge-type algorithm Cesa-Bianchi and Lugosi (2006) was shown in Lee et al. (2022) to have $O(\sqrt{T})$ regret.
Defining $\mathcal{Y}=\{e\}$ a singleton, and $d=2$, such that $\ell_{t,1}(x_{t},y_{t})=f_{t}(x_{t})$ and $\ell_{t,1}(x_{t},y_{t})=g_{t}(x_{t})$ for $y_{t}=e$, we recover our model where two convex functions $f_{t},g_{t}$ are revealed at time $t$, and for which the benchmark $W_{L}^{T}$ simplifies to
$W_{L}^{T}=\sum_{t=1}^{T}\min_{x_{t}\in\mathcal{X}}\max\{f_{t}(x_{t}),g_{t}(x_{t})\}).$ (10)
Thus, benchmark (10) is fundamentally different than our considered benchmark (4) since it is additive across time slots, and allows the benchmark to choose different actions across time slots. Even though (10) allows choosing different actions across time slots, surprisingly, we show in Appendix D that $W_{L}^{T}$ (10) can be $3/2$ times more than our benchmark $C_{\mathsf{OPT}}$ (4) and hence algorithms with $O(\sqrt{T})$ regret (9) do not yield a $O(\sqrt{T})$ min-max static regret (5).
In summary, compared to COCO and GlobalExpertsProblem, our considered problem has a wider scope, while our considered metric (5) is more meaningful and stringent compared to (8) and (10).
1.5 Applications
Multi-objective OCO has also been studied from an applications perspective, e.g. Lyu et al. (2019) studied the problem of multi-period multi-objective online ride-matching problem, where the multiple objectives included platform revenue, assigning drivers with higher service quality, minimizing the distance to the customer etc. Lyu et al. (2019) also used the COCO-type reformulation of multi-objective OCO for deriving theoretical guarantees. Other applications of multi-objective OCO include fair resource allocation Sinha et al. (2023), fair allocation in clouds Hao et al. (2016), fair assignment of reusable resources Liu and Hajiesmaili (2025) congestion-controlled transport Kouvelis and Yu (1997); Bertsimas et al. (2011), and dynamic resource provisioning in data centers Urgaonkar et al. (2010), multi-task learning Sener and Koltun (2018), multiple target tracking Milan et al. (2017), portfolio selection Das and Banerjee (2011) and safe reinforcement learning Chow et al. (2018), where decision-makers must make sequential decisions under uncertainty, balancing objectives such as cost, delivery time, and reliability without full knowledge of future demand or disruptions Kouvelis and Yu (1997); Bertsimas et al. (2011), where the min-max criterion naturally makes sense.
1.6 Our Contributions
  1. 1.
    Compared to the discussed prior work, which has primarily focused on bounding the ‘respective’ regret under worst-case or adversarial inputs, establishing bounds with the adversarial input on the min–max static regret (5) appears to be substantially more challenging. We highlight the difficulty for one instance in Section 3, where even when full functions $f_{t},g_{t}$ are revealed before choosing action $x_{t}$, the greedy algorithm still has a linear regret.
  2. 2.
    Thus, we consider the min-max static regret problem in the i.i.d. setting where both $f_{t},g_{t}$ are chosen independently across $t$ with identical (unknown) distribution $\mathcal{D}$. Note that all the discussed applications and use cases for multi-objective OCO (e.g. fair machine learning, fair resource allocation, multi-task learning, network scheduling, multi-objective online ride-matching problem etc.) are well motivated in the i.i.d. case as well. Moreover, analytically also, the problem remains challenging with the i.i.d. input. From a theoretical advancement point of view similar approach was followed for the COCO problem where the earliest results were derived for the i.i.d. input Mahdavi et al. (2013), which were then later generalized to the worst case input.
    For the i.i.d. model, the expected cost of $\mathcal{A}$ is ${\bar{C}}_{\mathcal{A}}$ that is just the expected value of $C_{\mathcal{A}}$ in (2), while the $C_{\mathsf{OPT}}$ simplifies to
    ${\bar{C}}_{\mathsf{OPT}}=T\cdot\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x)\quad\text{with}\quad(x^{\star},\theta^{\star})=\arg\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x),$ (11)
    where $\mu(x)={\mathbb{E}}\{f_{t}(x),g_{t}(x)\}$, $\theta=(\theta_{1},\theta_{2})\in\Delta_{2}$ and $\Delta_{2}$ is the 2D2-D simplex. The metric is now the expected min-max static regret
    ${\bar{R}}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}\{{\bar{C}}_{\mathcal{A}}-{\bar{C}}_{\mathsf{OPT}}\}.$ (12)
    Algorithm: For the i.i.d. model, the basic idea behind our solution approach for minimizing the expected min-max static regret (12) is as follows. From (11), the real goal for an algorithm is to find the optimal pair $(x^{\star},\theta^{\star})$. Since an online algorithm $\mathcal{A}$ does not know $\mathcal{D}$ ahead of time, finding $(x^{\star},\theta^{\star})$ is not possible. So we propose a simple algorithm that produces pair $(x_{t},\theta_{t})$ at time $t$ in an online manner that uses Hedge algorithm [^2] We describe the basic Experts problem and Hedge algorithm in Appendix E together with its regret guarantee for completeness purposes. Cesa-Bianchi and Lugosi (2006) to update $\theta_{t}$ to solve the inner maximization problem in (11) given $x_{t}$, and OGD to obtain actions $x_{t}$ given $\theta_{t}$ to solve the outer minimization (OCO) problem in (11) applied to the surrogate loss function $\theta_{t}^{\top}[f_{t}(x),g_{t}(x)]=\theta_{1t}f_{t}(x)+\theta_{2t}g_{t}(x)$ which is a proxy for $\theta^{\top}\mu(x)$, in an attempt to keep $(x_{t},\theta_{t})$ to stay close to $(x^{\star},\theta^{\star})$, so that the regret remains small.
    Note that (11) is not the usual online $\it min-max$} optimization problem Cardoso et al. (2019), where an online algorithm and adversary are playing against each other. Instead with (11) both $x$ and $\theta$ are algorithm’s choices depending on $\mathcal{D}$, that is under an adversarial control as defined in (12). The proposed alternating min-max algorithm is intuitive, however, the regret analysis is actually quite delicate, and heavily depends on our non-trivial decomposition of regret in (19) into three terms, which also motivates our algorithm.
    It is worth noting that there are several other candidate algorithms (e.g. primal-dual algorithm for COCO Yu et al. (2017) and OGD with surrogate cost function for COCO Sinha and Vaze (2024)) that appear to solve (11), however, analyzing them is rather challenging because of the outside $\max$ in the cost of an online algorithm in (5).
  3. 3.
    Guarantee: Using simple and elegant analysis via an insightful regret decomposition, we show that the expected min-max static regret (12) of the proposed algorithm is $O(\sqrt{T})$. This is similar to the short and elegant analysis of OGD algorithm’s Hazan (2019) performance for OCO in the adversarial input setting, that has been very useful for further progress. Even though we considered just two sequences $f_{t},g_{t}$ for simplicity, the result generalizes directly if there are $K$ sequences $f^{k}_{t},k=1,\dots,K$ and the expected min-max static regret (5) is $O(\sqrt{T\log K})$. See Remark 7 for more details.
  4. 4.
    Improved guarantee for GlobalExpertsProblem Even-Dar et al. (2009) in the i.i.d. input: As a special case of our result, we derive an improved regret guarantee of $O(\sqrt{T\log K})$ for the GlobalExpertsProblem Even-Dar et al. (2009) with the i.i.d. input compared to $O(\sqrt{T}\log K)$ of Even-Dar et al. (2009) which was applicable in the worst case/adversarial input setting.
  5. 5.
    With Bandit Feedback: We also extend our results to the bandit setting, where only function evaluations at the most recent actions are available without any gradient feedback. We show that the expected min-max static regret (12) is $O(T^{3/4}\sqrt{\log K})$ and $O(\sqrt{T\log K})$ for one-point and two-point bandit feedback, respectively, matching the best known bounds on OCO in the respective bandit settings.
1.7 Open Questions
  • We have only considered the i.i.d. input setting, and the major open question that remains is whether $o(T)$ min-max static regret (5) is achievable in the worst case input or not.
  • Lower bound on the expected min-max regret: Theoretically, deriving a lower bound on the expected min-max static regret matching the derived guarantee of $O(\sqrt{T})$ appears challenging, however, simulations (see Fig. 1) with $f_{t},g_{t}$ as random linear functions suggest that $\Theta(\sqrt{T})$ is the correct scaling for the expected min-max static regret.
  • Strongly Convex Functions: For OCO, the best-known static regret reduces to $O(\log{T})$ when functions $f_{t}$’s are strongly convex. Because of the use of Hedge in our proposed algorithm to update weights $\theta_{t}$ , its expected min-max static regret (12) remains $O(\sqrt{T})$ even if $f_{t},g_{t}$ are strongly convex. It is not clear whether this is a limitation of the proposed algorithm or the expected min-max static regret is $\Omega(\sqrt{T})$ even when $f_{t},g_{t}$ are strongly convex. Simulations (see Fig. 2) suggest that in fact expected min-max static regret is $\Omega(\sqrt{T})$ for the proposed algorithm.
  • Constrained Multi-Objective OCO: An immediate open question is how to solve the constrained version of the multi-objective OCO. For example, let $f_{t},g_{t}$ be the loss functions under the $\it min-max$} criteria while $h_{t}$ be the constraint function. Similar to COCO, let $\mathcal{X}^{\star}$ be the feasible set consisting of all admissible actions that satisfy all constraints $h_{t}(x)\leq 0,t\in[T]$. Then, the optimal expected static benchmark cost is
    ${\bar{C}}_{\mathsf{OPT}}=T\cdot\min_{x\in\mathcal{X}^{\star}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x)\quad\text{with}\quad(x^{\star},\theta^{\star})=\arg\min_{x\in\mathcal{X}^{\star}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x),$ (13)
    and the corresponding regret and CCV are
    ${\bar{R}}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}\{{\bar{C}}_{\mathcal{A}}-{\bar{C}}_{\mathsf{OPT}}\}\quad\bar{\textrm{CCV}}_{[1:T]}\equiv\sum_{t=1}^{T}{\mathbb{E}}\{\max(h_{t}(x_{t}),0)\},$ (14)
    respectively. It turns out that our proposed algorithm cannot be married readily with the prior best known approaches for COCO Sinha and Vaze (2024); Vaze and Sinha (2025); Sinha and Vaze (2025) for deriving simultaneous bounds on regret and CCV. Thus, how to solve the min-max OCO with constraints remains an interesting open problem.
2 Assumptions
We start by stating the assumptions that we will use throughout, that are standard in the OCO literature (Hazan, 2019).
Assumption 1 (Convexity).
$\mathcal{X}\subset{\mathbb{R}}^{d}$ is the admissible set that is closed, convex and has a finite Euclidean diameter $D$. Cost functions $f_{t},g_{t}:\mathcal{X}\mapsto\mathbb{R}$ are convex for all $t\geq 1$, and are bounded $|f_{t}(x)|\leq B,|g_{t}(x)|\leq B,\ \forall\ x\in\mathcal{X}$.
Assumption 2 (Lipschitzness).
[…]
Assumption 3.
Information Structure: On round $t,$ the online algorithm $\mathcal{A}$ first chooses an admissible action $x_{t}\in\mathcal{X}$ and then the adversary chooses two convex cost functions $f_{t}:\mathcal{X}\to\mathbb{R}$ and $g_{t}:\mathcal{X}\to\mathbb{R}$. Once the action $x_{t}$ has been chosen, both $f_{t}(x_{t}),\nabla f_{t}(x_{t})$ and $g_{t}(x_{t}),\nabla g_{t}(x_{t})$ are revealed, as is standard in the literature. We will also consider the bandit information structure in Section 6 where no gradient information is available.
3 min-max OCO Problem (5) in the Adversarial Setting
In this section, we highlight the difficulty in solving the min-max OCO Problem (5) in the adversarial input setting, where the input can be chosen adversarially depending on the choice of an online algorithm.
To illustrate the basic difficulty, we even allow the algorithm to know $f_{t},g_{t}$ completely before choosing its action $x_{t}$ at time $t$. Under this enhanced information setting, consider the greedy (locally optimal) algorithm that chooses
$x_{t}=\min_{x\in\mathcal{X}}\max\{f_{t}(x),g_{t}(x)\}.$
Recall that when $g_{t}=f_{t}\ \forall\ t$, i.e. in the OCO setting, this algorithm will have non-positive regret $R^{\text{s}ingle}_{\mathcal{A}}(T)$ (1) since $f_{t}(x_{t})\leq f_{t}(x)$ for any $x$. However, as we show next, the greedy algorithm suffers a linear min-max static regret (5).
Let $T=2N$ and consider the following linear functions on $\mathcal{X}=[0,1]$, for $j=1,\dots,N$
$f_{2j-1}(x)=1.2-0.2x,$ $\qquad f_{2j}(x)=x,$
$\hskip-72.26999ptg_{2j-1}(x)=x,$ $\qquad g_{2j}(x)=0.8+0.2x.$ (15)
Clearly,
$\arg\min_{x\in\mathcal{X}}\{\max(f_{2j-1}(x),g_{2j-1}(x))\}=1$ $\quad\text{and}\quad\arg\min_{x\in\mathcal{X}}\{\max(f_{2j}(x),g_{2j}(x))\}=0.$
Thus, the greedy algorithm chooses $x_{2j-1}=1$ and $x_{2j}=0,$ and the cost $C_{\mathcal{A}}$ (2) of the greedy algorithm is
$\max\left\{\sum_{j=1}^{N}\left(f_{2j-1}(x_{2j-1})+f_{2j}(x_{2j})\right),\sum_{j=1}^{N}\left(g_{2j-1}(x_{2j-1})+g_{2j}(x_{2j})\right)\right\}$
$=\max\left\{\sum_{j=1}^{N}\left(f_{2j-1}(1)+f_{2j}(0)\right),\sum_{j=1}^{N}\left(g_{2j-1}(1)+g_{2j}(0)\right)\right\}=\max\left\{\sum_{j=1}^{N}(1+0),\sum_{j=1}^{N}(1+0.8)\right\}=1.8N.$
However, if an algorithm chooses a single static action $x^{\star}=0$ for all times, its cost $C_{\mathcal{A}}$ (5) is $1.2N$, and hence
$C_{\mathsf{OPT}}\leq 1.2N.$
Thus, the min-max static regret (5) of the greedy algorithm is linear in $T$. This example highlights the key difficulty in minimizing the min-max static regret in the adversarial/worst case input. We note that this example does not preclude more intelligent algorithms that take a global view in trying to balance the two costs, such as primal-dual, OGD+Hedge or FTRL to achieve sub-linear min-max static regret (5), however, the stringent min-max static regret definition makes the analysis highly challenging. We will make additional remarks on the technical difficulty in Appendix F. Thus, from here on we consider the i.i.d. input that is defined next.
4 i.i.d. Input model
We consider the i.i.d. input model, where at time $t$, loss functions $(f_{t},g_{t})$ are chosen independently and are identically jointly distributed as $\mathcal{D}$ (unknown) across all $t$. We assume that Assumption 1 and 2 holds, and consider the same information structure as defined in Assumption 3.
Let $\lambda_{t}=(f_{t},g_{t})$ and recall that $\Delta_{2}$ is the 2D2-D simplex, $\theta=(\theta_{1},\theta_{2})\in\Delta_{2}$, and
$\Lambda_{t}(x,\theta):=\theta_{1}f_{t}(x)+\theta_{2}g_{t}(x).$
Following (2), the expected cost of any online algorithm $\mathcal{A}$ with the i.i.d. input which chooses actions $x_{t}$ is
${\bar{C}}_{\mathcal{A}}={\mathbb{E}}\left\{\max\Big\{\sum_{t=1}^{T}f_{t}(x_{t}),\sum_{t=1}^{T}g_{t}(x_{t})\Big\}\right\}={\mathbb{E}}\left\{\max_{\theta\in\Delta_{2}}\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta)\right\}.$ (16)
while from (11)
${\bar{C}}_{\mathsf{OPT}}=T\cdot\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x)\quad\text{with}\quad(x^{\star},\theta^{\star})=\arg\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x),$ (17)
where $\mu(x)={\mathbb{E}}\{f_{t}(x),g_{t}(x)\}$.
Consequently, the expected min-max static regret ${\bar{R}}_{T}(\mathcal{A})$ following (12) is
${\bar{R}}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}\left\{{\bar{C}}_{\mathcal{A}}-{\bar{C}}_{\mathsf{OPT}}\right\}.$ (18)
We refer to (18) as the i.i.d. min-max OCO problem and for the rest of the paper, consider it with the goal of deriving an algorithm that has $o(T)$ expected min-max static regret.
Remark 3.
Note that even i.i.d. min-max OCO problem is theoretically challenging, and no prior results apply directly.
Remark 4.
We note that only two loss function sequences are considered for simplicity of exposition since it captures the basic difficulty of the problem, and all the results that we derive apply directly to the case when there are $K$ function sequences (see details in Remark 7).
4.1 Algorithm
In this section, we present an algorithm to solve the i.i.d. min-max OCO problem (18). Towards that end, if suppose an online algorithm knew the optimizing $\theta^{\star}$ in (17), then by executing an OGD algorithm for OCO with a single function $\theta_{1}^{\star}f_{t}+\theta_{2}^{\star}g_{t}$ at time $t+1$, it would achieve a regret of $O(\sqrt{T})$ using standard results Hazan (2019). However, the optimizing $\theta^{\star}$ that depends on $\mathcal{D}$ (unknown) remains unknown, but, this observation motivates our algorithm that constructs
i) $\theta_{t}$ at time $t$ as a surrogate for $\theta^{\star}$ using the Hedge algorithm (gains version) treating the two function sequences $\{f_{t}\}_{t=1}^{T}$ and $\{g_{t}\}_{t=1}^{T}$ as the gains of the two experts (see Section E for basic definition of the Experts Problem and the Hedge algorithm), and
ii) for each fixed value of $\theta_{t}$, plays the action chosen by OGD applied to the single surrogate function $\theta_{t1}f_{t}+\theta_{t2}g_{t}$ at time $t+1$. Pseudocode is given in Algorithm 1.
Algorithm 1 Minimization of the Maximum of The Sum of Two Convex Loss Sequences
1:Input: step sizes $\eta_{x,t}>0$, $\eta_{\theta,t}>0$, feasible set $\mathcal{X}$
2:Initialize: $w_{1}=(1,1)$, $\theta_{1}=w_{1}/\|w_{1}\|_{1}$, $f_{0}\equiv 0,g_{0}\equiv 0$ and play arbitrary $x_{0}\in\mathcal{X}$
3:for $t=1,\dots,T$ do
4:  Observe $(f_{t-1}(x_{t-1}),g_{t-1}(x_{t-1}))$ and $(\nabla f_{t-1}(x_{t-1}),\nabla g_{t-1}(x_{t-1}))$
5:  Using current $\theta_{t}$, form convex function $x\mapsto{\tilde{\Lambda}}_{t}(x,\theta_{t}):=\theta_{t,1}f_{t-1}(x)+\theta_{t,2}g_{t-1}(x)$
6:  OGD: Gradient step: $y_{t}=x_{t-1}-\eta_{x,t}\nabla{\tilde{\Lambda}}_{t}(x_{t-1},\theta_{t})$ and find $x_{t}=\text{Proj}_{\mathcal{X}}(y_{t})$
7:  Play action $x_{t}$
8:  Evaluate $\lambda_{t}=(f_{t}(x_{t}),g_{t}(x_{t}))$
9:  Hedge: Update weights: $w_{t+1,i}=w_{t,i}\exp(\eta_{\theta,t}\lambda_{t,i})$ for $i=1,2$
10:  Normalize: $\theta_{t+1}=w_{t+1}/\|w_{t+1}\|_{1}$
11:end for
4.2 Guarantee
Theorem 4.
Under Assumptions 1 and 2, for $\eta_{\theta,t}\leq\sqrt{\frac{2\ln 2}{B^{2}t}},\eta_{x,t}=\frac{D}{G\sqrt{t}}$, the expected min-max static regret (18) of the proposed algorithm (Algorithm 1) is
${\bar{R}}_{\mathcal{A}}(T)=O(\sqrt{T}).$
Proof.
The main ingredient of the proof as well as the inspiration for Algorithm 1 is the following decomposition (19) of the expected min-max static regret (18). Recall the definition of
$\Lambda_{t}(x,\theta):=\theta_{1}f_{t}(x)+\theta_{2}g_{t}(x).$
where $\theta=(\theta_{1},\theta_{2})\in\Delta_{2}$.
By adding and subtracting two terms, $\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t})$ and $\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta_{t})$ inside the expectation in the RHS of (18), we write the expected min-max static regret (18) as
${\bar{R}}_{T}(\mathcal{A})=\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{1}\}+{\mathbb{E}}\{R_{2}\}+{\mathbb{E}}\{R_{3}\}\}\leq\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{1}\}\}+\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{2}\}\}+\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{3}\}\},$ (19)
where
$R_{1}=\max_{\theta\in\Delta_{2}}\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta)-\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t}),$
$R_{2}=\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t})-\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta_{t}),$
$R_{3}=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta_{t})-C_{\mathsf{OPT}}=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta_{t})-\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta).$
The interpretation of $R_{1}$ is that it is the regret for the gains version of $2$-Experts Problem Cesa-Bianchi and Lugosi (2006), where the two function sequences $f_{t},g_{t}$ are the gains of two experts, with respect to an optimal $\theta^{\star}\in\Delta_{2}$ for fixed actions $x_{t}$ across time slots. This motivates the use of Hedge algorithm Cesa-Bianchi and Lugosi (2006) (gains version) in our proposed algorithm that is tailor-made for updating $\theta_{t}$ for fixed actions $x_{t}$.
Similarly, $R_{2}$ is the static regret for an OCO minimization problem relative to static optimal action $x^{\star}$ for a single surrogate objective function sequence of $\Lambda_{t}(x,\theta_{t})=\theta_{t1}f_{t}(x)+\theta_{t2}g_{t}(x)$ for a fixed $\theta_{t}$. Hence using an OGD algorithm is natural for this purpose as done in the algorithm for each time $t$, with a fixed weight $\theta_{t}$.
The third term $R_{3}$ is more nuanced and captures the difference between the cost of $\mathsf{OPT}$ for a fixed sequence of $\theta_{t},t=1,\dots,T$ and the optimal benchmark cost ${\bar{C}}_{\mathsf{OPT}}$ (17), respectively.
We will bound $R_{1}$ and $R_{2}$ sample path wise for each realization of $f_{t},g_{t}$, while for $R_{3}$ we will bound ${\mathbb{E}}\{R_{3}\}$. In particular, we will show the following that is proven in the Appendix.
Lemma 5.
For $\eta_{\theta,t}\leq\sqrt{2\ln 2/(B^{2}t)}$, $\sup_{\mathcal{D}}\{R_{1}\}\leq O(\sqrt{T})$.
Lemma 6.
Choosing $\eta_{x,t}=\frac{D}{G\sqrt{t}}$, $\sup_{\mathcal{D}}\{R_{2}\}\leq O(\sqrt{T})$.
Lemma 7.
${\mathbb{E}}\{R_{3}\}\leq 0$.
Because of the connection described above, the proof for $R_{1}$ in Lemma 5 and for $R_{2}$ in Lemma 6 follows standard analysis Hazan (2019) for Hedge and OGD, respectively. Bounding $R_{3}$ needs more care and we prove it in Appendix C where we critically exploit the i.i.d. input model.
Combining Lemma 5, 6 and 7, and the regret decomposition (19), we get that the expected min-max static regret (12) of Algorithm 1 with
$\eta_{\theta,t}\leq\sqrt{\frac{2\ln 2}{B^{2}t}},\qquad\ \text{and}\qquad\eta_{x,t}=\frac{D}{G\sqrt{t}}$ (20)
is
${\bar{R}}_{\mathcal{A}}(T)=O(\sqrt{T}),$ (21)
proving the result. ∎
Following remarks are in order.
Remark 5.
The regret decomposition (19) does multiple things at the same time i) identifies the critical difficulty of the problem, ii) motivates a natural algorithm, and ii) keeps the proof modular, where the regret bound for each module is simple and elegant, almost first principles based.
Remark 6.
Although the proof of Theorem 4 is remarkably simple and elegant, it hinges on the critical regret decomposition (19) – which is not obvious a priori. This decomposition eliminates the need for sophisticated technical machinery needed in related prior work Mahdavi et al. (2013) that also considers the i.i.d. input and highlights the underlying simplicity of the result. We would like to emphasize that the elegance of the argument should not obscure its importance.
Remark 7.
When there are $K$ loss function sequences $f_{t}^{k},k=1,\dots,K$ that are chosen i.i.d. from $\mathcal{D}$, define $\theta\in\Delta_{K}$, (where $\Delta_{K}$ is the $K$-dimensional simplex) and $\Lambda_{t}(x,\theta)=\sum_{k=1}^{K}\theta_{k}f_{t}^{k}$. Then the expected min-max static regret of $\mathcal{A}$ is
${\bar{R}}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}{\mathbb{E}}\left\{\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta)-T\cdot\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{K}}\theta^{\top}\mu(x)\right\}$ (22)
where $\mu(x)={\mathbb{E}}\{f_{t}^{1}(x),\dots,f_{t}^{K}(x)\}$.
With this new definition of $\theta$ and $\Lambda_{t}$, Algorithm 1 remains as is with Hedge choosing $\theta_{t}\in\Delta_{K}$ while $x_{t}$ the single action chosen by OGD. Assuming all $f_{t}^{k},k=1,\dots,K$ satisfy Assumption 1 and 2, the only change in the guarantee with $K>2$ compared to Theorem 4 with $K=2$ is because of Lemma 5 that is now $O(\sqrt{T\log K})$ Hazan (2019) since there are $K$ experts corresponding to the $K$ loss function sequences $f_{t}^{k},k=1,\dots,K$ . The OGD guarantee of Lemma 6 stays the same since $\|\nabla\Lambda_{t}(x_{t},\theta)\|\leq G$ as before. Moreover, Lemma 7 holds as is. Thus, the overall expected min-max static regret of Algorithm 1 with $K$ loss function sequences is
$O(\sqrt{T\log K}).$ (23)
Remark 8.
We note that Algorithm 1 is well motivated even while considering the worst case/adversarial input. The technical problem in its analysis with the worst case/adversarial input is that Lemma 7 that is the critical result fails when considering the adversarial input. We highlight this difficulty in Appendix F and show that the ‘best possible’ upper bound on $R_{3}$ is $R_{3}=O(T^{3/2})$ with the worst case/adversarial input for Algorithm 1 by exploiting just the properties of the Hedge algorithm which is insufficient for achieving sub-linear regret.
Remark 9.
[Extension to Strongly Convex $f_{t}$’s and $g_{t}$’s] One limitation of our proposed algorithm and the proof is that because of the use of the Hedge algorithm for updating $\theta_{t}$, the regret $\sup_{\mathcal{D}}\{R_{1}\}$ can at best be $O(\sqrt{T})$. Thus, unlike OCO, where OGD algorithm achieves logarithmic regret in $T$ when $f_{t}$’s are strongly convex, our algorithm’s regret is only $O(\sqrt{T})$ even when both $f_{t},g_{t}$ are strongly convex. Is this a limitation of the algorithm or of that of the problem itself remains an interesting unresolved question.
5 Improving the result on GlobalExpertsProblem with the i.i.d. input Even-Dar et al. (2009)
Let $\mathcal{X}=\Delta_{K}$, the $K$-dimensional simplex, and writing $f_{t}^{k}(x)=x_{k}\ell_{k}(t)$, where $x_{k}$ is the $k$-th coordinate of $x\in\mathcal{X}$, $\ell_{k}(t)$ is the loss corresponding to the $k$-th expert, and $k=1,2\dots,K$, we get the GlobalExpertsProblem Even-Dar et al. (2009) with $K$ experts in the i.i.d. input case is a special case of the i.i.d. min-max OCO problem (18).
Thus, from (23), we get that the expected min-max static regret of Algorithm 1 is $O(\sqrt{T\log K})$ for the GlobalExpertsProblem with i.i.d. input. In comparison, Even-Dar et al. (2009) derived a min-max static regret (5) guarantee of $O(\sqrt{T}\log K)$ with adversarial input for the following algorithm called MULTI:
For $K=2$,
$x_{1}(t+1)=x_{1}(t)+\frac{x_{2}(t)\ell_{2}(t)-x_{1}(t)\ell_{1}(t)}{\sqrt{T}},$
and $x_{2}(t)=1-x_{1}(t)$. For general $K$, the $K=2$ algorithm was used recursively. Compared to our Algorithm 1, the MULTI algorithm Even-Dar et al. (2009) is very compute intensive since it recursively uses an algorithm for $K=2$ and requires the knowledge of time horizon $T$.
6 Extension to Bandit Feedback
We now analyze the expected min-max OCO problem (18) under the bandit feedback, where at time $t$, once the online algorithm $\mathcal{A}$ chooses its action $x_{t}$, it only observes the scalar values $\{f_{t-1}^{k}(.)\}_{k=1}^{K}$ at one (one-point model) or two points (two-point model) and does not receive any gradient information.
Algorithm 2 Bandit-MinMax (unified one-point/two-point variant)
1:Feasible Set $\mathcal{X}\subset\mathbb{R}^{d}$, smoothing parameters $\{\delta_{t}\}$, step-sizes $\{\eta_{x,t}\},\{\eta_{\theta,t}\}$, feedback type $\mathsf{mode}\in\{\text{one-point},\text{two-point}\}$
2:Initialize $w_{1}=\mathbf{1}_{K}$, $\theta_{1}=w_{1}/\|w_{1}\|_{1}$, choose $x_{0}\in X$
3:Define $\Lambda_{t}(x,\theta_{t})=\sum_{k}\theta_{t,k}f_{t}^{k}(x)$
4:for $t=1,\dots,T$ do
5:  Sample $u_{t}$ uniformly from unit sphere $\mathbb{S}^{d-1}$
6:  if $\mathsf{mode}=\text{one-point}$ then
7:   Play $\tilde{x}_{t}=\text{Proj}_{\mathcal{X}}(x_{t-1}+\delta_{t}u_{t})$ and observe $\{f_{t}^{k}(\tilde{x}_{t})\}_{k=1}^{K}$
8:   Gradient estimator:
$\widehat{\nabla}_{t}=\frac{d}{\delta_{t}}\Lambda_{t}(\tilde{x}_{t},\theta_{t})\,u_{t}$
9:  else if $\mathsf{mode}=\text{two-point}$ then
10:   Query $x_{t-1}+\delta_{t}u_{t}$ and $x_{t-1}-\delta_{t}u_{t}$, observe $\{f_{t}^{k}(x_{t-1}\pm\delta_{t}u_{t})\}_{k=1}^{K}$
11:   Gradient estimator:
$\widehat{\nabla}_{t}=\frac{d}{2\delta_{t}}\Big(\Lambda_{t}(x_{t-1}+\delta_{t}u_{t},\theta_{t})-\Lambda_{t}(x_{t-1}-\delta_{t}u_{t},\theta_{t})\Big)u_{t}$
12:  end if
13:  Gradient step: $y_{t}=x_{t-1}-\eta_{x,t}\widehat{\nabla}_{t}$,    $x_{t}=\text{Proj}_{\mathcal{X}}(y_{t})$
14:  Hedge update: $w_{t+1,k}=w_{t,k}\exp\big(\eta_{\theta,t}f_{t,k}(x_{t})\big)$,    $\theta_{t+1}=w_{t+1}/\|w_{t+1}\|_{1}$
15:end for
6.1 Algorithm
For this bandit model, we consider a simple adaption of Algorithm 1 by replacing its OGD part with the bandit information counterpart, while keeping the Hedge part as it is, given by Algorithm 2. Moreover, the analysis also follows the same script by replacing the guarantees by their respective bandit counterparts.
6.2 Guarantee
Theorem 8.
Under Assumption 1 and 2,
  1. 1.
    (One-point feedback) With $\delta_{t}=t^{-1/4}$, $\eta_{x,t}=D/(Gd^{1/2}t^{3/4})$, and $\eta_{\theta,t}=\sqrt{\log K}/(B\sqrt{t})$, the expected min-max static regret satisfies
    ${\bar{R}}_{\mathcal{A}}(T)\leq c_{1}\,d^{1/2}GD\,T^{3/4}\;+\;c_{2}B\sqrt{T\log K},$
    for universal constants $c_{1},c_{2}>0$.
  2. 2.
    (Two-point feedback) With $\delta_{t}=t^{-1/2}$, $\eta_{x,t}=D/(Gd^{1/2}\sqrt{t})$, and $\eta_{\theta,t}=\sqrt{\log K}/(B\sqrt{t})$, the expected min-max static regret satisfies
    ${\bar{R}}_{\mathcal{A}}(T)\leq c_{3}\,d^{1/2}GD\,\sqrt{T}\;+\;c_{4}B\sqrt{T\log K},$
    for universal constants $c_{3},c_{4}>0$.
6.3 Proof of Theorem 8
Recall the decomposition (19)
${\bar{R}}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{1}\}+{\mathbb{E}}\{R_{2}\}+{\mathbb{E}}\{R_{3}\}\}\leq\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{1}\}\}+\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{2}\}\}+\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{3}\}\}.$
Bounding $R_{1}$ and $R_{3}$.
For both one-point or two-point bandit feedback, the Hedge guarantee remains the same as in Lemma 5. In particular, with $\eta_{\theta,t}=\sqrt{\log K}/(B\sqrt{t})$, sample path wise,
$R_{1}\leq B\sqrt{2T\log K},$
and similar to Lemma 7,
${\mathbb{E}}\{R_{3}\}\leq 0.$
The only change is in bounding $R_{2}$.
Bounding $R_{2}$ (one-point feedback). The one-point estimator
$\widehat{\nabla}_{t}=\frac{d}{\delta_{t}}\Lambda_{t}(\tilde{x}_{t},\theta_{t})\,u_{t}$
is unbiased for the gradient of the $\delta_{t}$-smoothed loss, with variance at most $O(dB^{2}/\delta_{t}^{2})$. Using $\delta_{t}=t^{-1/4}$ and $\eta_{x,t}=D/(Gd^{1/2}t^{3/4})$, the bandit-OGD analysis Flaxman et al. (2005) yields
$R_{2}\leq c_{1}\,d^{1/2}GD\,T^{3/4},$
for some universal constant $c_{1}$.
Combining the three bounds we get
${\bar{R}}_{\mathcal{A}}(T)\leq c_{1}\,d^{1/2}GD\,T^{3/4}\;+\;2B\sqrt{2T\log K}.$
Bounding $R_{2}$ (two-point feedback). With two function evaluations per round, the estimator
$\widehat{\nabla}_{t}=\tfrac{d}{2\delta_{t}}\Big(\Lambda_{t}(x_{t-1}+\delta_{t}u_{t},\theta_{t})-\Lambda_{t}(x_{t-1}-\delta_{t}u_{t},\theta_{t})\Big)u_{t}$
is unbiased with variance $O(dB^{2}/\delta_{t}^{2})$. Choosing $\delta_{t}=t^{-1/2}$ and $\eta_{x,t}=D/(Gd^{1/2}\sqrt{t})$, the two-point bandit-OGD analysis Agarwal et al. (2010); Duchi et al. (2015); Shamir (2017) gives
$R_{2}\leq c_{3}\,d^{1/2}GD\,\sqrt{T},$
for some universal constant $c_{3}$. Thus, with two-point feedback,
${\bar{R}}_{\mathcal{A}}(T)\leq c_{3}\,d^{1/2}GD\,\sqrt{T}\;+\;2B\sqrt{2T\log K}.$
Discussion: Compared to the limited information setting considered in Section 2, where gradient information is available, with bandit feedback the regret guarantees suffer for the expected min-max OCO problem, however, the degraded guarantees match the best known guarantees for OGD in the bandit setting. The proof of the result is simple because of the modularity of our algorithm and the regret decomposition that decomposes the total regret into regret resulting from Hedge and OGD, and the fact that only regret for OGD changes with bandit feedback.
7 Experimental Results
In this section, we present numerical results on the expected min-max static regret of Algorithm 1 under relevant settings, that are well motivated from both theoretical as well applications point of view. In all the figures, we plot the expected min-max static regret as a function of time horizon $T$.
7.1 Basic Results
We begin our experimental results by considering the input model when all functions $f_{t}^{k}$ for $k=1,\dots,K,t=1,\dots,T$ are random linear functions in $d=10$ dimensions, where each coefficient is distributed uniformly between $[0,1]$. Random linear functions are important since they were sufficient to find lower bounds on the regret of any online algorithm for OCO Hazan (2019). In Fig. 1, we plot the expected min-max static regret of Algorithm 1 as a function of $K$ and $T$ for random linear functions, where as expected, we observe that the regret scales as $O(\sqrt{T\log K})$.
Next, we consider $f_{t}^{k}$’s as strongly convex functions, where each $f_{t}^{k}$ is independently and identically selected from a set of $\{(x-a)^{2},a=[-9,-8,\dots,0,1,\dots,9]\}$. Recall that for OCO with strongly convex functions, the step size $\eta$ needed for OGD to get a regret guarantee of $O(\log T)$ scales as $\eta=1/t$ and not $\eta=1/\sqrt{t}$ Hazan (2019). Thus, for this simulation of Algorithm 1333In contrast to $\eta_{x,t}=O(1/\sqrt{t})$ as needed to prove Theorem 4. we use $\eta_{\theta,t}=O(1/\sqrt{t})$ and $\eta_{x,t}=O(1/t)$.  [^4] We have checked that other combinations only increase the regret. As pointed out earlier, we are unable to improve the expected min-max static regret guarantee for Algorithm 1 when all functions $f_{t}^{k}$’s are strongly convex. We see that in fact that is true for Algorithm 1 even in simulations as shown in Fig. 2, where for $K=1$, which is the same as OCO (with $\eta_{x,t}=O(1/t)$) the regret scales as $\log(T)$, while for larger values of $K$, the expected min-max static regret scales as $O(\sqrt{T})$ when $\eta_{\theta,t}=O(1/\sqrt{t})$ and $\eta_{x,t}=O(1/t)$. Thus, if indeed expected min-max static regret of $O(\log T)$ is achievable then Algorithm 1 needs to be conceptually modified.
050501001001501502002002502503003000202040406060Time ($t$)Average Cumulative RegretExpected min-max static regret for random linear functionsK=1K=1K=10K=10K=50K=50K=100K=100
Figure 1: Expected min-max static regret of Algorithm 1 when $f_{t}^{k}$ are random linear functions for different values of $K$.
0505010010015015020020025025030030002002004004006006008008001,0001{,} [^0] Time Horizon ($T$)Average RegretExpected min-max static regret for quadratic loss functionsK=1K=1K=10K=10K=50K=50K=100K=100
Figure 2: Expected min-max static regret of Algorithm 1 when $f_{t}^{k}$ are random quadratic functions for different values of $K$.
7.2 GlobalExpertsProblem
In this section, we compare the performance of Algorithm 1 and MULTI Even-Dar et al. (2009) for the GlobalExpertsProblem with the i.i.d. input. For that purpose, we consider that the loss of each expert at time $t$ is drawn independently from a uniform distribution:
$\ell_{t,k}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathsf{Uniform}(a,b),\quad t=1,2,\dots,T,\;k=1,2,\dots,K,$
where $a,b\in[0,1]$ are fixed parameters defining the range of losses. In our experiments, we set $a=0.2$ and $b=0.8$ to ensure non-trivial loss variation while avoiding extreme loss values.
For this model, we plot the expected min-max static regret performance of Algorithm 1 and MULTI in Fig. 3, and see that both perform similarly, even though the regret guarantee we obtained for Algorithm 1 in i.i.d. case is slightly better than MULTI (which is valid even with the adversarial input). It appears that MULTI has better performance than its guarantee Even-Dar et al. (2009). However, MULTI is computationally far more intensive than Algorithm 1 since it uses the $K=2$ primitive recursively, and moreover, it needs to know the time horizon $T$ in contrast to Algorithm 1.
1,0001{,}0002,0002{,}0004,0004{,}0008,0008{,}0000551010151520202525303035354040K=2K=2K=4K=4K=16K=16K=32K=32K=64K=64Time horizon TTAverage RegretMULTIAlgorithm 1
Figure 3: Expected min-max static regret of MULTI Even-Dar et al. (2009) and Algorithm 1 for different numbers of experts $K$ and time horizons $T$ for the fair classification example.
7.3 Application of min-max formulation to Fair Classification Setting
Finally, in this subsection, we simulate the i.i.d. min-max OCO problem for a fair classification setting.
Data generation.
We consider $K$ groups, each corresponding to a convex loss function sequence. For each group $k\in[K]$, we sample a hidden parameter $w_{k}^{\star}\sim\mathcal{N}(0,I_{d})$. At each round $t=1,\dots,T$, for each group $k$ we draw $m$ feature vectors
$z_{t,k,i}\sim\mathcal{N}(\mu_{k},\sigma^{2}I_{d}),\quad i=1,\dots,m,$
with $\mu_{k}$ drawn once per group. Labels are generated as
$y_{t,k,i}\sim\mathrm{Bernoulli}\!\left(\sigma(\langle w_{k}^{\star},z_{t,k,i}\rangle)\right),$
where $\sigma(u)=1/(1+e^{-u})$ is the sigmoid function.
Loss functions.
The per-group convex loss at round $t$ is the averaged logistic loss with $L_{2}$ regularization:
$f_{t}^{k}(x)\;=\;\frac{1}{m}\sum_{i=1}^{m}\log\!\Big(1+\exp(-y_{t,k,i}\,\langle x,z_{t,k,i}\rangle)\Big)\;+\;\kappa\|x\|_{2}^{2}.$
For this application, we compare the performance of following algorithms.
  • Algorithm 1 (OGD+Hedge)
  • OGD on average loss: projected OGD on $\tfrac{1}{K}\sum_{k=1}^{K}f_{t}^{k}(x)$.
  • Follow-the-Regularized-Leader (FTRL): $x_{t}=\arg\min_{x\in\mathcal{X}}\left\{\max_{k}\sum_{s=1}^{t-1}f_{s}^{k}(x)+\mathcal{R}(x)\right\}$.
Parameters.
We use $d=20$, $K=10$, mini-batch size $m=50$, diameter $D=5$, and regularization $\kappa=10^{-3}$. Step sizes are set as $\eta_{x,t}=D/(G\sqrt{t})$ for OGD (with $G$ an upper bound on gradient norms) and $\eta_{\theta,t}=\sqrt{\ln K}/(B\sqrt{t})$ for Hedge, with $B$ an upper bound on per-round losses.
In Fig. 4, we plot the expected min-max static regret ${\bar{R}}_{T}$ for all the three algorithms. We see that all algorithms have sub-linear regret, and Algorithm 1 has larger regret than FTRL but better than average OGD. However, recall that neither FTRL and average OGD have any theoretical guarantees on their expected min-max static regret. Next, in Fig. 5, we will consider an input where the index of the most challenging sequence changes over time and show that Algorithm 1 outperforms FTRL in that case.
10010020020030030040040002020404060608080Time Horizon ($T$)Average Regret vs. Offline OptimalExpected min-max static regret for $K=10$ for the fair classification problemAlgorithm 1 OGD+HedgeAvg-OGDFTL
Figure 4: Expected min-max static regret comparison for $K=10$.
Non-stationary input
Next, we change the input described in Section 7.3 to model a highly non‑stationary environment specifically designed to challenge static optimization strategies such as FTRL and highlight the adaptability of Algorithm 1 that uses OGD+Hedge. The detailed changes with respect to Section 7.3 are as follows:
  • We consider $K=3$ separate objectives, each with its own logistic loss function defined by fixed but distinct parameter vectors $w_{\text{star}}[k]$ and mean feature shifts $\mu[k]$.
  • Every switch_interval steps (here set to 100), one objective becomes significantly harder by adding a large mean shift ($\texttt{shift\_magnitude}=5.0$) to its feature distribution. This creates abrupt changes in which objective dominates the cumulative loss objective function.
  • Data generation: For each round $t$, and for each objective $k$, a feature matrix $Z$ of size $m\times d$ is generated from a Gaussian distribution centered at the shifted mean. Labels $y$ are generated from a logistic model using $w_{\text{star}}[k]$ (is a hidden “true” parameter vector for objective $k$. It defines how that objective behaves.), with additional noise introduced via random sampling.
We clearly see in Fig. 5 that Algorithm 1 (OGD+Hedge) adapts better to changing conditions than FTRL, since worst‑case objective changes over time, making past loss information misleading for FTRL.
1010505010010030030050050002020404060608080Time Horizon ($T$)Average Regret (Avg. over 10 Seeds)Expected min-max static regret in switching environment for K=3K=3Algorithm 1 OGD+HedgeFTRL
Figure 5: Expected min-max static regret for Algorithm 1 vs. FTRL in a piecewise-stationary environment.
8 Conclusions
In this paper, we have made a basic advance in the area of multi-objective OCO by formulating the min-max static regret minimization problem, that requires an algorithm to closely track multiple loss function sequences simultaneously and at all times. Because of analytical challenges with the adversarial input, we considered the i.i.d. input setting, for which we proposed a simple algorithm that used a combination of the well known Hedge and OGD algorithm. Using an insightful regret decomposition, which made the analysis modular, we derived a $O(\sqrt{T\log K})$ regret bound on the proposed algorithm, which is shown to be best possible via simulations. We also extended the results to the bandit information setting. Certain questions that remain open are: i) what is the best regret possible with the adversarial input, ii) how to handle the case when all loss functions are strongly convex, and iii) handling constraints like COCO.
Appendix A Proof of Lemma 5
Let $S_{1}:=\sum_{t=1}^{T}f_{t}(x_{t})$ and $S_{2}:=\sum_{t=1}^{T}g_{t}(x_{t})$, and recall that $\lambda_{t}=(f_{t}(x_{t}),g_{t}(x_{t}))$ and $\sum_{t}\Lambda_{t}(x_{t},\theta_{t})=\theta_{t1}f_{t}+\theta_{t2}g_{t}=\sum_{t}\langle\theta_{t},\lambda_{t}\rangle$. Then
$\max_{\theta}\sum_{t}\Lambda_{t}(x_{t},\theta)=\max\{S_{1},S_{2}\}.$
Standard Hedge (gains version) [7] analysis implies that for any expert $i=1,2$.
$\sum_{t=1}^{T}\langle\theta_{t},\lambda_{t}\rangle\geq S_{i}-\frac{\ln 2}{\eta_{\theta,T}}-\sum_{t=1}^{T}\frac{\eta_{\theta,t}}{8}(\lambda_{t,1}-\lambda_{t,2})^{2}.$
Maximizing over $i$ and using $(\lambda_{t,1}-\lambda_{t,2})^{2}\leq(2B)^{2}=4B^{2}$ (since values are bounded by $B$ from Assumption 1) yields
$\sum_{t=1}^{T}\langle\theta_{t},\lambda_{t}\rangle\geq\max\{S_{1},S_{2}\}-\frac{\ln 2}{\eta_{\theta,T}}-\sum_{t=1}^{T}\frac{\eta_{\theta,t}4B^{2}}{8}.$
Rearranging gives
$R_{1}=\max_{\theta\in\Delta_{2}}\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta)-\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t})=\max\{S_{1},S_{2}\}-\sum_{t=1}^{T}\langle\theta_{t},\lambda_{t}\rangle\leq\frac{\ln 2}{\eta_{\theta,T}}+\sum_{t=1}^{T}\frac{\eta_{\theta,t}B^{2}}{2}.$
Appendix B Proof of Lemma 6
Define $\tilde{f}_{t}(x):=\Lambda_{t}(x,\theta_{t})$, which is convex for fixed $\theta_{t}$. Thus, the standard OGD (projected) guarantee from [19] implies, for any $x\in\mathcal{X}$,
$\sum_{t=1}^{T}\tilde{f}_{t}(x_{t})-\sum_{t=1}^{T}\tilde{f}_{t}(x)\leq\frac{\|x-x_{0}\|^{2}}{2\eta_{x,T}}+\sum_{t=1}^{T}\frac{\eta_{x,t}}{2}\|\nabla{\tilde{f}}_{t}(x_{t-1})\|^{2}.$
Using the finite diameter $\|x-x_{0}\|\leq D$ and the bounded gradient $\|\nabla{\tilde{f}}_{t}(x_{t-1})\|\leq G$ (since both $\|\nabla f_{t}(x_{t-1})\|\leq G$ and $\|\nabla g_{t}(x_{t-1})\|\leq G$ from Assumption 1) yields
$R_{2}\leq\frac{D^{2}}{2\eta_{x,T}}+\sum_{t=1}^{T}\frac{\eta_{x,t}G^{2}}{2}.$
With the choice $\eta_{x,t}=D/(G\sqrt{t})$
$R_{2}\leq DG\sqrt{T}.$
Appendix C Proof of Lemma 7
Recall that
$\lambda_{t}(x):=(f_{t}(x),g_{t}(x)),\qquad\text{and let}\quad\ \mu(x):=\mathbb{E}\{\lambda_{t}(x)\}.$
where $\mu(x)$ is independent of index $t$ on account of the i.i.d. assumption on functions $f_{t},g_{t}$. Let $\mathcal{F}_{t-1}$ denote the filtration generated by
$\mathcal{F}_{t-1}:=\sigma\big(x_{1},\dots,x_{t-1},\theta_{1},\dots,\theta_{t-1},\lambda_{1}(x_{1}),\dots,\lambda_{t-1}(x_{t-1})\big),$
which determines $\theta_{t}$ at round $t$.
Recall the definition of optimal pair $x^{\star}$ and $\theta^{\star}$ that defines the static offline optimal solution of the expected min-max problem (17):
$x^{\star}:=\arg\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\sum_{t=1}^{T}\theta^{\top}\mu(x),\quad\text{and}\quad\theta^{\star}:=\arg\max_{\theta\in\Delta_{2}}\sum_{t=1}^{T}\theta^{\top}\mu(x^{\star}).$ (24)
By the definition of $R_{3}$,
$R_{3}$ $=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\theta_{t}^{\top}\lambda_{t}(x)-\min_{x\in\mathcal{X}}\max_{\theta}\sum_{t=1}^{T}\theta^{\top}\lambda_{t}(x)$
$\stackrel{{\scriptstyle(a)}}{{=}}\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\theta_{t}^{\top}\lambda_{t}(x)-\sum_{t=1}^{T}\theta^{\star}{}^{\top}\lambda_{t}(x^{\star}),$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{t=1}^{T}\theta_{t}^{\top}\lambda_{t}(x^{\star})-\sum_{t=1}^{T}\theta^{\star}{}^{\top}\lambda_{t}(x^{\star}),$
$=\sum_{t=1}^{T}(\theta_{t}-\theta^{\star})^{\top}\lambda_{t}(x^{\star}),$ (25)
where $(a)$ follows from the definition of $x^{\star}$ and $(b)$ follows since $x^{\star}\in\mathcal{X}$ is feasible. Moreover, we have
$\mathbb{E}\left\{(\theta_{t}-\theta^{\star})^{\top}\lambda_{t}(x^{\star})\,\Big|\,\mathcal{F}_{t-1}\right\}=(\theta_{t}-\theta^{\star})^{\top}{\mathbb{E}}\{\lambda_{t}(x^{\star})|\mathcal{F}_{t-1}\}=(\theta_{t}-\theta^{\star})^{\top}\mu(x^{\star}),$ (26)
where the first equality follows since $\theta_{t}$ is completely determined by $\mathcal{F}_{t-1}$, while the second equality follows because ${\mathbb{E}}\{\lambda_{t}(x^{\star})|\mathcal{F}_{t-1}\}=\mu(x^{\star})$ from the i.i.d. assumption.
We also have
$(\theta_{t}-\theta^{\star})^{\top}\mu(x^{\star})\leq 0$
for each $t=1,\dots,T$ using the optimality of $\theta^{\star}$ (24) since
$\theta^{\star}\in\arg\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x^{\star})$
depends only on the expected vector $\mu(x^{\star})$ and not on any particular realization of $\lambda_{t}(x^{\star})$.
Thus, we get from (26)
$\mathbb{E}\left\{(\theta_{t}-\theta^{\star})^{\top}\lambda_{t}(x^{\star})\,\Big|\,\mathcal{F}_{t-1}\right\}\leq 0,$ (27)
Using (27) in (25) and taking the expectation (using the tower property of conditional expectation),
$\mathbb{E}\{R_{3}\}\leq\sum_{t=1}^{T}(\theta_{t}-\theta^{\star})^{\top}\mu(x^{\star})\leq 0.$ (28)
Appendix D Per-slot $\it{\it min-max}$} multi-objective optimization [24]
In this section, we show that benchmark $W_{L}^{T}$(10) can be $3/2$ times more than our benchmark $C_{\mathsf{OPT}}$ (4).
Input
Let $T=2N$ and $\mathcal{X}=[0,1]$. For $j=1,\dots,N$ let
$f_{2j-1}(x)=1.2-0.2x,\quad f_{2j}(x)=x,\qquad g_{2j-1}(x)=x,\quad g_{2j}(x)=0.8+0.2x.$
Odd time slots: $\min_{x\in[0,1]}\max\{1.2-0.2x,x\}=1$ (attained at $x=1$). Even time slots rounds: $\min_{x\in[0,1]}\max\{x,0.8+0.2x\}=0.8$ (attained at $x=0$). Thus the benchmark $W_{L}^{T}$(10) is
$\sum_{t=1}^{T}\min_{x\in\mathcal{X}}\max\{f_{t}(x),g_{t}(x)\}=N(1+0.8)=1.8N.$
For any fixed $x$, for a pair of time slots 2j1,2j2j-1,2j, the sequence $f_{t}$ and $g_{t}$’s cost is
$\sum_{t=2j-1}^{2j}f_{t}(x)=1.2+0.8x,\qquad\sum_{t=2j-1}^{2j}g_{t}(x)=0.8+1.2x.$
Thus, $C_{\mathsf{OPT}}$ (4) is
$C_{\mathsf{OPT}}=\min_{x\in\mathcal{X}}\max\left\{\sum_{j=1}^{N}\sum_{t=2j-1}^{2j}1.2+0.8x,\sum_{j=1}^{N}\sum_{t=2j-1}^{2j}0.8+1.2x\right\}=1.2N\quad(\text{at }x^{\star}=0).$
Hence,
$\frac{W_{L}^{T}}{C_{\mathsf{OPT}}}=\frac{1.8N}{1.2N}=1.5,$
as claimed. Therefore, even if an online algorithm $\mathcal{A}$ achieves sub-linear regret with respect to the benchmark $W_{L}^{T}$(10), it does not necessarily have sub-linear regret for our benchmark $C_{\mathsf{OPT}}$ (4).
Appendix E Experts Problem (Gains Version)
We are given $N$ experts. The interaction proceeds for $T$ rounds as follows:
  • On each round $t=1,2,\dots,T$:
    1. 1.
      Each expert $i\in\{1,\dots,N\}$ receives a gain $g_{t,i}\in[0,1]$.
    2. 2.
      The learner chooses a probability distribution
      $p_{t}=(p_{t,1},\dots,p_{t,N}),$
      where $p_{t,i}\geq 0$ and $\sum_{i=1}^{N}p_{t,i}=1$.
    3. 3.
      The learner’s gain is the expected value
      $G_{t}=\sum_{i=1}^{N}p_{t,i}\,g_{t,i}.$
The learner’s goal is to compete with the best fixed expert in hindsight by minimizing the regret
$R_{T}=\max_{i}\sum_{t=1}^{T}g_{t,i}\;-\;\sum_{t=1}^{T}G_{t}.$
Hedge Algorithm (Gains Version)
Algorithm 3 Hedge (Gains Version)
1:Number of experts NN, learning rate $\eta>0$
2:Initialize weights: $w_{1,i}\leftarrow 1$ for all $i\in\{1,\dots,N\}$
3:for $t=1,2,\dots,T$ do
4:  Form probabilities:
$p_{t,i}\leftarrow\frac{w_{t,i}}{\sum_{j=1}^{N}w_{t,j}}\quad\forall i$
5:  Play according to distribution $p_{t}=(p_{t,1},\dots,p_{t,N})$
6:  Observe gains $g_{t,i}\in[0,1]$ for each expert $i$
7:  Update weights:
$w_{t+1,i}\leftarrow w_{t,i}\,e^{\eta g_{t,i}}\quad\forall i$
8:end for
Guarantee
Theorem 9.
[7] With $\eta=\sqrt{\tfrac{2\ln N}{T}}$, Hedge achieves
$\max_{i}\sum_{t=1}^{T}g_{t,i}-\sum_{t=1}^{T}G_{t}\;\;\leq\;\;\sqrt{2T\ln N}.$
Appendix F Bounding $R_{3}$ with the adversarial input
In this section, we show the best bound one can obtain for $R_{3}$ on a sample path basis for Algorithm 1 with the adversarial input is $O(\eta_{\theta}T^{2})$.
Theorem 10.
Under Assumption 1, i.e., $\|\lambda_{t}(x)\|_{\infty}\leq B$, $R_{3}\leq 2B^{2}T\sum_{t=1}^{T}\eta_{\theta,t}$.
Recall the definitions $\lambda_{t}(x)=(f_{t}(x),g_{t}(x))$,
$R_{3}\;=\;\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\langle\theta_{t},\lambda_{t}(x)\rangle-C_{\mathsf{OPT}},\qquad\text{and}\qquad C_{\mathsf{OPT}}=\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\sum_{t=1}^{T}\langle\theta,\lambda_{t}(x)\rangle.$
Thus, an almost natural way (or really the only way for Algorithm 1 that combines OGD+Hedge) to bound $R_{3}$ is via exploiting the stability $\sum_{t}\|\theta_{t+1}-\theta_{t}\|_{1}$ of the Hedge algorithm. Following this path, we next derive Lemma 11, Lemma 12 and Lemma 13 which when combined imply Theorem 10.
However, it is not sufficient since by substituting $\eta_{\theta,t}=\frac{1}{\sqrt{T}}$ for all $t$, the usual Hedge step size, we get $R_{3}=\Omega(T^{3/2})$.
Lemma 11.
Define the time-averaged vector $\bar{\theta}:=\frac{1}{T}\sum_{t=1}^{T}\theta_{t}\in\Delta_{2}.$ Then,
$R_{3}\leq B\sum_{t=1}^{T}\|\theta_{t}-\bar{\theta}\|_{1}.$
Proof.
Since $\bar{\theta}$ is a particular (constant) element of $\Delta_{2}$, for every $x\in\mathcal{X}$
$\sum_{t=1}^{T}\langle\bar{\theta},\lambda_{t}(x)\rangle\leq\max_{\theta\in\Delta_{2}}\sum_{t=1}^{T}\langle\theta,\lambda_{t}(x)\rangle,$
and therefore taking the minimum over $x$ yields
$\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\langle\bar{\theta},\lambda_{t}(x)\rangle\leq C_{\mathsf{OPT}}.$
Thus,
$R_{3}$ $=\min_{x}\sum_{t=1}^{T}\langle\theta_{t},\lambda_{t}(x)\rangle-C_{\mathsf{OPT}}$
$\leq\min_{x}\sum_{t=1}^{T}\langle\theta_{t},\lambda_{t}(x)\rangle-\min_{x}\sum_{t=1}^{T}\langle\bar{\theta},\lambda_{t}(x)\rangle$
$\leq\max_{x}\sum_{t=1}^{T}\big\langle\theta_{t}-\bar{\theta},\lambda_{t}(x)\big\rangle,$
where the last inequality follows as below.
Let $x^{\star}$ satisfy $u(x^{\star})=\min_{x}u(x)$. Then
$\min_{x}u(x)-\min_{x}v(x)=u(x^{\star})-\min_{x}v(x)\leq u(x^{\star})-v(x^{\star})\leq\max_{x}(u(x)-v(x)),$
which proves the claim.
This proves
$R_{3}\leq\max_{x\in\mathcal{X}}\sum_{t=1}^{T}\langle\theta_{t}-\bar{\theta},\lambda_{t}(x)\rangle.$ (29)
Since we are in the bounded setting, for all rounds $t$ and all $x\in\mathcal{X}$ we have $\|\lambda_{t}(x)\|_{\infty}\leq B$. Thus, appyling the Hölder’s inequality (using duality of $||.||_{1}$ and $||.||_{\infty}$ norms) on (29),
$\sum_{t=1}^{T}\langle\theta_{t}-\bar{\theta},\lambda_{t}(x)\rangle\leq\sum_{t=1}^{T}\|\theta_{t}-\bar{\theta}\|_{1}\,\|\lambda_{t}(x)\|_{\infty}\leq B\sum_{t=1}^{T}\|\theta_{t}-\bar{\theta}\|_{1}.$
Thus,
$R_{3}\leq B\sum_{t=1}^{T}\|\theta_{t}-\bar{\theta}\|_{1}.$
This is the desired inequality. ∎
Next, we state an elementary inequality without proof.
Lemma 12 (Deviation from average via consecutive differences).
Let $\theta_{1},\dots,\theta_{T}\in\mathbb{R}^{d}$ be vectors, and define the average
$\bar{\theta}=\frac{1}{T}\sum_{t=1}^{T}\theta_{t}.$
Then
$\sum_{t=1}^{T}\|\theta_{t}-\bar{\theta}\|_{1}\;\leq\;2T\sum_{t=1}^{T-1}\|\theta_{t+1}-\theta_{t}\|_{1}.$
Lemma 13 (One-step TV bound for Hedge).
Let $\{\theta_{t}\}$ be the Hedge distributions updated by
$\theta_{t+1,i}\;=\;\frac{\theta_{t,i}\,e^{\eta_{\theta,t}\lambda_{t,i}}}{\sum_{j}\theta_{t,j}\,e^{\eta_{\theta,t}\lambda_{t,j}}},\qquad i=1,\dots,n,$
with learning rate $\eta_{\theta,t}>0$. Assume the losses satisfy $\lambda_{t,i}\in[0,B]$ for all $i$. Then
$\|\theta_{t+1}-\theta_{t}\|_{1}\leq\eta_{\theta,t}B.$
Proof.
Let
$Z_{t}:=\sum_{j}\theta_{t,j}e^{\eta_{\theta,t}\lambda_{t,j}}\;=\;\mathbb{E}_{i\sim\theta_{t}}\!\big[e^{\eta_{\theta,t}\lambda_{t,i}}\big].$
By the definition of Kullback–Leibler divergence,
$D_{\mathrm{KL}}(\theta_{t}\|\theta_{t+1})$ $=\sum_{i}\theta_{t,i}\ln\frac{\theta_{t,i}}{\theta_{t+1,i}}=\sum_{i}\theta_{t,i}\ln\frac{\theta_{t,i}}{\theta_{t,i}e^{\eta_{\theta,t}\lambda_{t,i}}/Z_{t}}$
$=\sum_{i}\theta_{t,i}(\ln Z_{t}-\eta_{\theta,t}\lambda_{t,i})=\ln Z_{t}-\eta_{\theta,t}\,\mathbb{E}_{i\sim\theta_{t}}[\lambda_{t,i}].$
Let $\mu:=\mathbb{E}_{i\sim\theta_{t}}[\lambda_{t,i}]$. Then $\ln Z_{t}=\ln\mathbb{E}_{\theta_{t}}\big[e^{\eta_{\theta,t}\lambda_{t,i}}\big]=\eta_{\theta,t}\mu+\ln\mathbb{E}_{\theta_{t}}\big[e^{\eta_{\theta,t}(\lambda_{t,i}-\mu)}\big]$, so (1) is equivalently
$D_{\mathrm{KL}}(\theta_{t}\|\theta_{t+1})=\ln\mathbb{E}_{\theta_{t}}\big[e^{\eta_{\theta,t}(\lambda_{t,i}-\mu)}\big].$ (30)
Thus to upper bound the one-step KL it suffices to bound the centered MGF $\ln\mathbb{E}[e^{\eta_{\theta,t}(\lambda_{t,i}-\mu)}]$.
Proposition 14 (Hoeffding’s MGF bound).
For a random variable $X$ that satisfies $X\in[a,b]$ almost surely and $\mathbb{E}[X]=\mu$, then for any $s\in\mathbb{R}$
$\ln\mathbb{E}[e^{s(X-\mu)}]\leq\frac{s^{2}(b-a)^{2}}{8}.$ (31)
Since $\lambda_{t,i}\in[0,B]$, using (31) in (30) with $X=\lambda_{t,i}$, [a,b]=[B,B][a,b]=[-B,B], $s=\eta_{\theta,t}$, and mean $\mu$, we get
$D_{\mathrm{KL}}(\theta_{t}\|\theta_{t+1})\leq\frac{\eta_{\theta,t}^{2}4B^{2}}{8}.$ (32)
Proposition 15 (Pinsker’s inequality [12]).
For any two probability vectors $p,q$,
$\|p-q\|_{1}\leq\sqrt{2\,D_{\mathrm{KL}}(p\|q)}.$ (33)
Since both $\theta_{t}$ and $\theta_{t+1}$ are distributions, applying (33) on (32), we get
$\|\theta_{t+1}-\theta_{t}\|_{1}\leq\sqrt{2\cdot\frac{\eta_{\theta,t}^{2}B^{2}}{2}}=\eta_{\theta,t}B,$
the claimed bound. ∎

### v3

@@
 Abstract
-In online convex optimization (OCO), a single loss function sequence is revealed over a time horizon of $T$, and an online algorithm has to choose its action at time $t$, before the loss function at time $t$ is revealed.
-The goal of the online algorithm is to incur minimal penalty (called regret) compared to a static optimal action made by an optimal offline algorithm knowing all functions of the sequence in advance.
-In this paper, we broaden the horizon of OCO, and consider multi-objective OCO, where there are $K$ distinct loss function sequences, and an algorithm has to choose its action at time $t$, before the $K$ loss functions at time $t$ are revealed.
+In this paper, we broaden the horizon of online convex optimization (OCO), and consider multi-objective OCO, where there are $K$ distinct loss function sequences, and an algorithm has to choose its action at time $t$, before the $K$ loss functions at time $t$ are revealed.
 To capture the tradeoff between tracking the $K$ different sequences, we consider the min-max regret, where the benchmark (optimal offline algorithm) takes a static action across all time slots that minimizes the maximum of the total loss (summed across time slots) incurred by each of the $K$ sequences.
 An online algorithm is allowed to change its action across time slots, and its min-max regret is defined as the difference between its min-max cost and that of the benchmark.
-The min-max regret is a stringent performance measure and an algorithm with small regret needs to ‘track’ all loss function sequences closely at all times.
-We consider this min-max regret in the i.i.d.
-input setting where all loss functions are i.i.d.
-generated from an unknown distribution.
-For the i.i.d.
-model we propose a simple algorithm that combines the well-known Hedge and online gradient descent (OGD) and show via a remarkably simple proof that its expected min-max regret is $O(\sqrt{T\log K})$.
+The min-max regret is a stringent performance measure and an algorithm with small regret needs to ‘track’ all loss functions simultaneously.
+We first show that with adversarial input, min-max regret scales linearly with the time horizon $T$ for any online algorithm.
+Consequently, we consider a stochastic i.i.d.
+input model where all loss functions are i.i.d.
+generated from an unknown joint distribution and propose a simple algorithm that combines the well-known Hedge and online gradient descent (OGD) and show via a remarkably simple proof that its expected min-max regret is $O(\sqrt{T\log(TK)})$.
+Analogous results are also derived for Martingale difference and Markov input models.
+The binary prediction problem with multiple binary sequences is also considered where an algorithm has to predict a binary action and the loss function is the maximum of the Hamming distance between the predicted actions and the binary sequences.
+For this problem, we again show that under adversarial input the min-max regret grows linearly with the time horizon $T$ for any online algorithm, while under the stochastic i.i.d.
+input model we propose an algorithm that achieves $O(1)$ min-max regret.
 1 Introduction
@@
 (2012) etc., where at time $t$, an online algorithm $\mathcal{A}$ has to choose an action $x_{t}\in\mathcal{X}$, after which the adversary reveals the convex objective/loss function $f_{t}$ on which $x_{t}$ is evaluated.
-As a benchmark, the offline optimal algorithm is considered, that knows the entire set of $f_{t}$, $t=1,\dots,T$, however, can only choose a single action across all times.
-To quantify the performance of online algorithms, the metric of static regret $R^{\text{s}ingle}_{T}$ is studied, that is defined as follows, where $x_{t}$’s are the actions taken by $\mathcal{A}$,
-$R^{\text{s}ingle}_{\mathcal{A}}(T)=\sup_{f_{t}}\left\{\sum_{t=1}^{T}f_{t}(x_{t})-\min_{x}\sum_{t=1}^{T}f_{t}(x)\right\}.$ (1)
-In a long line of work, comprehensively summarized in Hazan (2019), tight guarantees (both upper and lower bounds) have been obtained for the OCO problem, and $R^{\text{s}ingle}(T)=\inf_{\mathcal{A}}R^{\text{s}ingle}_{\mathcal{A}}(T)=\Theta(\sqrt{T})$.
+As a benchmark, the optimal offline algorithm is considered, that knows the entire set of $f_{t}$, $t=1,\dots,T$, however, can only choose a single action across all times.
+To quantify the performance of online algorithms with OCO, the metric of static regret $R^{\text{s}ingle}_{\mathcal{A}}(T)=\sup_{f_{t}}\left\{\sum_{t=1}^{T}f_{t}(x_{t})-\min_{x}\sum_{t=1}^{T}f_{t}(x)\right\}$ is studied, where $x_{t}$’s are the actions taken by $\mathcal{A}$.
+In a long line of work, comprehensively summarized in Hazan (2019), tight guarantees (both upper and lower bounds) have been obtained for the OCO problem, where optimal $R^{\text{s}ingle}_{\mathcal{A}}(T)=\Theta(\sqrt{T})$.
 In particular, online gradient descent (OGD), mirror descent, and follow-the-regularized-leader (FTRL) Hazan (2019) all achieve the optimal static regret performance for OCO.
@@
 We provide more details of this in the Related Work section.
-Thus, in this paper, we consider the robust alternative in the multi-objective case: the $\it min-max$} OCO, where the objective of an algorithm is to minimize the maximum of the total loss (summed across time slots) incurred by each of the multiple objective functions.
+Thus, in this paper, we consider the robust alternative in the multi-objective case: the min-max OCO, where the objective of an algorithm is to minimize the maximum of the total loss (summed across time slots) incurred by each of the multiple objective functions.
 A compelling example for considering the min-max approach is fair machine learning, where performance is typically evaluated across multiple demographic groups, which naturally induces a multi-objective optimization problem.
-While Pareto efficiency identifies trade-offs between groups, it may still permit highly unequal solutions, as a classifier can be Pareto optimal even if it performs very poorly on a disadvantaged group.
-To address this limitation, fairness is often modeled via a min–max objective: maximize the minimum accuracy (or minimize the maximum error) across groups Chen et al.
+While Pareto efficiency identifies trade-offs between groups, it may still permit highly unequal solutions.
+min-max objective effectively addresses this: it maximizes the minimum accuracy (or minimize the maximum error) across groups Chen et al.
 (2020); Chouldechova and Roth (2020)).
-The min-max approach is also well suited for modern networked systems—such as data centers, wireless edge networks, and shared cloud infrastructures—where online decisions must balance performance across multiple service classes, users, or flows whose objectives (throughput, latency, energy) often conflict.
-The min–max regret criterion provides a principled way to ensure worst-case fairness by minimizing the largest cumulative loss among competing entities, which is crucial in fair resource allocation Sinha et al.
-(2023), fair allocation in clouds Hao et al.
-(2016), fair assignment of reusable resources Liu and Hajiesmaili (2025) congestion-controlled transport Kouvelis and Yu (1997); Bertsimas et al.
-(2011), and dynamic resource provisioning in data centers Urgaonkar et al.
-(2010).
-Unlike average-regret formulations that can conceal systematic bias, minimizing min–max regret explicitly guards against persistent under-allocation to any group or flow—an increasingly important property in multi-tenant and QoS-sensitive networks.
+See Appendix A for more details on applications of min-max OCO problem.
 The formal min-max OCO problem that we consider is as follows.
-We consider that there are two convex loss function sequences $f_{t},g_{t},$ that are revealed after $\mathcal{A}$ chooses action $x_{t}$ at time $t$ for $t=1,\dots,T$.
+Let there be two convex loss function sequences $f_{t},g_{t},t=1,\dots,T$ that are revealed after $\mathcal{A}$ chooses action $x_{t}$ at time $t$.
 The min-max cost of $\mathcal{A}$ is
 $C_{\mathcal{A}}=\max\left\{\sum_{t=1}^{T}f_{t}(x_{t}),\sum_{t=1}^{T}g_{t}(x_{t})\right\}.$ (1)
-Two loss function sequences are considered for simplicity of exposition since they capture the basic difficulty.
+Two loss function sequences are considered for simplicity of exposition since they already capture the basic difficulty.
 Extension to $K>2$ sequences is straightforward and detailed in Remark 6.
-As a benchmark, we consider the static offline optimal algorithm $\mathsf{OPT}$ that knows $f_{t},g_{t}$ for $t=1,\dots,T$ at time $t=1$ itself but can only choose a single action $x^{\star}$ across all times such that
-$x^{\star}=\arg\min_{x\in\mathcal{X}}\max\left\{\sum_{t=1}^{T}f_{t}(x),\sum_{t=1}^{T}g_{t}(x)\right\},$ (3)
-and the optimal cost is
-$C_{\mathsf{OPT}}=\max\left\{\sum_{t=1}^{T}f_{t}(x^{\star}),\sum_{t=1}^{T}g_{t}(x^{\star})\right\}.$ (4)
-Consequently, we define the min-max static regret of $\mathcal{A}$ with actions $x_{t}$ as
+As a benchmark, we consider the static offline optimal algorithm $\mathsf{OPT}$ that knows $f_{t},g_{t}$ for $t=1,\dots,T$ at time $t=1$ itself but can only choose a single action $x^{\star}$ across all times and the optimal cost is
+$C_{\mathsf{OPT}}=\max\left\{\sum_{t=1}^{T}f_{t}(x^{\star}),\sum_{t=1}^{T}g_{t}(x^{\star})\right\},$ (2)
+where $x^{\star}=\arg\min_{x\in\mathcal{X}}\max\left\{\sum_{t=1}^{T}f_{t}(x),\sum_{t=1}^{T}g_{t}(x)\right\}.$ Consequently, we define the min-max static regret of $\mathcal{A}$ with actions $x_{t}$ as
 $R_{\mathcal{A}}(T)=\sup_{f_{t},g_{t}}\{C_{\mathcal{A}}-C_{\mathsf{OPT}}\}.$ (3)
 We refer to (3) as the min-max OCO problem.
-Remark 1 (On the regret definition).
-The chosen benchmark in (5) is deliberately stringent.
-If we considered a weaker benchmark, e.g., minimizing $\min_{x\in X}\sum_{t=1}^{T}\max\{f_{t}(x),g_{t}(x)\}$), then applying standard OGD algorithm on the surrogate function $h_{t}(x)=\max\{f_{t}(x),g_{t}(x)\}$ would yield regret $O(\sqrt{T})$ independent of the number of objectives, underestimating the intrinsic hardness of balancing multiple sequences.
-In particular:
-  • Our definition forces the algorithm to track both sequences globally and simultaneously.
-This means the online algorithm cannot focus on one objective at the expense of the other, nor can it switch actions adaptively across time as the per-slot benchmark allows.
-  • Our benchmark ensures robustness across objectives: the comparator $x^{\star}$ is the best fixed action that minimizes the worst cumulative cost.
-  • This definition aligns with fairness and robustness considerations (e.g., in machine learning across demographic groups), where good performance must be guaranteed for all objectives, not merely on average or in expectation.
-Thus, while the regret metric in (5) is harder to minimize, it is an appropriate notion when robustness across multiple objectives is essential.
-1.1 Related Work
-Because of its widespread applications, multi-objective OCO has been an object of immense interest, however, with limited success.
-For example, Mahdavi et al.
-(2013) wanted to study the multi-objective OCO and stated the difficulty as “A fundamental difference between single- and multi-objective optimization is that for the latter it is not obvious how to evaluate the optimization quality”.
-Further, it said “Since it is impossible to simultaneously minimize multiple loss functions and in order to avoid complications caused by handling more than one objective, we choose one function as the objective and try to bound other objectives by appropriate thresholds.
-This work gave rise to the now popular problem of constrained online convex optimization (COCO), where the problem is transformed into a constrained optimization problem with one objective and one or more constraint functions, that is described next.
-1.1.1 COCO
-In COCO, on every round $t,$ an online algorithm first chooses an admissible action $x_{t}\in\mathcal{X}\subset{\mathbb{R}}^{d},$ and then the adversary chooses a convex loss/cost function $f_{t}:\mathcal{X}\to\mathbb{R}$ and a constraint function of the form $g_{t}(x)\leq 0,$ where $g_{t}:\mathcal{X}\to\mathbb{R}$ is a convex function.
-Let $\mathcal{X}^{\star}$ be the feasible set consisting of all admissible actions that satisfy all constraints $g_{t}(x)\leq 0,t\in[T]$.
-A standard assumption is made that $\mathcal{X}^{\star}$ is not empty (called the feasibility assumption).
-Since $g_{t}$’s are revealed after the action $x_{t}$ is chosen, an online algorithm $\mathcal{A}$ need not take feasible actions on each round, and in addition to the static regret defined as
-$\textrm{R}^{\text{COCO}}_{\mathcal{A}}(T)\equiv\sup_{\{f_{t}\}_{t=1}^{T}}\left\{\sum_{t=1}^{T}f_{t}(x_{t})-\inf_{x\in\mathcal{X}^{\star}}\sum_{t=1}^{T}f_{t}(x^{\star})\right\},$ (6)
-an additional metric of interest is the total cumulative constraint violation (CCV) defined as
-$\textrm{CCV}_{\mathcal{A}}(T)\equiv\sum_{t=1}^{T}\max(g_{t}(x_{t}),0).$
-The goal with COCO is to design an online algorithm to simultaneously achieve a small regret (6) with respect to any admissible benchmark $x^{\star}\in\mathcal{X}^{\star}$ and a small CCV.
-1.1.2 Prior Work on COCO
-(A) i.i.d.
-loss and constraints: COCO with independent and identically distributed (i.i.d.) loss functions $f_{t}$’s and constraint functions $g_{t}$’s was studied in Mahdavi et al.
-(2013) while its extension to stationary and ergodic processes can be found in Uziel and El-Yaniv (2017), where both derived an algorithm with regret $O(\sqrt{T})$ and CCV $O(T^{3/4})$.
-(B) Time-invariant constraints: COCO with time-invariant constraints, i.e., $g_{t}=g,\forall\ t$ (Yuan and Lamperski, 2018; Jenatton et al., 2016; Mahdavi et al., 2012; Yi et al., 2021) has been considered extensively, where functions $g$ are assumed to be known to the algorithm a priori.
-The best known results in this case give static regret of $O(T^{\max\{c,1-c\}})$ and CCV of $O(T^{(1-c)/2})$ for any $c\in(0,1)$.
-The algorithm is allowed to take actions that are infeasible at any time to avoid the costly projection step of the vanilla projected OGD algorithm and the main objective was to design an efficient algorithm with a small regret and CCV while avoiding the explicit projection step.
-(C) Time-varying constraints: The more difficult question is solving COCO problem when the constraint functions, i.e., $g_{t}$’s, change arbitrarily with time $t$.
-One popular algorithm for solving COCO considered a Lagrangian function optimization that is updated using the primal and dual variables (Yu et al., 2017; Sun et al., 2017; Yi et al., 2023).
-Alternatively, (Neely and Yu, 2017) and (Liakopoulos et al., 2019) used the drift-plus-penalty (DPP) framework (Neely, 2010) to solve the COCO, but which needed additional assumption, e.g.
-the Slater’s condition in (Neely and Yu, 2017) and with weaker form of the feasibility assumption in (Neely and Yu, 2017).
-(Guo et al., 2022) obtained bounds similar to (Neely and Yu, 2017) but without assuming Slater’s condition.
-Very recently, the state of the art guarantees on simultaneous bounds on regret $O(\sqrt{T})$ and CCV $O(\sqrt{T}\log T)$ for COCO were derived in (Sinha and Vaze, 2024) with a very simple algorithm that combines the loss function at time $t$ and the CCV accrued till time $t$ in a single loss function, and then executes the online gradient descent (OGD) algorithm on the single loss function with an adaptive step-size.
-Most recently (Sinha and Vaze, 2025) showed lower CCV is possible with increased regret compared to (Sinha and Vaze, 2024) while (Vaze and Sinha, 2025) obtained instance dependent CCV bounds which can be $O(1)$ for specific cases while achieving regret of $O(\sqrt{T})$, respectively.
-The COCO problem has been considered in the dynamic setting as well (Chen and Giannakis, 2018; Cao and Liu, 2018; Vaze, 2022; Liu et al., 2022) where the benchmark $x^{\star}$ in (6) is replaced by $x_{t}^{\star}$ ($x_{t}^{\star}=\arg\min_{x}f_{t}(x)$) that is also allowed to change its actions over time.
 Remark 1.
-Even though the min-max OCO and the COCO problem appear related, but unfortunately results for one do not give results for the other in either direction.
+The chosen benchmark in (2) is deliberately stringent and forces the algorithm to track both sequences globally and simultaneously compared to $\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\max\{f_{t}(x),g_{t}(x)\}$, for which the OGD algorithm applied on the surrogate function $h_{t}(x)=\max\{f_{t}(x),g_{t}(x)\}$ would yield regret $O(\sqrt{T})$ independent of the number of distinct loss function sequences.
+2 Related Work
+Early attempts at multi-objective OCO highlighted the difficulty of defining an appropriate performance metric Mahdavi et al.
+(2013).
+A common resolution, introduced in Mahdavi et al.
+(2013), is constrained online convex optimization (COCO), where one objective is optimized while the remaining objectives are enforced as constraints.
+In COCO, an algorithm chooses $x_{t}\in\mathcal{X}$ before observing a convex loss $f_{t}$ and convex constraint $g_{t}$, and performance is measured by both static regret with respect to a feasible comparator and cumulative constraint violation.
+COCO has been extensively studied under i.i.d.
+inputs Mahdavi et al.
+(2013); Uziel and El-Yaniv (2017), time-invariant constraints Jenatton et al.
+(2016); Mahdavi et al.
+(2012); Yuan and Lamperski (2018); Yi et al.
+(2021), and fully time-varying constraints using primal–dual or drift-plus-penalty methods Neely and Yu (2017); Guo et al.
+(2022); Yu et al.
+(2017), with the best possible guarantee being regret of $O(\sqrt{T})$ and constraint violation of $O(\sqrt{T}\log T)$ with simple algorithms Sinha and Vaze (2024), with further refinements yielding instance-dependent or improved violation bounds Vaze and Sinha (2025); Sinha and Vaze (2025).
+Dynamic versions of COCO, where the benchmark varies over time, have also been considered Chen and Giannakis (2018); Cao and Liu (2018); Vaze (2022); Liu et al.
+(2022).
+Despite superficial similarities, COCO and min-max OCO are fundamentally different: COCO enforces feasibility over time, whereas min-max OCO compares against a static action that minimizes the worst total loss across objectives.
+Consequently, results for one setting do not directly imply results for the other.
+A more detailed discussion on COCO can be found in Appendix B.
 2.1 Global Cost Functions in Experts Setting
@@
 (2009), where there are $K$ experts, with expert $i$’s loss at time $t$, being $\ell_{t}(i)$.
-If expert $i$ is selected by an algorithm $\mathcal{A}$ with probability $p_{t}(i)$ at time $t$, then the overall expected loss for expert $i$ is $L_{i}=\sum_{t=1}^{T}p_{t}(i)\ell_{t}(i)$ where $T$ is the time horizon, and the cost of $\mathcal{A}$ is the global cost Even-Dar et al.
-(2009)
-$C_{\mathcal{A}}=\max_{i=1,\dots,K}\{L_{i}\}.$ (7)
-The regret of $\mathcal{A}$ is defined as usual with respect to the static optimal offline algorithm that is required to choose a fixed distribution across all time slots while knowing all losses in advance.
-The name global refers to the fact that the objective function (7) captures the end-to-end cost function rather than a summation of the instantaneous losses.
-For this formulation, which we refer to as GlobalExpertsProblem, Even-Dar et al.
-(2009) proposed an algorithm called MULTI and showed its regret to be $O(\sqrt{T}\log(K))$.
-GlobalExpertsProblem differs with our min-max OCO formulation, in two fundamental ways, i) the objective inside the $\max$ is a linear function of the actions $p_{t}(i)$, and ii) the action $p_{t}(i)$ taken at time $t$ affects each expert’s loss but only in an average sense since $\sum_{i=1}^{K}p_{t}(i)=1$ while with min-max OCO the action $x_{t}$ chosen at time $t$ affects both $\sum_{t=1}^{T}f_{t}(x_{t})$ and $\sum_{t=1}^{T}g_{t}(x_{t})$ in full.
-Even though this appears to only be a small difference, it changes everything.
-As a consequence the result on GlobalExpertsProblem does not apply for min-max OCO even when $f_{t},g_{t}$ are linear.
-Extensions of Even-Dar et al.
-(2009) with additive global functions can be found in Azar et al.
-(2014); Kesselheim and Singla (2020) while for non-additive ones in Rakhlin et al.
-(2011).
+If expert $i$ is selected by an algorithm $\mathcal{A}$ with probability $p_{t}(i)$ at time $t$, then the overall expected loss for expert $i$ is $L_{i}=\sum_{t=1}^{T}p_{t}(i)\ell_{t}(i)$, and the cost of $\mathcal{A}$ is the global cost $C_{\mathcal{A}}=\max_{i=1,\dots,K}\{L_{i}\}$.
+While Even-Dar et al.
+(2009) achieved $O(\sqrt{T}\log K)$ regret using the MULTI algorithm, their setting fundamentally differs from ours: their objective is linear in the actions $p_{t}$, and the probability mass $p_{t}(i)$ is split among experts (summing to 1).
+In contrast, with min-max OCO actions affect all objective functions simultaneously and in full.
+Consequently, results from Even-Dar et al.
+(2009) and its extensions Azar et al.
+(2014); Kesselheim and Singla (2020); Rakhlin et al.
+(2011) do not apply to our formulation, even for linear functions.
 2.2 Pareto Optimal Multi-Objective OCO
@@
 (2023) made a valid attempt at modelling the general multi-objective OCO, the weakness of the metric limited its applicability.
-1.4 Per-slot $\it{\it min-max}$} multi-objective optimization
+2.3 Per-slot min-max multi-objective optimization
 Compared to our ‘global’ min-max objective (2), in Lee et al.
-(2022), a per-slot $\it min-max$} multi-objective OCO was studied, where in particular, at each round $t=1,\dots,T$: i) an online algorithm chooses $x_{t}\in\mathcal{X}$ (convex compact set), ii) the adversary chooses $y_{t}\in\mathcal{Y}$ (convex compact set), and iii) the online algorithm suffers a vector loss:
-$\ell_{t}(x_{t},y_{t})=\big(\ell_{t,1}(x_{t},y_{t}),\dots,\ell_{t,d}(x_{t},y_{t})\big)\in[-1,1]^{d}.$
-The goal of an online algorithm is to solve $\min_{\{x_{t}\}}\max_{j\in[d]}\sum_{t=1}^{T}\ell_{t,j}(x_{t},y_{t})$, knowing $\ell_{t}(.,.)$, but not $y_{t}$, and its performance is compared against a per-round local minimax value
-$w_{L}^{t}:=\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\max_{j\in[d]}\ell_{t,j}(x,y),$
-where the cumulative benchmark is:
+(2022), a per-slot min-max multi-objective OCO was studied, where at each round $t=1,\dots,T$: i) an online algorithm chooses $x_{t}\in\mathcal{X}$, ii) the adversary chooses $y_{t}\in\mathcal{Y}$, and iii) the online algorithm suffers a vector loss $\ell_{t}(x_{t},y_{t})=\big(\ell_{t,1}(x_{t},y_{t}),\dots,\ell_{t,d}(x_{t},y_{t})\big)\in[-1,1]^{d}.$ The goal of an online algorithm is to solve $\min_{\{x_{t}\}}\max_{j\in[d]}\sum_{t=1}^{T}\ell_{t,j}(x_{t},y_{t})$, knowing $\ell_{t}(.,.)$, but not $y_{t}$, and its performance is compared against a per-round local minimax value $w_{L}^{t}:=\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\max_{j\in[d]}\ell_{t,j}(x,y),$ where the cumulative benchmark is:
 $W_{L}^{T}:=\sum_{t=1}^{T}w_{L}^{t},\quad\text{and regret}\quad\mathrm{R}^{d}_{\mathcal{A}}(T):=\max_{j\in[d]}\sum_{t=1}^{T}\ell_{t,j}(x_{t},y_{t})-W_{L}^{T}.$ (5)
-A Hedge-type algorithm Cesa-Bianchi and Lugosi (2006) was shown in Lee et al. (2022) to have $O(\sqrt{T})$ regret.
-Defining $\mathcal{Y}=\{e\}$ a singleton, and $d=2$, such that $\ell_{t,1}(x_{t},y_{t})=f_{t}(x_{t})$ and $\ell_{t,1}(x_{t},y_{t})=g_{t}(x_{t})$ for $y_{t}=e$, we recover our model where two convex functions $f_{t},g_{t}$ are revealed at time $t$, and for which the benchmark $W_{L}^{T}$ simplifies to
-$W_{L}^{T}=\sum_{t=1}^{T}\min_{x_{t}\in\mathcal{X}}\max\{f_{t}(x_{t}),g_{t}(x_{t})\}).$ (10)
-Thus, benchmark (10) is fundamentally different than our considered benchmark (4) since it is additive across time slots, and allows the benchmark to choose different actions across time slots.
-Even though (10) allows choosing different actions across time slots, surprisingly, we show in Appendix D that $W_{L}^{T}$ (10) can be $3/2$ times more than our benchmark $C_{\mathsf{OPT}}$ (4) and hence algorithms with $O(\sqrt{T})$ regret (9) do not yield a $O(\sqrt{T})$ min-max static regret (5).
-In summary, compared to COCO and GlobalExpertsProblem, our considered problem has a wider scope, while our considered metric (5) is more meaningful and stringent compared to (8) and (10).
-1.5 Applications
-Multi-objective OCO has also been studied from an applications perspective, e.g.
-Lyu et al.
-(2019) studied the problem of multi-period multi-objective online ride-matching problem, where the multiple objectives included platform revenue, assigning drivers with higher service quality, minimizing the distance to the customer etc.
-Lyu et al.
-(2019) also used the COCO-type reformulation of multi-objective OCO for deriving theoretical guarantees.
-Other applications of multi-objective OCO include fair resource allocation Sinha et al.
-(2023), fair allocation in clouds Hao et al.
-(2016), fair assignment of reusable resources Liu and Hajiesmaili (2025) congestion-controlled transport Kouvelis and Yu (1997); Bertsimas et al.
-(2011), and dynamic resource provisioning in data centers Urgaonkar et al.
-(2010), multi-task learning Sener and Koltun (2018), multiple target tracking Milan et al.
-(2017), portfolio selection Das and Banerjee (2011) and safe reinforcement learning Chow et al.
-(2018), where decision-makers must make sequential decisions under uncertainty, balancing objectives such as cost, delivery time, and reliability without full knowledge of future demand or disruptions Kouvelis and Yu (1997); Bertsimas et al.
-(2011), where the min-max criterion naturally makes sense.
+A Hedge-type algorithm Cesa-Bianchi and Lugosi (2006) was shown in Lee et al.
+(2022) to have $O(\sqrt{T})$ regret.
+Defining $\mathcal{Y}=\{e\}$ as a singleton, and $d=2$, such that $\ell_{t,1}(x_{t},y_{t})=f_{t}(x_{t})$ and $\ell_{t,2}(x_{t},y_{t})=g_{t}(x_{t})$ for $y_{t}=e$, we recover our model where two convex functions $f_{t},g_{t}$ are revealed at time $t$, and for which $W_{L}^{T}=\sum_{t=1}^{T}\min_{x_{t}\in\mathcal{X}}\max\{f_{t}(x_{t}),g_{t}(x_{t})\}).$ This benchmark is fundamentally different than our considered benchmark (2) since it is additive across time slots, and allows the benchmark to choose different actions across time slots.
+Even though this benchmark allows choosing different actions across time slots, surprisingly, we show in Appendix H that $W_{L}^{T}$ can be $3/2$ times more than our benchmark $C_{\mathsf{OPT}}$ (2) and hence algorithms with $O(\sqrt{T})$ regret (5) do not yield a $O(\sqrt{T})$ min-max static regret (3).
+2.4 $\alpha$-Fairness
+Finally, we review the most closely related problem of online $\alpha$-fairness studied by Si Salem et al.
+(2023).
+In this formulation with two users, and at each round $t\in[T]$ an online algorithm selects an action $x_{t}\in\mathcal{X}$.
+After $x_{t}$ is chosen, two nonnegative concave utility functions $f_{t},g_{t}:\mathcal{X}\to\mathbb{R}_{\geq 0}$ are revealed.
+Given a sequence of actions $\{x_{t}\}_{t=1}^{T}$, the time-averaged utilities are $\bar{f}(x_{1:T}):=\frac{1}{T}\sum_{t=1}^{T}f_{t}(x_{t})$, and $\bar{g}(x_{1:T}):=\frac{1}{T}\sum_{t=1}^{T}g_{t}(x_{t})$.
+Let $\bar{f}(x_{1:T})=\bar{f}(x)$ when $x_{t}=x$ for all $t$.
+Following Si Salem et al.
+(2023), performance is evaluated via the $\alpha$-fair utility
+$U_{\alpha}(\bar{f})+U_{\alpha}(\bar{g}),\qquad U_{\alpha}(z)=\begin{cases}\frac{z^{1-\alpha}}{1-\alpha},&\alpha\neq 1,\\ \log z,&\alpha=1,\end{cases}$ (6)
+which is a concave aggregation of the two user utilities.
+To compare with our min-max framework in this paper, define a cost function $\psi_{\alpha}:\mathbb{R}_{>0}^{2}\to\mathbb{R}$ where $\psi_{\alpha}(a,b):=-\,\Big(U_{\alpha}(a)+U_{\alpha}(b)\Big)$, which is convex in $(a,b)$ for all $\alpha\geq 0$.
+The offline benchmark of Si Salem et al.
+(2023) is then $C_{\mathsf{OPT}}(\alpha,T)=\min_{x\in\mathcal{X}}\psi_{\alpha}(\bar{f}(x),\bar{g}(x))$ and the cost of $\mathcal{A}$ is $C_{\mathcal{A}}(\alpha,T)=\psi_{\alpha}(\bar{f}(x_{1:T}),\bar{g}(x_{1:T}))$.
+In the limit $\alpha\to\infty$ costs $C_{\mathcal{A}}(\alpha,T)$ and $C_{\mathsf{OPT}}(\alpha,T)$ become equivalent to the min-max costs defined in (1) and (2), respectively.
+Thus, the online $\alpha$-fairness problem Si Salem et al.
+(2023) formally subsumes our min-max OCO problem as the limiting case $\alpha\to\infty$.
+Si Salem et al.
+(2023) show that with adversarial input, the regret $C_{\mathcal{A}}(\alpha,T)-C_{\mathsf{OPT}}(\alpha,T)=\Omega(T)$ for any online algorithm, even with fixed $\alpha$, ruling out sublinear regret in this model.
+Motivated by this impossibility, they consider a stochastic (smoothed) input model, under which they propose an online horizon-fair (OHF) policy and prove that its regret is $O(\alpha\sqrt{T})$.
+Notably, since the regime of interest for our min-max objective (1) corresponds to $\alpha\to\infty$, the stochastic regret guarantee of Si Salem et al.
+(2023) becomes vacuous.
+See Remark 11 for more details.
 2.5 Our Contributions
   1. 1.
-    Compared to the discussed prior work, which has primarily focused on bounding the ‘respective’ regret under worst-case or adversarial inputs, establishing bounds with the adversarial input on the min–max static regret (5) appears to be substantially more challenging.
-We highlight the difficulty for one instance in Section 3, where even when full functions $f_{t},g_{t}$ are revealed before choosing action $x_{t}$, the greedy algorithm still has a linear regret.
+    We begin by deriving a lower bound of $\Omega(T)$ on the min-max static regret (3) of any online algorithm with adversarial input.
+The input construction is inspired from Si Salem et al.
+(2023) and is modified to suit our specific problem.
   2. 2.
-    Thus, we consider the min-max static regret problem in the i.i.d.
-setting where both $f_{t},g_{t}$ are chosen independently across $t$ with identical (unknown) distribution $\mathcal{D}$.
-Note that all the discussed applications and use cases for multi-objective OCO (e.g.
-fair machine learning, fair resource allocation, multi-task learning, network scheduling, multi-objective online ride-matching problem etc.) are well motivated in the i.i.d.
-case as well.
-Moreover, analytically also, the problem remains challenging with the i.i.d.
+    Motivated by the impossibility of sublinear min-max static regret in the adversarial setting, we consider a stochastic i.i.d.
+model, [^2] We also consider the Martingale difference and Markov input models in Appendix N and O.
+where loss pairs $(f_{t},g_{t})$ are drawn from a fixed unknown distribution $\mathcal{D}$.
+This setting is standard for applications like fair resource allocation and machine learning and remains analytically challenging.
+    Under the i.i.d.
+model, the benchmark is ${\mathbb{E}}\{C_{\mathsf{OPT}}\}$ where $C_{\mathsf{OPT}}$ has been defined in (2).
+Directly analyzing the min-max static regret (3) with respect to ${\mathbb{E}}\{C_{\mathsf{OPT}}\}$ is technically challenging, since $C_{\mathsf{OPT}}$ depends on the entire realized sequence of losses.
+Instead, we introduce a deterministic surrogate benchmark $\bar{C}_{\mathsf{OPT}}$ obtained by averaging out the randomness in the losses:
+    $\bar{C}_{\mathsf{OPT}}=T\cdot\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x),\qquad\mu(x)={\mathbb{E}}\{[f_{t}(x),g_{t}(x)]\},$ (7)
+    where $\Delta_{K}$ is the $K$-dimensional simplex. Importantly, as shown in Appendix D,
+    ${\mathbb{E}}\!\left\{\,\big|C_{\mathsf{OPT}}-\bar{C}_{\mathsf{OPT}}\big|\,\right\}=O\!\left(\sqrt{T\log T}\right).$
+    Using $\bar{C}_{\mathsf{OPT}}$ as the benchmark, we define the expected min-max static regret of an online algorithm $\mathcal{A}$ as
+    $\bar{R}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}\left\{{\mathbb{E}}\{C_{\mathcal{A}}\}-\bar{C}_{\mathsf{OPT}}\right\}.$ (8)
+    For clarity, we refer to regret measured with respect to ${\mathbb{E}}\{C_{\mathsf{OPT}}\}$ as the true min-max static regret, and regret measured with respect to $\bar{C}_{\mathsf{OPT}}$ as the expected min-max static regret.
+    In the remainder of the paper, we bound $\bar{R}_{\mathcal{A}}(T)$.
+From the above concentration bound, the true min-max static regret satisfies ${\mathbb{E}}\{C_{\mathcal{A}}\}-{\mathbb{E}}\{C_{\mathsf{OPT}}\}\;\leq\;\bar{R}_{\mathcal{A}}(T)+O\!\left(\sqrt{T\log T}\right)$.
+[^4] With $K$ sequences, the additive term becomes $O\!\left(\sqrt{T\log(TK)}\right)$.
+    Even under the i.i.d.
+model, the lower bound on the true min–max static regret is $\Omega(\sqrt{T})$.
+This follows by taking $f_{t}\equiv g_{t}$ for all $t$, which reduces the problem to standard OCO, for which ${\mathbb{E}}\{R^{\text{single}}_{\mathcal{A}}(T)\}=\Omega(\sqrt{T})$ even under i.i.d.
+inputs (e.g., random linear losses; see Hazan (2019)).
+  3. 3.
+    Algorithm While considering expected min-max static regret (8), the structure of (7) reveals the core intuition for our approach: the optimal policy corresponds to a saddle point $(x^{\star},\theta^{\star})$ that solves ${\bar{C}}_{\mathsf{OPT}}$ in (7).
+Since the distribution $\mathcal{D}$ is unknown, we cannot compute this pair offline.
+Instead, we propose a natural online algorithm (Algorithm 1) that interlaces two classic strategies: i) Hedge to update the weights $\theta_{t}$ (solving the inner maximization over objectives in (7)), and ii) OGD to update the action $x_{t}$ (solving the outer minimization on the surrogate loss $\theta_{t}^{\top}[f_{t},g_{t}]$ in (7)).
+    While (7) resembles a zero-sum game Cardoso et al.
+(2019), the problem is fundamentally different: here, the learner controls both $x_{t}$ and $\theta_{t}$ to track an unknown distribution.
+The novelty of our approach lies not in the individual choice of Hedge or OGD, but in their principled combination which simplifies an otherwise intractable analysis.
+While other candidates like FTRL or algorithms for COCO Yu et al.
+(2017); Sinha and Vaze (2024) have an intuitive appeal for minimizing (8), there are obvious analytical roadblocks caused by the global $\max$ in the cost function (1) even with stochastic i.i.d.
 input.
-From a theoretical advancement point of view similar approach was followed for the COCO problem where the earliest results were derived for the i.i.d.
-input Mahdavi et al.
-(2013), which were then later generalized to the worst case input.
-    For the i.i.d. model, the expected cost of $\mathcal{A}$ is ${\bar{C}}_{\mathcal{A}}$ that is just the expected value of $C_{\mathcal{A}}$ in (2), while the $C_{\mathsf{OPT}}$ simplifies to
-    ${\bar{C}}_{\mathsf{OPT}}=T\cdot\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x)\quad\text{with}\quad(x^{\star},\theta^{\star})=\arg\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x),$ (11)
-    where $\mu(x)={\mathbb{E}}\{f_{t}(x),g_{t}(x)\}$, $\theta=(\theta_{1},\theta_{2})\in\Delta_{2}$ and $\Delta_{2}$ is the 2D2-D simplex. The metric is now the expected min-max static regret
-    ${\bar{R}}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}\{{\bar{C}}_{\mathcal{A}}-{\bar{C}}_{\mathsf{OPT}}\}.$ (12)
-    Algorithm: For the i.i.d.
-model, the basic idea behind our solution approach for minimizing the expected min-max static regret (12) is as follows.
-From (11), the real goal for an algorithm is to find the optimal pair $(x^{\star},\theta^{\star})$.
-Since an online algorithm $\mathcal{A}$ does not know $\mathcal{D}$ ahead of time, finding $(x^{\star},\theta^{\star})$ is not possible.
-So we propose a simple algorithm that produces pair $(x_{t},\theta_{t})$ at time $t$ in an online manner that uses Hedge algorithm [^2] We describe the basic Experts problem and Hedge algorithm in Appendix E together with its regret guarantee for completeness purposes.
-Cesa-Bianchi and Lugosi (2006) to update $\theta_{t}$ to solve the inner maximization problem in (11) given $x_{t}$, and OGD to obtain actions $x_{t}$ given $\theta_{t}$ to solve the outer minimization (OCO) problem in (11) applied to the surrogate loss function $\theta_{t}^{\top}[f_{t}(x),g_{t}(x)]=\theta_{1t}f_{t}(x)+\theta_{2t}g_{t}(x)$ which is a proxy for $\theta^{\top}\mu(x)$, in an attempt to keep $(x_{t},\theta_{t})$ to stay close to $(x^{\star},\theta^{\star})$, so that the regret remains small.
-    Note that (11) is not the usual online $\it min-max$} optimization problem Cardoso et al.
-(2019), where an online algorithm and adversary are playing against each other.
-Instead with (11) both $x$ and $\theta$ are algorithm’s choices depending on $\mathcal{D}$, that is under an adversarial control as defined in (12).
-The proposed alternating min-max algorithm is intuitive, however, the regret analysis is actually quite delicate, and heavily depends on our non-trivial decomposition of regret in (19) into three terms, which also motivates our algorithm.
-    It is worth noting that there are several other candidate algorithms (e.g.
-primal-dual algorithm for COCO Yu et al.
-(2017) and OGD with surrogate cost function for COCO Sinha and Vaze (2024)) that appear to solve (11), however, analyzing them is rather challenging because of the outside $\max$ in the cost of an online algorithm in (5).
-  3. 3.
-    Guarantee: Using simple and elegant analysis via an insightful regret decomposition, we show that the expected min-max static regret (12) of the proposed algorithm is $O(\sqrt{T})$.
-This is similar to the short and elegant analysis of OGD algorithm’s Hazan (2019) performance for OCO in the adversarial input setting, that has been very useful for further progress.
-Even though we considered just two sequences $f_{t},g_{t}$ for simplicity, the result generalizes directly if there are $K$ sequences $f^{k}_{t},k=1,\dots,K$ and the expected min-max static regret (5) is $O(\sqrt{T\log K})$.
-See Remark 7 for more details.
+Our framework resolves this difficulty through a modular regret decomposition and decouples the dynamics of Hedge and OGD that renders the proof both modular and elegant.
+    We prove that our algorithm achieves an expected min-max static regret (8) of $O(\sqrt{T})$ (or $O(\sqrt{T\log(K)})$ for $K$ objectives).
+Analogous results for Martingale difference and Markov input models are provided in Appendix N and O.
+    Recall that the $\alpha$-fair model of Si Salem et al.
+(2023) is equivalent to our min-max setting when $\alpha\rightarrow\infty$.
+Thus, importantly, compared to the $O(\alpha\sqrt{T})$ regret bound of Si Salem et al.
+(2023) we obtain an $\alpha$-independent regret bound.
   4. 4.
-    Improved guarantee for GlobalExpertsProblem Even-Dar et al.
-(2009) in the i.i.d.
-input: As a special case of our result, we derive an improved regret guarantee of $O(\sqrt{T\log K})$ for the GlobalExpertsProblem Even-Dar et al.
-(2009) with the i.i.d.
-input compared to $O(\sqrt{T}\log K)$ of Even-Dar et al.
-(2009) which was applicable in the worst case/adversarial input setting.
+    With Bandit Feedback: Exploiting our modular proof approach, we easily extend our results to the bandit setting, where only function evaluations at the most recent actions are available without any gradient feedback.
+We show that the expected min-max static regret (8) is $O(T^{3/4}\sqrt{\log K})$ and $O(\sqrt{T\log K})$ for one-point and two-point bandit feedback, respectively, matching the best known bounds on OCO in the respective bandit settings.
   5. 5.
-    With Bandit Feedback: We also extend our results to the bandit setting, where only function evaluations at the most recent actions are available without any gradient feedback.
-We show that the expected min-max static regret (12) is $O(T^{3/4}\sqrt{\log K})$ and $O(\sqrt{T\log K})$ for one-point and two-point bandit feedback, respectively, matching the best known bounds on OCO in the respective bandit settings.
-1.7 Open Questions
-  • We have only considered the i.i.d. input setting, and the major open question that remains is whether $o(T)$ min-max static regret (5) is achievable in the worst case input or not.
-  • Lower bound on the expected min-max regret: Theoretically, deriving a lower bound on the expected min-max static regret matching the derived guarantee of $O(\sqrt{T})$ appears challenging, however, simulations (see Fig.
-1) with $f_{t},g_{t}$ as random linear functions suggest that $\Theta(\sqrt{T})$ is the correct scaling for the expected min-max static regret.
+    We also consider the min-max generalization of the universal prediction of binary sequences Feder et al.
+(2002); Cover (1966), where at each time $t$, an online algorithm $\mathcal{A}$ has to choose a binary action $x_{t}$ before two binary sequences $b_{t}^{1},b_{t}^{2}$’s are revealed.
+The cost of $\mathcal{A}$ on sequence $k$ is its Hamming distance between $\{x_{t}\}_{t=1}^{T}$ and $\{b_{t}^{k}\}_{t=1}^{T},k=1,2$, and the overall cost of $\mathcal{A}$ is the maximum of its cost on the two sequences.
+Even for this case, we show that the min-max static regret of any online algorithm is $\Omega(T)$ with adversarial input.
+With stochastic input where $(b_{t}^{1},b_{t}^{2})$ are i.i.d.
+across time, we propose an algorithm with $O(1)$ true min-max static regret.
+In comparison, for the classical version Feder et al.
+(2002); Cover (1966) with only a single sequence $b_{t}^{1}$, Follow-the-Leader (FTL) algorithm achieves $O(1)$ regret when the input is i.i.d.
+Feder et al.
+(2002), while Cover (1966) achieves $O(\sqrt{T})$ regret in the adversarial case.
+We discuss the applications of multi-objective OCO and research questions left open in Appendix A, while numerical results are provided in Appendix 10.
+3 Assumptions
+We make the following assumptions that are standard in the OCO literature Hazan (2019).
+Assumption 1 (Convexity).
+$\mathcal{X}\subset{\mathbb{R}}^{d}$ is the admissible set that is closed, convex and has a finite Euclidean diameter $D$.
+Cost functions $f_{t},g_{t}:\mathcal{X}\mapsto\mathbb{R}$ are convex for all $t\geq 1$, and are bounded $|f_{t}(x)|\leq B,|g_{t}(x)|\leq B,\ \forall\ x\in\mathcal{X}$.
+Assumption 2 (Lipschitzness).
+[…]
+Assumption 3.
+Information Structure: On round $t,$ the online algorithm $\mathcal{A}$ first chooses an admissible action $x_{t}\in\mathcal{X}$ and then the adversary chooses two convex cost functions $f_{t}:\mathcal{X}\to\mathbb{R}$ and $g_{t}:\mathcal{X}\to\mathbb{R}$.
+Once the action $x_{t}$ has been chosen, both $f_{t}(x_{t}),\nabla f_{t}(x_{t})$ and $g_{t}(x_{t}),\nabla g_{t}(x_{t})$ are revealed, as is standard in the literature.
+We will also consider the bandit information structure in Section 7 where no gradient information is available.
+4 Warmup
+In this section, we highlight the difficulty in solving the min-max OCO Problem (3) in the adversarial input setting.
+To illustrate the basic difficulty, we even allow an online algorithm to know $f_{t},g_{t}$ completely before choosing its action $x_{t}$ at time $t$.
+Under this enhanced information setting, consider the greedy (locally optimal) algorithm that chooses $x_{t}=\min_{x\in\mathcal{X}}\max\{f_{t}(x),g_{t}(x)\}.$ Recall that when $g_{t}=f_{t}\ \forall\ t$, i.e.
+in the OCO setting, this algorithm will have non-positive regret $R^{\text{s}ingle}_{\mathcal{A}}(T)$ since $f_{t}(x_{t})\leq f_{t}(x)$ for any $x$.
+However, as we show next, the greedy algorithm suffers a linear min-max static regret (3).
+Let $T=2N$ and for $j=1,\dots,N$, consider the following linear functions on $\mathcal{X}=[0,1]$,
+$f_{2j-1}(x)=1.2-0.2x,$ $\qquad f_{2j}(x)=x,$
+$\hskip-72.26999ptg_{2j-1}(x)=x,$ $\qquad g_{2j}(x)=0.8+0.2x.$ (9)
+Clearly, $\arg\min_{x\in\mathcal{X}}\{\max(f_{2j-1}(x),g_{2j-1}(x))\}=1\ \text{and}\arg\min_{x\in\mathcal{X}}\{\max(f_{2j}(x),g_{2j}(x))\}=0.$
+Thus, the greedy algorithm chooses $x_{2j-1}=1$ and $x_{2j}=0,$ and the cost $C_{\mathcal{A}}$ (1) of the greedy algorithm is $1.8N$.
+However, if an algorithm $\mathcal{A}_{1}$ chooses a single static action $x^{\star}=0$ for all times, its cost $C_{\mathcal{A}_{1}}$ (3) is $1.2N$, and hence $C_{\mathsf{OPT}}\leq 1.2N.$ Thus, the min-max static regret (3) of the greedy algorithm is linear in $T$.
+5 Lower bound on min-max static regret (3) for all online algorithms
+We formalize the intuition presented in Section 4 as follows.
+Theorem 4.
+The min-max static regret (3) for any online algorithm is $\Omega(T)$.
+The proof of Theorem 4 can be found in Appendix C.
+The lower bound derived in Theorem 4 uses the fact that both $f_{t}$ and $g_{t}$ are convex and not strongly convex.
+With strongly convex functions, the lower bound may weaken.
+In light of Theorem 4, next, we consider the stochastic i.i.d.
+case for solving the min-max OCO problem which is also non-trivial.
+6 Stochastic i.i.d. input model
+We consider the i.i.d.
+input model, where at time $t$, loss function pair $(f_{t},g_{t})$ is chosen independently and is identically (jointly) distributed as $\mathcal{D}$ (unknown) across all $t$.
+The objective is to minimize the expected min-max static regret (8).
+Remark 2.
+We note that only two loss function sequences are considered for simplicity of exposition since it captures the basic difficulty of the problem, and all the results that we derive apply directly to the case when there are $K$ function sequences (see details in Remark 6).
+6.1 Algorithm
+In this section, we present an algorithm to solve the i.i.d. min-max OCO problem (8).
+Algorithm 1 Hedge+OGD
+1:Input: step sizes $\eta_{x,t}>0$, $\eta_{\theta,t}>0$, feasible set $\mathcal{X}$
+2:Initialize: $w_{1}=(1,1)$, $\theta_{1}=w_{1}/\|w_{1}\|_{1}$, $f_{0}\equiv 0,g_{0}\equiv 0$ and play arbitrary $x_{0}\in\mathcal{X}$
+3:for $t=1,\dots,T$ do
+4:  Observe $(f_{t-1}(x_{t-1}),g_{t-1}(x_{t-1}))$ and $(\nabla f_{t-1}(x_{t-1}),\nabla g_{t-1}(x_{t-1}))$
+5:  Form convex function $x\mapsto{\tilde{\Lambda}}_{t}(x,\theta_{t}):=\theta_{t,1}f_{t-1}(x)+\theta_{t,2}g_{t-1}(x)$
+6:  OGD: $y_{t}=x_{t-1}-\eta_{x,t}\nabla{\tilde{\Lambda}}_{t}(x_{t-1},\theta_{t})$ and find $x_{t}=\text{Proj}_{\mathcal{X}}(y_{t})$
+7:  Play action $x_{t}$  Evaluate $\lambda_{t}=(f_{t}(x_{t}),g_{t}(x_{t}))$
+8:  Hedge: $w_{t+1,i}=w_{t,i}\exp(\eta_{\theta,t}\lambda_{t,i}),i=1,2$
+9:  Normalize $\theta_{t+1}=w_{t+1}/\|w_{t+1}\|_{1}$
+10:end for
+Towards that end, if suppose an online algorithm knew the optimizing $\theta^{\star}$ in (7), then by executing an OGD algorithm for OCO with a single function $\theta_{1}^{\star}f_{t}+\theta_{2}^{\star}g_{t}$ at time $t+1$, it would achieve a regret of $O(\sqrt{T})$ using standard results Hazan (2019).
+However, the optimizing $\theta^{\star}$ that depends on $\mathcal{D}$ (unknown) remains unknown, but, this observation motivates our algorithm that constructs
+i) $\theta_{t}$ at time $t$ as a surrogate for $\theta^{\star}$ using the Hedge algorithm (gains version) treating the two function sequences $\{f_{t}\}_{t=1}^{T}$ and $\{g_{t}\}_{t=1}^{T}$ as the gains of the two experts (see Section J for basic definition of the Experts Problem and the Hedge algorithm), and
+ii) for each fixed value of $\theta_{t}$, plays the action chosen by OGD applied to the single surrogate function $\theta_{t1}f_{t}+\theta_{t2}g_{t}$ at time $t+1$. Pseudocode is given in Algorithm 1.
+6.2 Guarantee
+Theorem 5.
+Under Assumptions 1 and 2, for $\eta_{\theta,t}\leq\sqrt{\frac{2\ln 2}{B^{2}t}},\eta_{x,t}=\frac{D}{G\sqrt{t}}$, the expected min-max static regret (8) of the proposed algorithm (Algorithm 1) is
+${\bar{R}}_{\mathcal{A}}(T)=O(\sqrt{T}).$
+Proof.
+The main ingredient of the proof as well as the inspiration for Algorithm 1 is the following decomposition (10) of the expected min-max static regret (8).
+Let $\Delta_{K}$ be the $K$-dimensional simplex.
+Recall the definition of $\Lambda_{t}(x,\theta):=\theta_{1}f_{t}(x)+\theta_{2}g_{t}(x)$, where $\theta=(\theta_{1},\theta_{2})\in\Delta_{2}$.
+By adding and subtracting two terms, $\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t})$ and $\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta_{t})$ inside the expectation in the RHS of (8), we write the expected min-max static regret (8) as
+${\bar{R}}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{1}\}+{\mathbb{E}}\{R_{2}\}+{\mathbb{E}}\{R_{3}\}\}\leq\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{1}\}\}+\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{2}\}\}+\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{3}\}\},$ (10)
+where
+$R_{1}=\max_{\theta\in\Delta_{2}}\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta)-\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t}),$
+$R_{2}=\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t})-\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta_{t}),$
+and
+$R_{3}=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta_{t})-{\bar{C}}_{\mathsf{OPT}}=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta_{t})-T\cdot\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x).$
+The interpretation of $R_{1}$ is that it is the regret for the gains version of $2$-Experts Problem Cesa-Bianchi and Lugosi (2006), where the two function sequences $f_{t},g_{t}$ are the gains of two experts, with respect to an optimal $\theta^{\star}\in\Delta_{2}$ for fixed actions $x_{t}$ across time slots.
+This motivates the use of Hedge algorithm (gains version) in our proposed algorithm that is tailor-made for updating $\theta_{t}$ for fixed actions $x_{t}$.
+Similarly, $R_{2}$ is the static regret for the OCO problem relative to static optimal action $x^{\star}$ for a single surrogate objective function sequence of $\Lambda_{t}(x,\theta_{t})=\theta_{t1}f_{t}(x)+\theta_{t2}g_{t}(x)$ for a fixed $\theta_{t}$.
+Hence using an OGD algorithm is natural for this purpose as done in the algorithm for each time $t$, with a fixed weight $\theta_{t}$.
+The third term $R_{3}$ is more nuanced and captures the difference between the cost of $\mathsf{OPT}$ for a fixed sequence of $\theta_{t},t=1,\dots,T$ and the optimal benchmark cost ${\bar{C}}_{\mathsf{OPT}}$ (7), respectively.
+We bound $R_{1}$ and $R_{2}$ sample-path wise for each realization of $f_{t},g_{t}$, while for $R_{3}$ we will bound ${\mathbb{E}}\{R_{3}\}$ as follows whose proofs are provided in Appendix E, F, and G, respectively
+Lemma 6.
+For $\eta_{\theta,t}\leq\sqrt{2\ln 2/(B^{2}t)}$, $\sup_{\mathcal{D}}\{R_{1}\}\leq O(\sqrt{T})$.
+Lemma 7.
+Choosing $\eta_{x,t}=\frac{D}{G\sqrt{t}}$, $\sup_{\mathcal{D}}\{R_{2}\}\leq O(\sqrt{T})$.
+Lemma 8.
+${\mathbb{E}}\{R_{3}\}\leq 0$.
+Combining Lemma 6, 7 and 8, and the regret decomposition (10), we get that the expected min-max static regret ${\bar{R}}_{\mathcal{A}}(T)=O(\sqrt{T})$ (8) with $\eta_{\theta,t}\leq\sqrt{\frac{2\ln 2}{B^{2}t}},\text{and}\ \eta_{x,t}=\frac{D}{G\sqrt{t}}$.
+∎
+Remark 3.
+The regret decomposition (10) does multiple things at the same time i) identifies the critical difficulty of the problem, ii) motivates a natural algorithm, and ii) keeps the proof modular, where the regret bound for each module is simple and elegant, almost first principles based.
+Remark 4.
+Although the proof of Theorem 5 is remarkably simple and elegant, it hinges on the critical regret decomposition (10) – which is not obvious a priori.
+This decomposition eliminates the need for sophisticated technical machinery needed in related prior work Mahdavi et al.
+(2013) that also considers the i.i.d.
+input and highlights the underlying simplicity of the result.
+We would like to emphasize that the elegance of the argument should not obscure its importance.
+Remark 5.
+For the $\alpha$-fair problem, Si Salem et al.
+(2023) derived a regret guarantee of $O(\alpha\sqrt{T})$, which becomes vacuous for our regime of interest since $\alpha$-fair problem is equivalent to our min-max setting when $\alpha\rightarrow\infty$.
+Compared to Si Salem et al.
+(2023), Algorithm 1, has an $\alpha$-independent regret guarantee.
+It is tempting to use Algorithm 1 for the $\alpha$-fair problem for finite $\alpha$, however, as we point out in Remark 11, the $\alpha$-fair problem with finite $\alpha$ and $\alpha\rightarrow\infty$ are fundamentally different.
+Remark 6.
+When there are $K$ loss function sequences $f_{t}^{k},k=1,\dots,K$ that are chosen i.i.d.
+from $\mathcal{D}$, define $\theta\in\Delta_{K}$, (where $\Delta_{K}$ is the $K$-dimensional simplex) and $\Lambda_{t}(x,\theta)=\sum_{k=1}^{K}\theta_{k}f_{t}^{k}$.
+With $\mu(x)={\mathbb{E}}\{f_{t}^{1}(x),\dots,f_{t}^{K}(x)\}$, analogous to (8), the expected min-max static regret of $\mathcal{A}$ is
+${\bar{R}}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}{\mathbb{E}}\left\{\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta)-T\cdot\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{K}}\theta^{\top}\mu(x)\right\}.$ (11)
+With this new definition of $\theta$ and $\Lambda_{t}$, Algorithm 1 remains as is with Hedge choosing $\theta_{t}\in\Delta_{K}$ while $x_{t}$ the single action chosen by OGD with respect to the surrogate loss function $\Lambda_{t}(x,\theta_{t})$ with fixed $\theta_{t}$.
+Assuming all $f_{t}^{k},k=1,\dots,K$ satisfy Assumption 1 and 2, the only change in the guarantee with $K>2$ compared to Theorem 5 with $K=2$ is because of Lemma 6 that is now $O(\sqrt{T\log K})$ Hazan (2019) since there are $K$ experts corresponding to the $K$ loss function sequences $f_{t}^{k},k=1,\dots,K$ .
+The OGD guarantee of Lemma 7 stays the same since $\|\nabla\Lambda_{t}(x_{t},\theta)\|\leq G$ as before.
+Moreover, Lemma 8 holds as is.
+Thus, the overall expected min-max static regret of Algorithm 1 with $K$ loss function sequences is $O(\sqrt{T\log K}).$
+Remark 7.
+[Extension to Strongly Convex $f_{t}$’s and $g_{t}$’s] One limitation of our proposed algorithm and the proof is that because of the use of the Hedge algorithm for updating $\theta_{t}$, the regret $\sup_{\mathcal{D}}\{R_{1}\}$ can at best be $O(\sqrt{T})$ with stochastic i.i.d.
+input without any further assumptions.
+Thus, unlike OCO, where OGD algorithm achieves logarithmic regret in $T$ when $f_{t}$’s are strongly convex, our algorithm’s regret is only $O(\sqrt{T})$ even when both $f_{t},g_{t}$ are strongly convex.
+Is this a limitation of the algorithm or of that of the problem itself remains an interesting unresolved question.
+7 Extension to Bandit Feedback
+We now analyze the expected min-max OCO problem (8) under bandit feedback: at each time $t$, after choosing $x_{t}$, the algorithm observes only the scalar losses $\{f_{t-1}^{k}(\cdot)\}_{k=1}^{K}$ at one point (one-point model) or two points (two-point model), and receives no gradient information.
+For this bandit model, we consider a simple adaptation of Algorithm 1 by replacing its OGD part with the bandit information counterpart, while keeping the Hedge part as it is, given by Algorithm 2.
+Exploiting the modular nature of the proof in Section 6, the results follow almost immediately.
+Algorithm 2 Bandit-min-max
+1:Input:Feasible set $\mathcal{X}\subset\mathbb{R}^{d}$, parameter $\{\delta,\rho\}$, stepsizes $\{\eta_{x,t}\},\{\eta_{\theta,t}\}$, $\mathsf{mode}\in\{\text{one-point},\text{two-point}\}$ Initialize $w_{1}=(1,1,\dots,1)_{K}$, $\theta_{1}=w_{1}/\|w_{1}\|_{1}$, choose $x_{1}\in(1-\rho)\mathcal{X}$
+2:for $t=1,\dots,T$ do Sample $u_{t}\sim\mathrm{Unif}(\mathbb{S}^{d-1})$
+3:  if $\mathsf{mode}=\text{one-point}$ then Query $\tilde{x}_{t}=x_{t}+\delta u_{t}$  Observe $\{f_{t}^{k}(\tilde{x}_{t})\}_{k=1}^{K}$$\widehat{\nabla}_{t}=\frac{d}{\delta}\Lambda_{t}(\tilde{x}_{t},\theta_{t})u_{t}$ where $\Lambda_{t}(\tilde{x}_{t},\theta_{t})=\sum_{k=1}^{K}\theta_{t,k}f_{t}^{k}(\tilde{x}_{t})$$\ell_{t,k}=f_{t}^{k}(\tilde{x}_{t})$
+4:  elseQuery $x_{t}^{+}=x_{t}+\delta u_{t}$, $x_{t}^{-}=x_{t}-\delta u_{t}$  Observe $\{f_{t}^{k}(x_{t}^{+})\}_{k=1}^{K}$ and $\{f_{t}^{k}(x_{t}^{-})\}_{k=1}^{K}$$\widehat{\nabla}_{t}=\frac{d}{2\delta}\big(\Lambda_{t}(x_{t}^{+},\theta_{t})-\Lambda_{t}(x_{t}^{-},\theta_{t})\big)u_{t}$$\ell_{t,k}=\tfrac{1}{2}\left(f_{t}^{k}(x_{t}^{+})+f_{t}^{k}(x_{t}^{-})\right)$
+[…]
+6:end for
+Theorem 9.
+Let $\mathsf{r}$} be such that $r\mathcal{B}_{1}\subseteq\mathcal{X}\subseteq D\mathcal{B}_{1}$ and $\mathcal{B}_{1}$ is the unit ball.
+  1. 1.
+    (One-point feedback) With $\delta=T^{-1/4}$, $\eta_{x,t}=D/(Gd^{1/2}t^{3/4})$, $\rho=\delta/{\mathsf{r}}$, and $\eta_{\theta,t}=\sqrt{\log K}/(B\sqrt{t})$, expected min-max static regret ${\bar{R}}_{\mathcal{A}}(T)\leq c_{1}\,d^{1/2}GD\,T^{3/4}\;+\;c_{2}B\sqrt{T\log K}$, for constants $c_{1},c_{2}>0$.
+  2. 2.
+    (Two-point feedback) With $\delta=T^{-1/2}$, $\rho=\delta/{\mathsf{r}}$, $\eta_{x,t}=D/(Gd^{1/2}\sqrt{t})$, and $\eta_{\theta,t}=\sqrt{\log K}/(B\sqrt{t})$, expected min-max static regret satisfies ${\bar{R}}_{\mathcal{A}}(T)\leq c_{3}\,d^{1/2}GD\,\sqrt{T}\;+\;c_{4}B\sqrt{T\log K}$, for constants $c_{3},c_{4}>0$.
+Proof of Theorem 9 can be found in Appendix I.
+Discussion: Compared to the limited information setting considered in Section 3, where gradient information is available, with bandit feedback the expected min-max regret guarantees suffer, however, the degraded guarantees match the best known guarantees for OGD in the bandit setting.
+The proof of the result is simple because of the modularity of our algorithm and the regret decomposition that decomposes the total regret into regret resulting from Hedge and OGD, and the fact that only regret for OGD changes with bandit feedback.
+8 Binary Prediction Problem
+In this section, we consider the min-max universal prediction of binary sequences, where at each time $t$, two binary sequences $b_{t}^{1},b_{t}^{2}\in\{0,1\}$ are revealed after an online algorithm $\mathcal{A}$ (predicts) chooses an action $x_{t}\in\{0,1\}$, and cost of $\mathcal{A}$ is
+$C^{b}_{\mathcal{A}}\;=\;\max\left\{\sum_{t=1}^{T}\mathbf{1}\{x_{t}\neq b_{t}^{1}\},\;\sum_{t=1}^{T}\mathbf{1}\{x_{t}\neq b_{t}^{2}\}\right\}.$
+Similar to (2),
+$C^{b}_{\mathsf{OPT}}\;=\;\min_{x\in\{0,1\}}\max\Big\{\sum_{t=1}^{T}\mathbf{1}\{x\neq b_{t}^{1}\},\;\sum_{t=1}^{T}\mathbf{1}\{x\neq b_{t}^{2}\}\Big\},$
+and the regret is defined as $R^{b}_{\mathcal{A}}(T)=\max_{b_{t}^{1},b_{t}^{2}}\{C^{b}_{\mathcal{A}}-C^{b}_{\mathsf{OPT}}\}.$
+Similar to Theorem 4, we show that the min-max static regret $R^{b}_{\mathcal{A}}(T)$ for any online algorithm is $\Omega(T)$ with adversarial input in contrast to the classical setting Cover (1966); Feder et al.
+(2002) where with a single binary sequence the regret is $O(\sqrt{T})$ with adversarial input.
+Theorem 10.
+The min-max static regret $R^{b}_{\mathcal{A}}(T)$ for any online algorithm is $\Omega(T)$.
+The proof of Theorem 10 can be found in Appendix K. Thus, once again we consider the stochastic i.i.d. input as follows.
+8.1 $(b_{t}^{1},b_{t}^{2})$ i.i.d. across time
+In this section, we consider that $(b_{t}^{1},b_{t}^{2})$ are generated from an unknown distribution with i.i.d.
+across time.
+Under this setting, we can directly use Theorem 5 to show that Algorithm 1 can be used to get a true min-max static regret bound of ${\mathbb{E}}\{R^{b}_{\mathcal{A}}(T)\}=O(\sqrt{T\log T})$, see details in Appendix L.
+Next, we show that in this binary case, we can do better than Algorithm 1 and get true min-max static regret bound ${\mathbb{E}}\{R^{b}_{\mathcal{A}}(T)\}=O(1)$ without involving the averaged-out benchmark ${\bar{C}}_{\mathsf{OPT}}^{b}$ that is defined similar to (7).
+The main challenge in this binary min-max prediction setting arises in disagreement regimes, where the two sequences $\{b_{t}^{1}\}_{t=1}^{T},\{b_{t}^{2}\}_{t=1}^{T}$ favor opposite actions.
+In such cases, simply following a majority or resolving ties arbitrarily can lead to persistent oscillations, since each sequence individually advocates a different decision.
+To achieve low expected min-max static regret, an algorithm must determine which of these conflicting preferences induces the larger worst-case cost and commit to that action.
+To address this we propose Adaptive Stronger-Bias Double Majority (ASB-DM) algorithm (Algorithm 4 in Appendix M) that exploits the strength of the statistical evidence, rather than only its direction.
+When the empirical preferences of the two sequences disagree, ASB-DM aggregates the evidence by comparing the combined bias toward the two actions and selects the one with the smaller inferred worst-case loss.
+As the empirical estimates concentrate, this stronger-bias rule allows the algorithm to reliably distinguish the optimal action from the suboptimal one, ensuring that the decision stabilizes and the total number of mistakes remains finite.
+Theorem 11.
+Assume the pairs $(b_{t}^{1},b_{t}^{2})$ are generated i.i.d.
+across time $t$ from a joint distribution with marginals $b_{t}^{1}\sim\mathrm{Ber}(q_{1})$ and $b_{t}^{2}\sim\mathrm{Ber}(q_{2})$.
+If the preferences are asymmetric (i.e., $q_{1},q_{2}\neq 1/2$ and $q_{1}+q_{2}\neq 1$), then the true min-max regret of Algorithm 4 satisfies ${\mathbb{E}}\{R^{b}_{\mathcal{A}}(T)\}=O(1)$.
+Proof of Theorem 11 is provided in Appendix M. Note that extension to the case when there are $K\geq 2$ sequences $b_{t}^{k},k=1,\dots,K$ is immediate with $O(1)$ term scaling linearly with $K$.
+Remark 8 (Singularity at $q_{1}+q_{2}=1$).
+The necessity of the condition $q_{1}+q_{2}\neq 1$ is best illustrated by the example where $q_{1}=0.4$ and $q_{2}=0.6$.
+To determine the static optimal action, we compare the worst-case expected costs: action 0 incurs $0.6T$ (driven by Sequence 2), which is identical to the $0.6T$ cost incurred by action $1$ (driven by Sequence 1).
+Since the worst-case costs are identical, the static optimal action is not unique.
+At this singularity, any online algorithm faces an inherent lack of statistical separation between the actions.
+Because the expected costs are perfectly balanced, empirical observations will fluctuate around the true parameters due to random sampling noise.
+Without a strictly positive gap to stabilize the decision process, any algorithm will be forced to switch between actions repeatedly—effectively behaving as a symmetric random walk—leading to an accumulation of regret at a rate of $\Omega(\sqrt{T})$.
+9 Conclusions
+In this paper, we have made a basic advance in the area of multi-objective OCO by formulating the min-max static regret minimization problem, that requires an algorithm to closely track multiple loss function sequences simultaneously and at all times.
+Because of the lower bound on regret that scales linearly with the time horizon under an adversarial input, we considered the i.i.d.
+input setting, for which we proposed a simple algorithm that used a combination of the well-known Hedge and OGD algorithm.
+Using an insightful regret decomposition, which made the analysis modular, we derived an $O(\sqrt{T\log(TK)})$ regret bound on the proposed algorithm.
+We also extended the results to the bandit information setting.
+Certain questions that remain open are: i) how to handle the case when all loss functions are strongly convex, and ii) handling constraints like COCO.
+We also generalized the classical problem of universal prediction of binary sequences to the case when there are multiple sequences.
+Even for this binary case, we showed that the min-max static regret of any online algorithm scales linearly with the time horizon, exposing the inherent difficulty of the min-max setting.
+With stochastic input where $(b_{t}^{1},b_{t}^{2})$ are i.i.d.
+across time, we proposed an algorithm that chooses its action on the strength of statistical evidence and showed that it has $O(1)$ expected min-max static regret.
+Appendix A Applications of Multi-objective OCO and Open Questions
+A.1 Applications
+The min-max approach is also well suited for modern networked systems—such as data centers, wireless edge networks, and shared cloud infrastructures—where online decisions must balance performance across multiple service classes, users, or flows whose objectives (throughput, latency, energy) often conflict.
+The min-max regret criterion provides a principled way to ensure worst-case fairness by minimizing the largest cumulative loss among competing entities, which is crucial in fair resource allocation [43], fair allocation in clouds [19], fair assignment of reusable resources [28] congestion-controlled transport [24, 3], and dynamic resource provisioning in data centers [45].
+Unlike average-regret formulations that can conceal systematic bias, minimizing min-max regret explicitly guards against persistent under-allocation to any group or flow—an increasingly important property in multi-tenant and QoS-sensitive networks.
+Multi-objective OCO has also been studied from an applications perspective, e.g.
+[30] studied the problem of multi-period multi-objective online ride-matching problem, where the multiple objectives included platform revenue, assigning drivers with higher service quality, minimizing the distance to the customer etc.
+[30] also used the COCO-type reformulation of multi-objective OCO for deriving theoretical guarantees.
+Other applications of multi-objective OCO include multi-task learning [38], multiple target tracking [33], portfolio selection [12] and safe reinforcement learning [10], where decision-makers must make sequential decisions under uncertainty, balancing objectives such as cost, delivery time, and reliability without full knowledge of future demand or disruptions.
+A.2 Open Questions
   • Strongly Convex Functions: For OCO, the best-known static regret reduces to $O(\log{T})$ when functions $f_{t}$’s are strongly convex.
@@
     respectively.
-It turns out that our proposed algorithm cannot be married readily with the prior best known approaches for COCO Sinha and Vaze (2024); Vaze and Sinha (2025); Sinha and Vaze (2025) for deriving simultaneous bounds on regret and CCV.
+It turns out that our proposed algorithm cannot be married readily with the prior best known approaches for COCO [41, 48, 42] for deriving simultaneous bounds on regret and CCV.
 Thus, how to solve the min-max OCO with constraints remains an interesting open problem.
-2 Assumptions
-We start by stating the assumptions that we will use throughout, that are standard in the OCO literature (Hazan, 2019).
-Assumption 1 (Convexity).
-$\mathcal{X}\subset{\mathbb{R}}^{d}$ is the admissible set that is closed, convex and has a finite Euclidean diameter $D$.
-Cost functions $f_{t},g_{t}:\mathcal{X}\mapsto\mathbb{R}$ are convex for all $t\geq 1$, and are bounded $|f_{t}(x)|\leq B,|g_{t}(x)|\leq B,\ \forall\ x\in\mathcal{X}$.
-Assumption 2 (Lipschitzness).
-[…]
-Assumption 3.
-Information Structure: On round $t,$ the online algorithm $\mathcal{A}$ first chooses an admissible action $x_{t}\in\mathcal{X}$ and then the adversary chooses two convex cost functions $f_{t}:\mathcal{X}\to\mathbb{R}$ and $g_{t}:\mathcal{X}\to\mathbb{R}$.
-Once the action $x_{t}$ has been chosen, both $f_{t}(x_{t}),\nabla f_{t}(x_{t})$ and $g_{t}(x_{t}),\nabla g_{t}(x_{t})$ are revealed, as is standard in the literature.
-We will also consider the bandit information structure in Section 6 where no gradient information is available.
-3 min-max OCO Problem (5) in the Adversarial Setting
-In this section, we highlight the difficulty in solving the min-max OCO Problem (5) in the adversarial input setting, where the input can be chosen adversarially depending on the choice of an online algorithm.
-To illustrate the basic difficulty, we even allow the algorithm to know $f_{t},g_{t}$ completely before choosing its action $x_{t}$ at time $t$.
-Under this enhanced information setting, consider the greedy (locally optimal) algorithm that chooses
-$x_{t}=\min_{x\in\mathcal{X}}\max\{f_{t}(x),g_{t}(x)\}.$
-Recall that when $g_{t}=f_{t}\ \forall\ t$, i.e.
-in the OCO setting, this algorithm will have non-positive regret $R^{\text{s}ingle}_{\mathcal{A}}(T)$ (1) since $f_{t}(x_{t})\leq f_{t}(x)$ for any $x$.
-However, as we show next, the greedy algorithm suffers a linear min-max static regret (5).
-Let $T=2N$ and consider the following linear functions on $\mathcal{X}=[0,1]$, for $j=1,\dots,N$
-$f_{2j-1}(x)=1.2-0.2x,$ $\qquad f_{2j}(x)=x,$
-$\hskip-72.26999ptg_{2j-1}(x)=x,$ $\qquad g_{2j}(x)=0.8+0.2x.$ (15)
-Clearly,
-$\arg\min_{x\in\mathcal{X}}\{\max(f_{2j-1}(x),g_{2j-1}(x))\}=1$ $\quad\text{and}\quad\arg\min_{x\in\mathcal{X}}\{\max(f_{2j}(x),g_{2j}(x))\}=0.$
-Thus, the greedy algorithm chooses $x_{2j-1}=1$ and $x_{2j}=0,$ and the cost $C_{\mathcal{A}}$ (2) of the greedy algorithm is
-$\max\left\{\sum_{j=1}^{N}\left(f_{2j-1}(x_{2j-1})+f_{2j}(x_{2j})\right),\sum_{j=1}^{N}\left(g_{2j-1}(x_{2j-1})+g_{2j}(x_{2j})\right)\right\}$
-$=\max\left\{\sum_{j=1}^{N}\left(f_{2j-1}(1)+f_{2j}(0)\right),\sum_{j=1}^{N}\left(g_{2j-1}(1)+g_{2j}(0)\right)\right\}=\max\left\{\sum_{j=1}^{N}(1+0),\sum_{j=1}^{N}(1+0.8)\right\}=1.8N.$
-However, if an algorithm chooses a single static action $x^{\star}=0$ for all times, its cost $C_{\mathcal{A}}$ (5) is $1.2N$, and hence
-$C_{\mathsf{OPT}}\leq 1.2N.$
-Thus, the min-max static regret (5) of the greedy algorithm is linear in $T$.
-This example highlights the key difficulty in minimizing the min-max static regret in the adversarial/worst case input.
-We note that this example does not preclude more intelligent algorithms that take a global view in trying to balance the two costs, such as primal-dual, OGD+Hedge or FTRL to achieve sub-linear min-max static regret (5), however, the stringent min-max static regret definition makes the analysis highly challenging.
-We will make additional remarks on the technical difficulty in Appendix F.
-Thus, from here on we consider the i.i.d.
-input that is defined next.
-4 i.i.d. Input model
-We consider the i.i.d.
-input model, where at time $t$, loss functions $(f_{t},g_{t})$ are chosen independently and are identically jointly distributed as $\mathcal{D}$ (unknown) across all $t$.
-We assume that Assumption 1 and 2 holds, and consider the same information structure as defined in Assumption 3.
-Let $\lambda_{t}=(f_{t},g_{t})$ and recall that $\Delta_{2}$ is the 2D2-D simplex, $\theta=(\theta_{1},\theta_{2})\in\Delta_{2}$, and
-$\Lambda_{t}(x,\theta):=\theta_{1}f_{t}(x)+\theta_{2}g_{t}(x).$
-Following (2), the expected cost of any online algorithm $\mathcal{A}$ with the i.i.d. input which chooses actions $x_{t}$ is
-${\bar{C}}_{\mathcal{A}}={\mathbb{E}}\left\{\max\Big\{\sum_{t=1}^{T}f_{t}(x_{t}),\sum_{t=1}^{T}g_{t}(x_{t})\Big\}\right\}={\mathbb{E}}\left\{\max_{\theta\in\Delta_{2}}\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta)\right\}.$ (16)
-while from (11)
-${\bar{C}}_{\mathsf{OPT}}=T\cdot\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x)\quad\text{with}\quad(x^{\star},\theta^{\star})=\arg\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x),$ (17)
-where $\mu(x)={\mathbb{E}}\{f_{t}(x),g_{t}(x)\}$.
-Consequently, the expected min-max static regret ${\bar{R}}_{T}(\mathcal{A})$ following (12) is
-${\bar{R}}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}\left\{{\bar{C}}_{\mathcal{A}}-{\bar{C}}_{\mathsf{OPT}}\right\}.$ (18)
-We refer to (18) as the i.i.d. min-max OCO problem and for the rest of the paper, consider it with the goal of deriving an algorithm that has $o(T)$ expected min-max static regret.
+Appendix B Related Work on COCO
+Because of its widespread applications, multi-objective OCO has been an object of immense interest, however, with limited success.
+For example, [32] wanted to study the multi-objective OCO and stated the difficulty as “A fundamental difference between single- and multi-objective optimization is that for the latter it is not obvious how to evaluate the optimization quality”.
+Further, it said “Since it is impossible to simultaneously minimize multiple loss functions and in order to avoid complications caused by handling more than one objective, we choose one function as the objective and try to bound other objectives by appropriate thresholds.
+This work gave rise to the now popular problem of constrained online convex optimization (COCO), where the problem is transformed into a constrained optimization problem with one objective and one or more constraint functions, that is described next.
+B.0.1 COCO
+In COCO, on every round $t,$ an online algorithm first chooses an admissible action $x_{t}\in\mathcal{X}\subset{\mathbb{R}}^{d},$ and then the adversary chooses a convex loss/cost function $f_{t}:\mathcal{X}\to\mathbb{R}$ and a constraint function of the form $g_{t}(x)\leq 0,$ where $g_{t}:\mathcal{X}\to\mathbb{R}$ is a convex function.
+Let $\mathcal{X}^{\star}$ be the feasible set consisting of all admissible actions that satisfy all constraints $g_{t}(x)\leq 0,t\in[T]$.
+A standard assumption is made that $\mathcal{X}^{\star}$ is not empty (called the feasibility assumption).
+Since $g_{t}$’s are revealed after the action $x_{t}$ is chosen, an online algorithm $\mathcal{A}$ need not take feasible actions on each round, and in addition to the static regret defined as
+$\textrm{R}^{\text{COCO}}_{\mathcal{A}}(T)\equiv\sup_{\{f_{t}\}_{t=1}^{T}}\left\{\sum_{t=1}^{T}f_{t}(x_{t})-\inf_{x\in\mathcal{X}^{\star}}\sum_{t=1}^{T}f_{t}(x^{\star})\right\},$ (14)
+an additional metric of interest is the total cumulative constraint violation (CCV) defined as
+$\textrm{CCV}_{\mathcal{A}}(T)\equiv\sum_{t=1}^{T}\max(g_{t}(x_{t}),0).$
+The goal with COCO is to design an online algorithm to simultaneously achieve a small regret (14) with respect to any admissible benchmark $x^{\star}\in\mathcal{X}^{\star}$ and a small CCV.
+B.0.2 Prior Work on COCO
+(A) i.i.d.
+loss and constraints: COCO with independent and identically distributed (i.i.d.) loss functions $f_{t}$’s and constraint functions $g_{t}$’s was studied in [32] while its extension to stationary and ergodic processes can be found in [46], where both derived an algorithm with regret $O(\sqrt{T})$ and CCV $O(T^{3/4})$.
+(B) Time-invariant constraints: COCO with time-invariant constraints, i.e., $g_{t}=g,\forall\ t$ [53, 21, 31, 50] has been considered extensively, where functions $g$ are assumed to be known to the algorithm a priori.
+The best known results in this case give static regret of $O(T^{\max\{c,1-c\}})$ and CCV of $O(T^{(1-c)/2})$ for any $c\in(0,1)$.
+The algorithm is allowed to take actions that are infeasible at any time to avoid the costly projection step of the vanilla projected OGD algorithm and the main objective was to design an efficient algorithm with a small regret and CCV while avoiding the explicit projection step.
+(C) Time-varying constraints: The more difficult question is solving COCO problem when the constraint functions, i.e., $g_{t}$’s, change arbitrarily with time $t$.
+One popular algorithm for solving COCO considered a Lagrangian function optimization that is updated using the primal and dual variables [52, 44, 51].
+Alternatively, [35] and [26] used the drift-plus-penalty (DPP) framework [34] to solve the COCO, but which needed additional assumption, e.g.
+the Slater’s condition in [35] and with weaker form of the feasibility assumption in [35].
+[18] obtained bounds similar to [35] but without assuming Slater’s condition.
+Very recently, the state of the art guarantees on simultaneous bounds on regret $O(\sqrt{T})$ and CCV $O(\sqrt{T}\log T)$ for COCO were derived in [41] with a very simple algorithm that combines the loss function at time $t$ and the CCV accrued till time $t$ in a single loss function, and then executes the online gradient descent (OGD) algorithm on the single loss function with an adaptive step-size.
+Most recently [42] showed lower CCV is possible with increased regret compared to [41] while [48] obtained instance dependent CCV bounds which can be $O(1)$ for specific cases while achieving regret of $O(\sqrt{T})$, respectively.
+The COCO problem has been considered in the dynamic setting as well [8, 4, 47, 29] where the benchmark $x^{\star}$ in (14) is replaced by $x_{t}^{\star}$ ($x_{t}^{\star}=\arg\min_{x}f_{t}(x)$) that is also allowed to change its actions over time.
 Remark 9.
-Note that even i.i.d. min-max OCO problem is theoretically challenging, and no prior results apply directly.
+Even though the min-max OCO and the COCO problem appear related, but unfortunately results for one do not give results for the other in either direction.
+Appendix C Proof of Theorem 4
+Proof.
+Let the decision set be $\mathcal{X}=[-1,1]$.
+The adversary constructs two inputs $\sigma_{1}$ and $\sigma_{2}$ with common prefix until $t=T/2$ as follows.
+For both inputs $\sigma_{1}$ and $\sigma_{2}$, for rounds $t=1,\dots,T/2$,
+$f_{t}(x)=-x,\qquad g_{t}(x)=x.$
+Let $x_{t}$ be the actions of an online algorithm $\mathcal{A}$, and let
+${\bar{x}}:=\sum_{t=1}^{T/2}x_{t}$
+denote the cumulative action of $\mathcal{A}$ during the prefix. After observing $\bar{x}$}, the adversary completes the input as follows:
+  • If ${\bar{x}}<0$, the adversary selects input $\sigma_{1}$.
+  • If ${\bar{x}}\geq 0$, the adversary selects input $\sigma_{2}$, where the respective suffixes for $t=T/2+1,\dots,T$ are defined as follows.
+Suffix of $\sigma_{1}$.
+For $t=T/2+1,\dots,T$,
+$f_{t}(x)=1,\qquad g_{t}(x)=0.$
+Suffix of $\sigma_{2}$.
+For $t=T/2+1,\dots,T$,
+$f_{t}(x)=0,\qquad g_{t}(x)=1.$
+For $\sigma_{1}$, the cumulative losses for a fixed $x\in[-1,1]$ are
+$\sum_{t=1}^{T}f_{t}(x)=-\frac{T}{2}x+\frac{T}{2},\qquad\sum_{t=1}^{T}g_{t}(x)=\frac{T}{2}x.$
+The two quantities are equal at $x=\tfrac{1}{2}$, yielding
+$C_{\mathsf{OPT}}(\sigma_{1})=\frac{T}{4}.$
+Similarly, for $\sigma_{2}$,
+$\sum_{t=1}^{T}f_{t}(x)=-\frac{T}{2}x,\qquad\sum_{t=1}^{T}g_{t}(x)=\frac{T}{2}x+\frac{T}{2},$
+which are equal at $x=-\tfrac{1}{2}$, giving
+$C_{\mathsf{OPT}}(\sigma_{2})=\frac{T}{4}.$
+Thus, in both cases,
+$C_{\mathsf{OPT}}(\sigma_{i})=\frac{T}{4},\qquad i=1,2.$
+We now lower bound the cost incurred by $\mathcal{A}$.
+Case 1: ${\bar{x}}<0$ (input $\sigma_{1}$).
+$\sum_{t=1}^{T}f_{t}(x_{t})=-\sum_{t=1}^{T/2}x_{t}+\frac{T}{2}=\frac{T}{2}-{\bar{x}},$
+$\sum_{t=1}^{T}g_{t}(x_{t})=\sum_{t=1}^{T/2}x_{t}={\bar{x}}.$
+Hence,
+$C_{\mathcal{A}}(\sigma_{1})=\max\Bigl\{\frac{T}{2}-{\bar{x}},\;{\bar{x}}\Bigr\}\geq\frac{T}{2},$
+since ${\bar{x}}<0$.
+Case 2: ${\bar{x}}\geq 0$ (input $\sigma_{2}$).
+$\sum_{t=1}^{T}f_{t}(x_{t})=-\sum_{t=1}^{T/2}x_{t}=-{\bar{x}},$
+$\sum_{t=1}^{T}g_{t}(x_{t})=\sum_{t=1}^{T/2}x_{t}+\frac{T}{2}.$
+Thus,
+$C_{\mathcal{A}}(\sigma_{2})=\max\Bigl\{-{\bar{x}},\;\frac{T}{2}+{\bar{x}}\Bigr\}\;\geq\;\frac{T}{2},$
+since ${\bar{x}}\geq 0$.
+Thus,
+$\mathrm{Regret}_{\mathcal{A}}(T)=C_{\mathcal{A}}(\sigma_{i})-C_{\mathsf{OPT}}(\sigma_{i})\geq\frac{T}{4}$
+for at least one $i=1,2$. ∎
+Appendix D Connecting the two benchmarks ${\mathbb{E}}\{C_{\mathsf{OPT}}\}$ and ${\bar{C}}_{\mathsf{OPT}}$ (7)
+Without loss of generality, let $\chi\subseteq[0,1]^{d}$. We will invoke Assumptions 1 and 2 throughout without repeatedly mentioning it. Thus, for each $t=1,\dots,T$ and $x\in[0,1]^{d}$, we have
+$\lambda_{t}(x)=(f_{t}(x),g_{t}(x))\in[0,B]^{2},$
+where $\{\lambda_{t}(\cdot)\}_{t=1}^{T}$ are i.i.d. across $t$. Since $f_{t}(x),g_{t}(x)$ is $G$-Lipschitz, each coordinate of $\lambda_{t}$ is GG–Lipschitz in $x$. Let the empirical mean be
+$\hat{\mu}_{T}(x)=\frac{1}{T}\sum_{t=1}^{T}\lambda_{t}(x),$
+while as before
+$\mu(x)=\mathbb{E}\{\lambda_{t}(x)\}.$
+Let $\Phi=\min_{x}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x).$ Then from (2) and (7)
+$C_{\mathsf{OPT}}=\min_{x}\max\Big\{\sum_{t}f_{t}(x),\,\sum_{t}g_{t}(x)\Big\},\qquad{\bar{C}}_{\mathsf{OPT}}=T\Phi.$
+Let $\phi(v):=\min_{x}\max_{\theta\in\Delta_{2}}\theta^{\top}v(x)=\min_{x}\max\{v_{1}(x),v_{2}(x)\}$ for $v:[0,1]^{d}\to\mathbb{R}^{2}$.
+Lemma 12 (1–Lipschitz property of $\phi$).
+For any $v,w:[0,1]^{d}\to\mathbb{R}^{2}$,
+$\big|\phi(v)-\phi(w)\big|\;\leq\;\sup_{x\in[0,1]^{d}}\ \|v(x)-w(x)\|_{\infty}.$
+Proof.
+For each $x\in[0,1]^{d}$ define
+$h_{v}(x)\;:=\;\max_{\theta\in\Delta_{2}}\theta^{\top}v(x)\;=\;\|v(x)\|_{\infty}.$
+Then $\phi(v)=\min_{x}h_{v}(x)$.
+First, note the elementary inequality valid for any $a,b\in\mathbb{R}^{2}$:
+$\max_{k}a_{k}-\max_{k}b_{k}\;\leq\;\max_{k}(a_{k}-b_{k})\;=\;\|a-b\|_{\infty}.$
+Using ($\ast$) pointwise with $a=v(x)$ and $b=w(x)$ gives
+$h_{v}(x)-h_{w}(x)\;=\;\sup_{k}v_{k}(x)-\sup_{k}w_{k}(x)\;\leq\;\sup_{k}\big(v_{k}(x)-w_{k}(x)\big)\;=\;\|v(x)-w(x)\|_{\infty}.$
+Thus, we get
+$\phi(v)-\phi(w)\;=\;\inf_{x}h_{v}(x)-\inf_{x}h_{w}(x)\;\leq\;\sup_{x}\big(h_{v}(x)-h_{w}(x)\big)\;\leq\;\sup_{x}\|v(x)-w(x)\|_{\infty}.$
+Swapping the roles of $v$ and $w$ gives the reverse inequality $\phi(w)-\phi(v)\leq\sup_{x}\|v(x)-w(x)\|_{\infty}$. Combining the two displays proves
+$\big|\phi(v)-\phi(w)\big|\;\leq\;\sup_{x\in[0,1]^{d}}\ \|v(x)-w(x)\|_{\infty}.$
+Hence, from Lemma 12
+$\Big|\frac{C_{\mathsf{OPT}}}{T}-\Phi\Big|=|\phi(\hat{\mu}_{T})-\phi(\mu)|\leq\sup_{x\in[0,1]^{d}}\|\hat{\mu}_{T}(x)-\mu(x)\|_{\infty}.$ (14)
+Let $\mathcal{N}_{r}$ be an $\ell_{\infty}$ rr–net of $0,1]^{d} [$49], then $|\mathcal{N}_{r}|\leq(3/r)^{d}$ Lemma 5.2 [49]. Using the GG–Lipschitzness of $f_{t},g_{t}$, from Lemma 5.2 [49]
+$\sup_{x}\|\hat{\mu}_{T}(x)-\mu(x)\|_{\infty}\leq\max_{x\in\mathcal{N}_{r}}\|\hat{\mu}_{T}(x)-\mu(x)\|_{\infty}+2G\,r.$ (15)
+For each fixed $x\in\mathcal{N}_{r}$ and coordinate $j\in\{1,2\}$, Hoeffding’s inequality for [0,B][0,B]–bounded variables gives
+$\Pr\!\left(\big|\hat{\mu}_{T}^{(j)}(x)-\mu^{(j)}(x)\big|>\gamma\right)\leq 2\exp\!\left(-\frac{2T\gamma^{2}}{B^{2}}\right).$
+A union bound over $\mathcal{N}_{r}$ and the two coordinates yields
+$\Pr\!\left(\max_{x\in\mathcal{N}_{r}}\|\hat{\mu}_{T}(x)-\mu(x)\|_{\infty}>\gamma\right)\leq 4|\mathcal{N}_{r}|\exp\!\left(-\frac{2T\gamma^{2}}{B^{2}}\right).$ (16)
+Integrating the tail probability in (16) gives
+$\mathbb{E}\!\left\{\max_{x\in\mathcal{N}_{r}}\|\hat{\mu}_{T}(x)-\mu(x)\|_{\infty}\right\}$ $\leq B\sqrt{\frac{\log(4|\mathcal{N}_{r}|)}{2T}}+\frac{\sqrt{\pi}\,B}{2\sqrt{2T}}.$ (17)
+Combining (15) and (17) yields
+$\mathbb{E}\!\left\{\sup_{x}\|\hat{\mu}_{T}(x)-\mu(x)\|_{\infty}\right\}\leq B\sqrt{\frac{\log(4|\mathcal{N}_{r}|)}{2T}}+\frac{\sqrt{\pi}\,B}{2\sqrt{2T}}+2G\,r.$ (18)
+Take $r=T^{-1/2}$, so $|\mathcal{N}_{r}|\leq(3\sqrt{T})^{d}$ and $\log(4|\mathcal{N}_{r}|)\leq c_{0}d+\tfrac{d}{2}\log T$ with $c_{0}=\log 12$. Then (18) gives
+$\mathbb{E}\!\left\{\sup_{x}\|\hat{\mu}_{T}(x)-\mu(x)\|_{\infty}\right\}\leq C_{1}\,B\sqrt{\frac{d\log T}{T}}+C_{2}\frac{1}{\sqrt{T}}+C_{3}\,\frac{G}{\sqrt{T}},$
+for constants $C_{1},C_{2},C_{3}$.
+Taking expectations in (14) and multiplying by $T$,
+$\mathbb{E}\left\{\big|C_{\mathsf{OPT}}\}-T\Phi\big|\right\}\leq T\,\mathbb{E}\!\left\{\sup_{x}\|\hat{\mu}_{T}(x)-\mu(x)\|_{\infty}\right\}\leq C_{1}\,B\,\sqrt{d\,T\,\log T}+C_{2}\sqrt{T}+C_{3}\,G\,\sqrt{T}.$ (19)
 Remark 10.
-We note that only two loss function sequences are considered for simplicity of exposition since it captures the basic difficulty of the problem, and all the results that we derive apply directly to the case when there are $K$ function sequences (see details in Remark 7).
-4.1 Algorithm
-In this section, we present an algorithm to solve the i.i.d.
-min-max OCO problem (18).
-Towards that end, if suppose an online algorithm knew the optimizing $\theta^{\star}$ in (17), then by executing an OGD algorithm for OCO with a single function $\theta_{1}^{\star}f_{t}+\theta_{2}^{\star}g_{t}$ at time $t+1$, it would achieve a regret of $O(\sqrt{T})$ using standard results Hazan (2019).
-However, the optimizing $\theta^{\star}$ that depends on $\mathcal{D}$ (unknown) remains unknown, but, this observation motivates our algorithm that constructs
-i) $\theta_{t}$ at time $t$ as a surrogate for $\theta^{\star}$ using the Hedge algorithm (gains version) treating the two function sequences $\{f_{t}\}_{t=1}^{T}$ and $\{g_{t}\}_{t=1}^{T}$ as the gains of the two experts (see Section E for basic definition of the Experts Problem and the Hedge algorithm), and
-ii) for each fixed value of $\theta_{t}$, plays the action chosen by OGD applied to the single surrogate function $\theta_{t1}f_{t}+\theta_{t2}g_{t}$ at time $t+1$. Pseudocode is given in Algorithm 1.
-Algorithm 1 Minimization of the Maximum of The Sum of Two Convex Loss Sequences
-1:Input: step sizes $\eta_{x,t}>0$, $\eta_{\theta,t}>0$, feasible set $\mathcal{X}$
-2:Initialize: $w_{1}=(1,1)$, $\theta_{1}=w_{1}/\|w_{1}\|_{1}$, $f_{0}\equiv 0,g_{0}\equiv 0$ and play arbitrary $x_{0}\in\mathcal{X}$
-3:for $t=1,\dots,T$ do
-4:  Observe $(f_{t-1}(x_{t-1}),g_{t-1}(x_{t-1}))$ and $(\nabla f_{t-1}(x_{t-1}),\nabla g_{t-1}(x_{t-1}))$
-5:  Using current $\theta_{t}$, form convex function $x\mapsto{\tilde{\Lambda}}_{t}(x,\theta_{t}):=\theta_{t,1}f_{t-1}(x)+\theta_{t,2}g_{t-1}(x)$
-6:  OGD: Gradient step: $y_{t}=x_{t-1}-\eta_{x,t}\nabla{\tilde{\Lambda}}_{t}(x_{t-1},\theta_{t})$ and find $x_{t}=\text{Proj}_{\mathcal{X}}(y_{t})$
-7:  Play action $x_{t}$
-8:  Evaluate $\lambda_{t}=(f_{t}(x_{t}),g_{t}(x_{t}))$
-9:  Hedge: Update weights: $w_{t+1,i}=w_{t,i}\exp(\eta_{\theta,t}\lambda_{t,i})$ for $i=1,2$
-10:  Normalize: $\theta_{t+1}=w_{t+1}/\|w_{t+1}\|_{1}$
-11:end for
-4.2 Guarantee
-Theorem 4.
-Under Assumptions 1 and 2, for $\eta_{\theta,t}\leq\sqrt{\frac{2\ln 2}{B^{2}t}},\eta_{x,t}=\frac{D}{G\sqrt{t}}$, the expected min-max static regret (18) of the proposed algorithm (Algorithm 1) is
-${\bar{R}}_{\mathcal{A}}(T)=O(\sqrt{T}).$
-Proof.
-The main ingredient of the proof as well as the inspiration for Algorithm 1 is the following decomposition (19) of the expected min-max static regret (18). Recall the definition of
-$\Lambda_{t}(x,\theta):=\theta_{1}f_{t}(x)+\theta_{2}g_{t}(x).$
-where $\theta=(\theta_{1},\theta_{2})\in\Delta_{2}$.
-By adding and subtracting two terms, $\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t})$ and $\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta_{t})$ inside the expectation in the RHS of (18), we write the expected min-max static regret (18) as
-${\bar{R}}_{T}(\mathcal{A})=\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{1}\}+{\mathbb{E}}\{R_{2}\}+{\mathbb{E}}\{R_{3}\}\}\leq\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{1}\}\}+\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{2}\}\}+\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{3}\}\},$ (19)
-where
-$R_{1}=\max_{\theta\in\Delta_{2}}\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta)-\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t}),$
-$R_{2}=\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t})-\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta_{t}),$
-$R_{3}=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta_{t})-C_{\mathsf{OPT}}=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta_{t})-\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta).$
-The interpretation of $R_{1}$ is that it is the regret for the gains version of $2$-Experts Problem Cesa-Bianchi and Lugosi (2006), where the two function sequences $f_{t},g_{t}$ are the gains of two experts, with respect to an optimal $\theta^{\star}\in\Delta_{2}$ for fixed actions $x_{t}$ across time slots.
-This motivates the use of Hedge algorithm Cesa-Bianchi and Lugosi (2006) (gains version) in our proposed algorithm that is tailor-made for updating $\theta_{t}$ for fixed actions $x_{t}$.
-Similarly, $R_{2}$ is the static regret for an OCO minimization problem relative to static optimal action $x^{\star}$ for a single surrogate objective function sequence of $\Lambda_{t}(x,\theta_{t})=\theta_{t1}f_{t}(x)+\theta_{t2}g_{t}(x)$ for a fixed $\theta_{t}$.
-Hence using an OGD algorithm is natural for this purpose as done in the algorithm for each time $t$, with a fixed weight $\theta_{t}$.
-The third term $R_{3}$ is more nuanced and captures the difference between the cost of $\mathsf{OPT}$ for a fixed sequence of $\theta_{t},t=1,\dots,T$ and the optimal benchmark cost ${\bar{C}}_{\mathsf{OPT}}$ (17), respectively.
-We will bound $R_{1}$ and $R_{2}$ sample path wise for each realization of $f_{t},g_{t}$, while for $R_{3}$ we will bound ${\mathbb{E}}\{R_{3}\}$.
-In particular, we will show the following that is proven in the Appendix.
-Lemma 5.
-For $\eta_{\theta,t}\leq\sqrt{2\ln 2/(B^{2}t)}$, $\sup_{\mathcal{D}}\{R_{1}\}\leq O(\sqrt{T})$.
-Lemma 6.
-Choosing $\eta_{x,t}=\frac{D}{G\sqrt{t}}$, $\sup_{\mathcal{D}}\{R_{2}\}\leq O(\sqrt{T})$.
-Lemma 7.
-${\mathbb{E}}\{R_{3}\}\leq 0$.
-Because of the connection described above, the proof for $R_{1}$ in Lemma 5 and for $R_{2}$ in Lemma 6 follows standard analysis Hazan (2019) for Hedge and OGD, respectively.
-Bounding $R_{3}$ needs more care and we prove it in Appendix C where we critically exploit the i.i.d.
-input model.
-Combining Lemma 5, 6 and 7, and the regret decomposition (19), we get that the expected min-max static regret (12) of Algorithm 1 with
-$\eta_{\theta,t}\leq\sqrt{\frac{2\ln 2}{B^{2}t}},\qquad\ \text{and}\qquad\eta_{x,t}=\frac{D}{G\sqrt{t}}$ (20)
-is
-${\bar{R}}_{\mathcal{A}}(T)=O(\sqrt{T}),$ (21)
-proving the result. ∎
-Following remarks are in order.
-Remark 5.
-The regret decomposition (19) does multiple things at the same time i) identifies the critical difficulty of the problem, ii) motivates a natural algorithm, and ii) keeps the proof modular, where the regret bound for each module is simple and elegant, almost first principles based.
-Remark 6.
-Although the proof of Theorem 4 is remarkably simple and elegant, it hinges on the critical regret decomposition (19) – which is not obvious a priori.
-This decomposition eliminates the need for sophisticated technical machinery needed in related prior work Mahdavi et al.
-(2013) that also considers the i.i.d.
-input and highlights the underlying simplicity of the result.
-We would like to emphasize that the elegance of the argument should not obscure its importance.
-Remark 7.
-When there are $K$ loss function sequences $f_{t}^{k},k=1,\dots,K$ that are chosen i.i.d.
-from $\mathcal{D}$, define $\theta\in\Delta_{K}$, (where $\Delta_{K}$ is the $K$-dimensional simplex) and $\Lambda_{t}(x,\theta)=\sum_{k=1}^{K}\theta_{k}f_{t}^{k}$.
-Then the expected min-max static regret of $\mathcal{A}$ is
-${\bar{R}}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}{\mathbb{E}}\left\{\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta)-T\cdot\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{K}}\theta^{\top}\mu(x)\right\}$ (22)
-where $\mu(x)={\mathbb{E}}\{f_{t}^{1}(x),\dots,f_{t}^{K}(x)\}$.
-With this new definition of $\theta$ and $\Lambda_{t}$, Algorithm 1 remains as is with Hedge choosing $\theta_{t}\in\Delta_{K}$ while $x_{t}$ the single action chosen by OGD.
-Assuming all $f_{t}^{k},k=1,\dots,K$ satisfy Assumption 1 and 2, the only change in the guarantee with $K>2$ compared to Theorem 4 with $K=2$ is because of Lemma 5 that is now $O(\sqrt{T\log K})$ Hazan (2019) since there are $K$ experts corresponding to the $K$ loss function sequences $f_{t}^{k},k=1,\dots,K$ .
-The OGD guarantee of Lemma 6 stays the same since $\|\nabla\Lambda_{t}(x_{t},\theta)\|\leq G$ as before.
-Moreover, Lemma 7 holds as is.
-Thus, the overall expected min-max static regret of Algorithm 1 with $K$ loss function sequences is
-$O(\sqrt{T\log K}).$ (23)
-Remark 8.
-We note that Algorithm 1 is well motivated even while considering the worst case/adversarial input.
-The technical problem in its analysis with the worst case/adversarial input is that Lemma 7 that is the critical result fails when considering the adversarial input.
-We highlight this difficulty in Appendix F and show that the ‘best possible’ upper bound on $R_{3}$ is $R_{3}=O(T^{3/2})$ with the worst case/adversarial input for Algorithm 1 by exploiting just the properties of the Hedge algorithm which is insufficient for achieving sub-linear regret.
-Remark 9.
-[Extension to Strongly Convex $f_{t}$’s and $g_{t}$’s] One limitation of our proposed algorithm and the proof is that because of the use of the Hedge algorithm for updating $\theta_{t}$, the regret $\sup_{\mathcal{D}}\{R_{1}\}$ can at best be $O(\sqrt{T})$.
-Thus, unlike OCO, where OGD algorithm achieves logarithmic regret in $T$ when $f_{t}$’s are strongly convex, our algorithm’s regret is only $O(\sqrt{T})$ even when both $f_{t},g_{t}$ are strongly convex.
-Is this a limitation of the algorithm or of that of the problem itself remains an interesting unresolved question.
-5 Improving the result on GlobalExpertsProblem with the i.i.d. input Even-Dar et al. (2009)
-Let $\mathcal{X}=\Delta_{K}$, the $K$-dimensional simplex, and writing $f_{t}^{k}(x)=x_{k}\ell_{k}(t)$, where $x_{k}$ is the $k$-th coordinate of $x\in\mathcal{X}$, $\ell_{k}(t)$ is the loss corresponding to the $k$-th expert, and $k=1,2\dots,K$, we get the GlobalExpertsProblem Even-Dar et al.
-(2009) with $K$ experts in the i.i.d.
-input case is a special case of the i.i.d.
-min-max OCO problem (18).
-Thus, from (23), we get that the expected min-max static regret of Algorithm 1 is $O(\sqrt{T\log K})$ for the GlobalExpertsProblem with i.i.d.
-input.
-In comparison, Even-Dar et al.
-(2009) derived a min-max static regret (5) guarantee of $O(\sqrt{T}\log K)$ with adversarial input for the following algorithm called MULTI:
-For $K=2$,
-$x_{1}(t+1)=x_{1}(t)+\frac{x_{2}(t)\ell_{2}(t)-x_{1}(t)\ell_{1}(t)}{\sqrt{T}},$
-and $x_{2}(t)=1-x_{1}(t)$.
-For general $K$, the $K=2$ algorithm was used recursively.
-Compared to our Algorithm 1, the MULTI algorithm Even-Dar et al.
-(2009) is very compute intensive since it recursively uses an algorithm for $K=2$ and requires the knowledge of time horizon $T$.
-6 Extension to Bandit Feedback
-We now analyze the expected min-max OCO problem (18) under the bandit feedback, where at time $t$, once the online algorithm $\mathcal{A}$ chooses its action $x_{t}$, it only observes the scalar values $\{f_{t-1}^{k}(.)\}_{k=1}^{K}$ at one (one-point model) or two points (two-point model) and does not receive any gradient information.
-Algorithm 2 Bandit-MinMax (unified one-point/two-point variant)
-1:Feasible Set $\mathcal{X}\subset\mathbb{R}^{d}$, smoothing parameters $\{\delta_{t}\}$, step-sizes $\{\eta_{x,t}\},\{\eta_{\theta,t}\}$, feedback type $\mathsf{mode}\in\{\text{one-point},\text{two-point}\}$
-2:Initialize $w_{1}=\mathbf{1}_{K}$, $\theta_{1}=w_{1}/\|w_{1}\|_{1}$, choose $x_{0}\in X$
-3:Define $\Lambda_{t}(x,\theta_{t})=\sum_{k}\theta_{t,k}f_{t}^{k}(x)$
-4:for $t=1,\dots,T$ do
-5:  Sample $u_{t}$ uniformly from unit sphere $\mathbb{S}^{d-1}$
-6:  if $\mathsf{mode}=\text{one-point}$ then
-7:   Play $\tilde{x}_{t}=\text{Proj}_{\mathcal{X}}(x_{t-1}+\delta_{t}u_{t})$ and observe $\{f_{t}^{k}(\tilde{x}_{t})\}_{k=1}^{K}$
-8:   Gradient estimator:
-$\widehat{\nabla}_{t}=\frac{d}{\delta_{t}}\Lambda_{t}(\tilde{x}_{t},\theta_{t})\,u_{t}$
-9:  else if $\mathsf{mode}=\text{two-point}$ then
-10:   Query $x_{t-1}+\delta_{t}u_{t}$ and $x_{t-1}-\delta_{t}u_{t}$, observe $\{f_{t}^{k}(x_{t-1}\pm\delta_{t}u_{t})\}_{k=1}^{K}$
-11:   Gradient estimator:
-$\widehat{\nabla}_{t}=\frac{d}{2\delta_{t}}\Big(\Lambda_{t}(x_{t-1}+\delta_{t}u_{t},\theta_{t})-\Lambda_{t}(x_{t-1}-\delta_{t}u_{t},\theta_{t})\Big)u_{t}$
-12:  end if
-13:  Gradient step: $y_{t}=x_{t-1}-\eta_{x,t}\widehat{\nabla}_{t}$,    $x_{t}=\text{Proj}_{\mathcal{X}}(y_{t})$
-14:  Hedge update: $w_{t+1,k}=w_{t,k}\exp\big(\eta_{\theta,t}f_{t,k}(x_{t})\big)$,    $\theta_{t+1}=w_{t+1}/\|w_{t+1}\|_{1}$
-15:end for
-6.1 Algorithm
-For this bandit model, we consider a simple adaption of Algorithm 1 by replacing its OGD part with the bandit information counterpart, while keeping the Hedge part as it is, given by Algorithm 2.
-Moreover, the analysis also follows the same script by replacing the guarantees by their respective bandit counterparts.
-6.2 Guarantee
-Theorem 8.
-Under Assumption 1 and 2,
-  1. 1.
-    (One-point feedback) With $\delta_{t}=t^{-1/4}$, $\eta_{x,t}=D/(Gd^{1/2}t^{3/4})$, and $\eta_{\theta,t}=\sqrt{\log K}/(B\sqrt{t})$, the expected min-max static regret satisfies
-    ${\bar{R}}_{\mathcal{A}}(T)\leq c_{1}\,d^{1/2}GD\,T^{3/4}\;+\;c_{2}B\sqrt{T\log K},$
-    for universal constants $c_{1},c_{2}>0$.
-  2. 2.
-    (Two-point feedback) With $\delta_{t}=t^{-1/2}$, $\eta_{x,t}=D/(Gd^{1/2}\sqrt{t})$, and $\eta_{\theta,t}=\sqrt{\log K}/(B\sqrt{t})$, the expected min-max static regret satisfies
-    ${\bar{R}}_{\mathcal{A}}(T)\leq c_{3}\,d^{1/2}GD\,\sqrt{T}\;+\;c_{4}B\sqrt{T\log K},$
-    for universal constants $c_{3},c_{4}>0$.
-6.3 Proof of Theorem 8
-Recall the decomposition (19)
-${\bar{R}}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{1}\}+{\mathbb{E}}\{R_{2}\}+{\mathbb{E}}\{R_{3}\}\}\leq\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{1}\}\}+\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{2}\}\}+\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{3}\}\}.$
-Bounding $R_{1}$ and $R_{3}$.
-For both one-point or two-point bandit feedback, the Hedge guarantee remains the same as in Lemma 5. In particular, with $\eta_{\theta,t}=\sqrt{\log K}/(B\sqrt{t})$, sample path wise,
-$R_{1}\leq B\sqrt{2T\log K},$
-and similar to Lemma 7,
-${\mathbb{E}}\{R_{3}\}\leq 0.$
-The only change is in bounding $R_{2}$.
-Bounding $R_{2}$ (one-point feedback). The one-point estimator
-$\widehat{\nabla}_{t}=\frac{d}{\delta_{t}}\Lambda_{t}(\tilde{x}_{t},\theta_{t})\,u_{t}$
-is unbiased for the gradient of the $\delta_{t}$-smoothed loss, with variance at most $O(dB^{2}/\delta_{t}^{2})$.
-Using $\delta_{t}=t^{-1/4}$ and $\eta_{x,t}=D/(Gd^{1/2}t^{3/4})$, the bandit-OGD analysis Flaxman et al.
-(2005) yields
-$R_{2}\leq c_{1}\,d^{1/2}GD\,T^{3/4},$
-for some universal constant $c_{1}$.
-Combining the three bounds we get
-${\bar{R}}_{\mathcal{A}}(T)\leq c_{1}\,d^{1/2}GD\,T^{3/4}\;+\;2B\sqrt{2T\log K}.$
-Bounding $R_{2}$ (two-point feedback). With two function evaluations per round, the estimator
-$\widehat{\nabla}_{t}=\tfrac{d}{2\delta_{t}}\Big(\Lambda_{t}(x_{t-1}+\delta_{t}u_{t},\theta_{t})-\Lambda_{t}(x_{t-1}-\delta_{t}u_{t},\theta_{t})\Big)u_{t}$
-is unbiased with variance $O(dB^{2}/\delta_{t}^{2})$.
-Choosing $\delta_{t}=t^{-1/2}$ and $\eta_{x,t}=D/(Gd^{1/2}\sqrt{t})$, the two-point bandit-OGD analysis Agarwal et al.
-(2010); Duchi et al.
-(2015); Shamir (2017) gives
-$R_{2}\leq c_{3}\,d^{1/2}GD\,\sqrt{T},$
-for some universal constant $c_{3}$. Thus, with two-point feedback,
-${\bar{R}}_{\mathcal{A}}(T)\leq c_{3}\,d^{1/2}GD\,\sqrt{T}\;+\;2B\sqrt{2T\log K}.$
-Discussion: Compared to the limited information setting considered in Section 2, where gradient information is available, with bandit feedback the regret guarantees suffer for the expected min-max OCO problem, however, the degraded guarantees match the best known guarantees for OGD in the bandit setting.
-The proof of the result is simple because of the modularity of our algorithm and the regret decomposition that decomposes the total regret into regret resulting from Hedge and OGD, and the fact that only regret for OGD changes with bandit feedback.
-7 Experimental Results
-In this section, we present numerical results on the expected min-max static regret of Algorithm 1 under relevant settings, that are well motivated from both theoretical as well applications point of view.
-In all the figures, we plot the expected min-max static regret as a function of time horizon $T$.
-7.1 Basic Results
-We begin our experimental results by considering the input model when all functions $f_{t}^{k}$ for $k=1,\dots,K,t=1,\dots,T$ are random linear functions in $d=10$ dimensions, where each coefficient is distributed uniformly between $[0,1]$.
-Random linear functions are important since they were sufficient to find lower bounds on the regret of any online algorithm for OCO Hazan (2019).
-In Fig.
-1, we plot the expected min-max static regret of Algorithm 1 as a function of $K$ and $T$ for random linear functions, where as expected, we observe that the regret scales as $O(\sqrt{T\log K})$.
-Next, we consider $f_{t}^{k}$’s as strongly convex functions, where each $f_{t}^{k}$ is independently and identically selected from a set of $\{(x-a)^{2},a=[-9,-8,\dots,0,1,\dots,9]\}$.
-Recall that for OCO with strongly convex functions, the step size $\eta$ needed for OGD to get a regret guarantee of $O(\log T)$ scales as $\eta=1/t$ and not $\eta=1/\sqrt{t}$ Hazan (2019).
-Thus, for this simulation of Algorithm 1333In contrast to $\eta_{x,t}=O(1/\sqrt{t})$ as needed to prove Theorem 4.
-we use $\eta_{\theta,t}=O(1/\sqrt{t})$ and $\eta_{x,t}=O(1/t)$.
-[^4] We have checked that other combinations only increase the regret.
-As pointed out earlier, we are unable to improve the expected min-max static regret guarantee for Algorithm 1 when all functions $f_{t}^{k}$’s are strongly convex.
-We see that in fact that is true for Algorithm 1 even in simulations as shown in Fig.
-2, where for $K=1$, which is the same as OCO (with $\eta_{x,t}=O(1/t)$) the regret scales as $\log(T)$, while for larger values of $K$, the expected min-max static regret scales as $O(\sqrt{T})$ when $\eta_{\theta,t}=O(1/\sqrt{t})$ and $\eta_{x,t}=O(1/t)$.
-Thus, if indeed expected min-max static regret of $O(\log T)$ is achievable then Algorithm 1 needs to be conceptually modified.
-050501001001501502002002502503003000202040406060Time ($t$)Average Cumulative RegretExpected min-max static regret for random linear functionsK=1K=1K=10K=10K=50K=50K=100K=100
-Figure 1: Expected min-max static regret of Algorithm 1 when $f_{t}^{k}$ are random linear functions for different values of $K$.
-0505010010015015020020025025030030002002004004006006008008001,0001{,} [^0] Time Horizon ($T$)Average RegretExpected min-max static regret for quadratic loss functionsK=1K=1K=10K=10K=50K=50K=100K=100
-Figure 2: Expected min-max static regret of Algorithm 1 when $f_{t}^{k}$ are random quadratic functions for different values of $K$.
-7.2 GlobalExpertsProblem
-In this section, we compare the performance of Algorithm 1 and MULTI Even-Dar et al.
-(2009) for the GlobalExpertsProblem with the i.i.d.
-input.
-For that purpose, we consider that the loss of each expert at time $t$ is drawn independently from a uniform distribution:
-$\ell_{t,k}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathsf{Uniform}(a,b),\quad t=1,2,\dots,T,\;k=1,2,\dots,K,$
-where $a,b\in[0,1]$ are fixed parameters defining the range of losses. In our experiments, we set $a=0.2$ and $b=0.8$ to ensure non-trivial loss variation while avoiding extreme loss values.
-For this model, we plot the expected min-max static regret performance of Algorithm 1 and MULTI in Fig.
-3, and see that both perform similarly, even though the regret guarantee we obtained for Algorithm 1 in i.i.d.
-case is slightly better than MULTI (which is valid even with the adversarial input).
-It appears that MULTI has better performance than its guarantee Even-Dar et al.
-(2009).
-However, MULTI is computationally far more intensive than Algorithm 1 since it uses the $K=2$ primitive recursively, and moreover, it needs to know the time horizon $T$ in contrast to Algorithm 1.
-1,0001{,}0002,0002{,}0004,0004{,}0008,0008{,}0000551010151520202525303035354040K=2K=2K=4K=4K=16K=16K=32K=32K=64K=64Time horizon TTAverage RegretMULTIAlgorithm 1
-Figure 3: Expected min-max static regret of MULTI Even-Dar et al. (2009) and Algorithm 1 for different numbers of experts $K$ and time horizons $T$ for the fair classification example.
-7.3 Application of min-max formulation to Fair Classification Setting
-Finally, in this subsection, we simulate the i.i.d. min-max OCO problem for a fair classification setting.
-Data generation.
-We consider $K$ groups, each corresponding to a convex loss function sequence.
-For each group $k\in[K]$, we sample a hidden parameter $w_{k}^{\star}\sim\mathcal{N}(0,I_{d})$.
-At each round $t=1,\dots,T$, for each group $k$ we draw $m$ feature vectors
-$z_{t,k,i}\sim\mathcal{N}(\mu_{k},\sigma^{2}I_{d}),\quad i=1,\dots,m,$
-with $\mu_{k}$ drawn once per group. Labels are generated as
-$y_{t,k,i}\sim\mathrm{Bernoulli}\!\left(\sigma(\langle w_{k}^{\star},z_{t,k,i}\rangle)\right),$
-where $\sigma(u)=1/(1+e^{-u})$ is the sigmoid function.
-Loss functions.
-The per-group convex loss at round $t$ is the averaged logistic loss with $L_{2}$ regularization:
-$f_{t}^{k}(x)\;=\;\frac{1}{m}\sum_{i=1}^{m}\log\!\Big(1+\exp(-y_{t,k,i}\,\langle x,z_{t,k,i}\rangle)\Big)\;+\;\kappa\|x\|_{2}^{2}.$
-For this application, we compare the performance of following algorithms.
-  • Algorithm 1 (OGD+Hedge)
-  • OGD on average loss: projected OGD on $\tfrac{1}{K}\sum_{k=1}^{K}f_{t}^{k}(x)$.
-  • Follow-the-Regularized-Leader (FTRL): $x_{t}=\arg\min_{x\in\mathcal{X}}\left\{\max_{k}\sum_{s=1}^{t-1}f_{s}^{k}(x)+\mathcal{R}(x)\right\}$.
-Parameters.
-We use $d=20$, $K=10$, mini-batch size $m=50$, diameter $D=5$, and regularization $\kappa=10^{-3}$.
-Step sizes are set as $\eta_{x,t}=D/(G\sqrt{t})$ for OGD (with $G$ an upper bound on gradient norms) and $\eta_{\theta,t}=\sqrt{\ln K}/(B\sqrt{t})$ for Hedge, with $B$ an upper bound on per-round losses.
-In Fig.
-4, we plot the expected min-max static regret ${\bar{R}}_{T}$ for all the three algorithms.
-We see that all algorithms have sub-linear regret, and Algorithm 1 has larger regret than FTRL but better than average OGD.
-However, recall that neither FTRL and average OGD have any theoretical guarantees on their expected min-max static regret.
-Next, in Fig.
-5, we will consider an input where the index of the most challenging sequence changes over time and show that Algorithm 1 outperforms FTRL in that case.
-10010020020030030040040002020404060608080Time Horizon ($T$)Average Regret vs.
-Offline OptimalExpected min-max static regret for $K=10$ for the fair classification problemAlgorithm 1 OGD+HedgeAvg-OGDFTL
-Figure 4: Expected min-max static regret comparison for $K=10$.
-Non-stationary input
-Next, we change the input described in Section 7.3 to model a highly non‑stationary environment specifically designed to challenge static optimization strategies such as FTRL and highlight the adaptability of Algorithm 1 that uses OGD+Hedge.
-The detailed changes with respect to Section 7.3 are as follows:
-  • We consider $K=3$ separate objectives, each with its own logistic loss function defined by fixed but distinct parameter vectors $w_{\text{star}}[k]$ and mean feature shifts $\mu[k]$.
-  • Every switch_interval steps (here set to 100), one objective becomes significantly harder by adding a large mean shift ($\texttt{shift\_magnitude}=5.0$) to its feature distribution.
-This creates abrupt changes in which objective dominates the cumulative loss objective function.
-  • Data generation: For each round $t$, and for each objective $k$, a feature matrix $Z$ of size $m\times d$ is generated from a Gaussian distribution centered at the shifted mean.
-Labels $y$ are generated from a logistic model using $w_{\text{star}}[k]$ (is a hidden “true” parameter vector for objective $k$.
-It defines how that objective behaves.), with additional noise introduced via random sampling.
-We clearly see in Fig. 5 that Algorithm 1 (OGD+Hedge) adapts better to changing conditions than FTRL, since worst‑case objective changes over time, making past loss information misleading for FTRL.
-1010505010010030030050050002020404060608080Time Horizon ($T$)Average Regret (Avg. over 10 Seeds)Expected min-max static regret in switching environment for K=3K=3Algorithm 1 OGD+HedgeFTRL
-Figure 5: Expected min-max static regret for Algorithm 1 vs. FTRL in a piecewise-stationary environment.
-8 Conclusions
-In this paper, we have made a basic advance in the area of multi-objective OCO by formulating the min-max static regret minimization problem, that requires an algorithm to closely track multiple loss function sequences simultaneously and at all times.
-Because of analytical challenges with the adversarial input, we considered the i.i.d.
-input setting, for which we proposed a simple algorithm that used a combination of the well known Hedge and OGD algorithm.
-Using an insightful regret decomposition, which made the analysis modular, we derived a $O(\sqrt{T\log K})$ regret bound on the proposed algorithm, which is shown to be best possible via simulations.
-We also extended the results to the bandit information setting.
-Certain questions that remain open are: i) what is the best regret possible with the adversarial input, ii) how to handle the case when all loss functions are strongly convex, and iii) handling constraints like COCO.
-Appendix A Proof of Lemma 5
+If there are $K$ functions $f_{t}^{k}$, $k=1,\dots,K$, instead of just two $f_{t},g_{t}$, the bound (19) changes by an additive term $O(\sqrt{T\log K})$, resulting in the final bound of $O(\sqrt{T\log(TK)})$.
+Appendix E Proof of Lemma 6
 Let $S_{1}:=\sum_{t=1}^{T}f_{t}(x_{t})$ and $S_{2}:=\sum_{t=1}^{T}g_{t}(x_{t})$, and recall that $\lambda_{t}=(f_{t}(x_{t}),g_{t}(x_{t}))$ and $\sum_{t}\Lambda_{t}(x_{t},\theta_{t})=\theta_{t1}f_{t}+\theta_{t2}g_{t}=\sum_{t}\langle\theta_{t},\lambda_{t}\rangle$.
@@
 $R_{1}=\max_{\theta\in\Delta_{2}}\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta)-\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t})=\max\{S_{1},S_{2}\}-\sum_{t=1}^{T}\langle\theta_{t},\lambda_{t}\rangle\leq\frac{\ln 2}{\eta_{\theta,T}}+\sum_{t=1}^{T}\frac{\eta_{\theta,t}B^{2}}{2}.$
-Appendix B Proof of Lemma 6
+Appendix F Proof of Lemma 7
 Define $\tilde{f}_{t}(x):=\Lambda_{t}(x,\theta_{t})$, which is convex for fixed $\theta_{t}$. Thus, the standard OGD (projected) guarantee from [20] implies, for any $x\in\mathcal{X}$,
@@
 $R_{2}\leq DG\sqrt{T}.$
-Appendix C Proof of Lemma 7
+Appendix G Proof of Lemma 8
 Recall that
@@
 By the definition of $R_{3}$,
-$R_{3}$ $=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\theta_{t}^{\top}\lambda_{t}(x)-\min_{x\in\mathcal{X}}\max_{\theta}\sum_{t=1}^{T}\theta^{\top}\lambda_{t}(x)$
-$\stackrel{{\scriptstyle(a)}}{{=}}\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\theta_{t}^{\top}\lambda_{t}(x)-\sum_{t=1}^{T}\theta^{\star}{}^{\top}\lambda_{t}(x^{\star}),$
-$\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{t=1}^{T}\theta_{t}^{\top}\lambda_{t}(x^{\star})-\sum_{t=1}^{T}\theta^{\star}{}^{\top}\lambda_{t}(x^{\star}),$
-$=\sum_{t=1}^{T}(\theta_{t}-\theta^{\star})^{\top}\lambda_{t}(x^{\star}),$ (25)
-where $(a)$ follows from the definition of $x^{\star}$ and $(b)$ follows since $x^{\star}\in\mathcal{X}$ is feasible. Moreover, we have
-$\mathbb{E}\left\{(\theta_{t}-\theta^{\star})^{\top}\lambda_{t}(x^{\star})\,\Big|\,\mathcal{F}_{t-1}\right\}=(\theta_{t}-\theta^{\star})^{\top}{\mathbb{E}}\{\lambda_{t}(x^{\star})|\mathcal{F}_{t-1}\}=(\theta_{t}-\theta^{\star})^{\top}\mu(x^{\star}),$ (26)
-where the first equality follows since $\theta_{t}$ is completely determined by $\mathcal{F}_{t-1}$, while the second equality follows because ${\mathbb{E}}\{\lambda_{t}(x^{\star})|\mathcal{F}_{t-1}\}=\mu(x^{\star})$ from the i.i.d.
-assumption.
+$R_{3}$ $=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\theta_{t}^{\top}\lambda_{t}(x)-T\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x)$
+$=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\theta_{t}^{\top}\lambda_{t}(x)-T\,\theta^{\star}{}^{\top}\mu(x^{\star}),$
+$\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{t=1}^{T}\theta_{t}^{\top}\lambda_{t}(x^{\star})-T\,\theta^{\star}{}^{\top}\mu(x^{\star}),$ (21)
+where $(b)$ follows since $x^{\star}\in\mathcal{X}$ is feasible.
+Now rewrite the right-hand side of (21) as
+$\sum_{t=1}^{T}\theta_{t}^{\top}\lambda_{t}(x^{\star})-T\,\theta^{\star}{}^{\top}\mu(x^{\star})$ $=\sum_{t=1}^{T}(\theta_{t}-\theta^{\star})^{\top}\lambda_{t}(x^{\star})\;+\;\sum_{t=1}^{T}\theta^{\star}{}^{\top}\big(\lambda_{t}(x^{\star})-\mu(x^{\star})\big).$ (22)
+Moreover, we have
+$\mathbb{E}\left\{(\theta_{t}-\theta^{\star})^{\top}\lambda_{t}(x^{\star})\,\Big|\,\mathcal{F}_{t-1}\right\}=(\theta_{t}-\theta^{\star})^{\top}{\mathbb{E}}\{\lambda_{t}(x^{\star})\mid\mathcal{F}_{t-1}\}=(\theta_{t}-\theta^{\star})^{\top}\mu(x^{\star}),$ (23)
+where the first equality follows since $\theta_{t}$ is completely determined by $\mathcal{F}_{t-1}$, while the second equality follows from the i.i.d. assumption.
 We also have
 $(\theta_{t}-\theta^{\star})^{\top}\mu(x^{\star})\leq 0$
-for each $t=1,\dots,T$ using the optimality of $\theta^{\star}$ (24) since
-$\theta^{\star}\in\arg\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x^{\star})$
-depends only on the expected vector $\mu(x^{\star})$ and not on any particular realization of $\lambda_{t}(x^{\star})$.
-Thus, we get from (26)
-$\mathbb{E}\left\{(\theta_{t}-\theta^{\star})^{\top}\lambda_{t}(x^{\star})\,\Big|\,\mathcal{F}_{t-1}\right\}\leq 0,$ (27)
-Using (27) in (25) and taking the expectation (using the tower property of conditional expectation),
-$\mathbb{E}\{R_{3}\}\leq\sum_{t=1}^{T}(\theta_{t}-\theta^{\star})^{\top}\mu(x^{\star})\leq 0.$ (28)
-Appendix D Per-slot $\it{\it min-max}$} multi-objective optimization [24]
+for each $t=1,\dots,T$, by the optimality of $\theta^{\star}$ in
+$\theta^{\star}\in\arg\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x^{\star}).$
+Further, note that
+${\mathbb{E}}\!\left\{\theta^{\star}{}^{\top}\big(\lambda_{t}(x^{\star})-\mu(x^{\star})\big)\,\Big|\,\mathcal{F}_{t-1}\right\}=0,$
+again by the i.i.d. assumption.
+Combining the above and taking conditional expectations in (22), we obtain
+${\mathbb{E}}\!\left\{\sum_{t=1}^{T}\theta_{t}^{\top}\lambda_{t}(x^{\star})-T\,\theta^{\star}{}^{\top}\mu(x^{\star})\right\}\leq 0.$ (24)
+Hence,
+$\mathbb{E}\{R_{3}\}\leq 0.$ (25)
+Appendix H Per-slot $\it{\it min-max}$} multi-objective optimization [25]
 In this section, we show that benchmark $W_{L}^{T}$ (5) can be $3/2$ times more than our benchmark $C_{\mathsf{OPT}}$ (2).
@@
 Therefore, even if an online algorithm $\mathcal{A}$ achieves sub-linear regret with respect to the benchmark $W_{L}^{T}$(5), it does not necessarily have sub-linear regret for our benchmark $C_{\mathsf{OPT}}$ (2).
-Appendix E Experts Problem (Gains Version)
+Remark 11.
+Detailed comparison with $\alpha$-fair model of [40] The transition from $\alpha$-fairness (with finite $\alpha$) (6) to the min-max objective (1) marks a shift from a smooth, strictly concave optimization landscape to a non-smooth, polyhedral one.
+For any finite $\alpha\geq 0$, the $\alpha$-fair utility function $U_{\alpha}(z)$ (6) acts as a differentiable, strictly concave regularizer that “softens” the trade-offs between competing objectives.
+In the finite $\alpha$ regime, the gradient of the aggregate utility changes continuously with the input, allowing algorithms to rely on the local curvature of the utility function to guide the learning process.
+However, as $\alpha\to\infty$, the utility function converges to the min-max objective, which is equivalent to an $L_{\infty}$ norm.
+Geometrically, the smooth, rounded contours of the $\alpha$-fair utility collapse into a sharp wedge where the derivative is discontinuous at the points where multiple objectives have equal costs.
+This loss of differentiability explains why prior regret bounds of $O(\alpha\sqrt{T})$ [40] become vacuous in the min-max limit; they are inversely proportional to the curvature of the utility, which vanishes as the objective becomes piecewise linear.
+In comparison, the Hedge+OGD algorithmic approach succeeds as $\alpha\rightarrow\infty$ because it does not attempt to “smoothen” this non-differentiable cusp.
+Instead, by employing the Hedge algorithm to update weights $\theta_{t}$, it treats the selection of the dominant objective as a discrete experts problem.
+This effectively bypasses the numerical instability inherent in the $\alpha\to\infty$ transition and provides a regret guarantee that remains independent of the parameter $\alpha$.
+Hedge+OGD algorithm is also applicable for finite $\alpha$, however, since it is designed to remain stable in the non-smooth limit ($\alpha\to\infty$), it lacks the specialized mechanics required to exploit the curvature present when $\alpha$ is small.
+For finite $\alpha$, the $\alpha$-fair utility function is strictly concave, offering a smooth landscape for optimization.
+The OHF algorithm [40] is specifically tailored to exploit this strict concavity, allowing the algorithm to achieve a regret bound of $O(\alpha\sqrt{T})$.
+When $\alpha$ is small, the utility function possesses significant curvature, which provides a strong, directional gradient signal that accelerates convergence.
+Hedge on the other hand is a first-order expert algorithm designed for the simplex; it does not naturally incorporate the second-order information or the strong concavity of the underlying utility.
+Consequently, it cannot “speed up” when the landscape is smooth.
+The main takeaway is that the finite $\alpha$ and $\alpha\rightarrow\infty$ regimes are fundamentally different and require different treatment.
+Appendix I Proof of Theorem 9
+Recall the decomposition (10),
+${\bar{R}}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{1}\}+{\mathbb{E}}\{R_{2}\}+{\mathbb{E}}\{R_{3}\}\}\;\leq\;\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{1}\}\}+\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{2}\}\}+\sup_{\mathcal{D}}\{{\mathbb{E}}\{R_{3}\}\}.$
+Bounding $R_{1}$ and $R_{3}$.
+Recall that
+$R_{1}:=\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta)-\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t}),\qquad\Lambda_{t}(x,\theta)=\sum_{k=1}^{K}\theta_{k}f_{t}^{k}(x).$
+In the bandit setting, the Hedge update observes the queried losses $\{f_{t-1}^{k}(\tilde{x}_{t})\}_{k=1}^{K}$. To relate $R_{1}$ to the Hedge regret on the queried points, we proceed as follows.
+Step 1 (add and subtract queried losses inside the maximization). For any fixed $\theta\in\Delta_{K}$,
+$\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta)=\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta)+\sum_{t=1}^{T}\bigl(\Lambda_{t}(x_{t},\theta)-\Lambda_{t}(\tilde{x}_{t},\theta)\bigr).$
+Substituting this identity into the definition of $R_{1}$ yields
+$R_{1}$ $=\max_{\theta\in\Delta_{K}}\Bigg[\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta)-\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t})$
+$\hskip 85.35826pt+\sum_{t=1}^{T}\bigl(\Lambda_{t}(x_{t},\theta)-\Lambda_{t}(\tilde{x}_{t},\theta)\bigr)\Bigg].$
+Step 2 (split the maximization). Using the elementary inequality
+$\max_{\theta}(a_{\theta}+b_{\theta})\;\leq\;\max_{\theta}a_{\theta}+\max_{\theta}b_{\theta},$
+we obtain
+$R_{1}$ $\leq\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta)-\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t})$
+$\quad+\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\bigl(\Lambda_{t}(x_{t},\theta)-\Lambda_{t}(\tilde{x}_{t},\theta)\bigr).$
+Step 3 (Lipschitz bound, uniform in $\theta$). By Assumption 3, each $f_{t}^{k}$ is $G$-Lipschitz, and since $\theta\in\Delta_{K}$ is a convex combination,
+$\bigl|\Lambda_{t}(x_{t},\theta)-\Lambda_{t}(\tilde{x}_{t},\theta)\bigr|\leq G\|x_{t}-\tilde{x}_{t}\|,\qquad\forall\,\theta\in\Delta_{K}.$
+Therefore,
+$\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\bigl(\Lambda_{t}(x_{t},\theta)-\Lambda_{t}(\tilde{x}_{t},\theta)\bigr)\leq G\sum_{t=1}^{T}\|x_{t}-\tilde{x}_{t}\|.$
+Combining the above bounds gives
+$R_{1}\leq\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta)-\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t})+G\sum_{t=1}^{T}\|x_{t}-\tilde{x}_{t}\|.$
+Step 4 (remove the remaining mismatch outside the maximization). Adding and subtracting $\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta_{t})$ yields
+$R_{1}$ $\leq\Bigg[\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta)-\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta_{t})\Bigg]$
+$\quad+\Bigg[\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta_{t})-\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t})\Bigg]+G\sum_{t=1}^{T}\|x_{t}-\tilde{x}_{t}\|.$
+The second bracket is again bounded by Lipschitz continuity:
+$\bigl|\Lambda_{t}(\tilde{x}_{t},\theta_{t})-\Lambda_{t}(x_{t},\theta_{t})\bigr|\leq G\|x_{t}-\tilde{x}_{t}\|.$
+Final bound on $R_{1}$.
+$R_{1}\leq\underbrace{\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta)-\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta_{t})}_{\text{Hedge regret on queried losses}}\;+\;2G\sum_{t=1}^{T}\|x_{t}-\tilde{x}_{t}\|.$
+The first term is exactly the regret of the $K$-experts problem with gains $f_{t-1}^{k}(\tilde{x}_{t})$, and hence with $\eta_{\theta,t}=\sqrt{\log K}/(B\sqrt{t})$,
+$\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta)-\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta_{t})\leq B\sqrt{2T\log K}.$
+Moreover, $\|x_{t}-\tilde{x}_{t}\|\leq\delta$, so the additional term $2G\sum_{t}\|x_{t}-\tilde{x}_{t}\|$ is of order $\sum_{t}\delta$. We upper bound this in the two cases as follows.
+  • One-point feedback. With $\delta=T^{-1/4}$,
+    $2G\sum_{t=1}^{T}\delta=O\!\left(G\sum_{t=1}^{T}T^{-1/4}\right)=O(GT^{3/4}).$
+  • Two-point feedback. With $\delta=T^{-1/2}$,
+    $2G\sum_{t=1}^{T}\delta=O\!\left(G\sum_{t=1}^{T}T^{-1/2}\right)=O(G\sqrt{T}).$
+Note that the term $R_{3}$ depends only on the true loss functions $(f_{t},g_{t})$ and the sequence $\{\theta_{t}\}_{t=1}^{T}$, and not on how gradients are estimated; hence the proof of ${\mathbb{E}}\{R_{3}\}\leq 0$ from Appendix G applies verbatim in the bandit setting as well.
+Thus,
+${\mathbb{E}}\{R_{3}\}\leq 0.$
+We bound $R_{2}$ as follows.
+Bounding $R_{2}$ (one-point feedback).
+Using $\delta=T^{-1/4}$ and $\eta_{x,t}=D/(Gd^{1/2}t^{3/4})$, $\rho=\delta/{\mathsf{r}}$ where $r\mathcal{B}_{1}\subseteq\mathcal{X}\subseteq D\mathcal{B}_{1}$ and $\mathcal{B}_{1}$ is the unit ball, the bandit-OGD analysis [16] yields
+$R_{2}\leq c_{1}\,d^{1/2}GD\,T^{3/4},$
+for some universal constant $c_{1}$.
+Combining the three bounds we get
+${\bar{R}}_{\mathcal{A}}(T)\leq c_{1}\,d^{1/2}GD\,T^{3/4}\;+\;2B\sqrt{2T\log K}.$
+Bounding $R_{2}$ (two-point feedback). Choosing $\delta=T^{-1/2}$, $\rho=\delta/{\mathsf{r}}$ and $\eta_{x,t}=D/(Gd^{1/2}\sqrt{t})$, the two-point bandit-OGD analysis [1, 13, 39] gives
+$R_{2}\leq c_{3}\,d^{1/2}GD\,\sqrt{T},$
+for some universal constant $c_{3}$. Thus, with two-point feedback,
+${\bar{R}}_{\mathcal{A}}(T)\leq c_{3}\,d^{1/2}GD\,\sqrt{T}\;+\;2B\sqrt{2T\log K}.$
+Appendix J Experts Problem (Gains Version)
 We are given $N$ experts. The interaction proceeds for $T$ rounds as follows:
@@
 $\max_{i}\sum_{t=1}^{T}g_{t,i}-\sum_{t=1}^{T}G_{t}\;\;\leq\;\;\sqrt{2T\ln N}.$
-Appendix F Bounding $R_{3}$ with the adversarial input
-In this section, we show the best bound one can obtain for $R_{3}$ on a sample path basis for Algorithm 1 with the adversarial input is $O(\eta_{\theta}T^{2})$.
+Appendix K Proof of Theorem 10
+Given the context is clear in this Appendix, we drop superscript $b$ from both cost $C^{b}_{\mathcal{A}}$ and $C^{b}_{\mathsf{OPT}}$ and write it simply as $C_{\mathcal{A}}$ and $C_{\mathsf{OPT}}$, respectively.
+Proof.
+Recall that at each round $t=1,\dots,T$, the algorithm $\mathcal{A}$ predicts $x_{t}\in\{0,1\}$ and observes bits $(b_{t}^{1},b_{t}^{2})\in\{0,1\}^{2}$. The losses are
+$\ell_{t}^{1}(x_{t})=\mathbf{1}\{x_{t}\neq b_{t}^{1}\},\qquad\ell_{t}^{2}(x_{t})=\mathbf{1}\{x_{t}\neq b_{t}^{2}\}.$
+The algorithm’s total cost on an input sequence $\sigma$ is
+$C_{\mathcal{A}}(\sigma)=\max\Bigl\{\sum_{t=1}^{T}\ell_{t}^{1}(x_{t}),\;\sum_{t=1}^{T}\ell_{t}^{2}(x_{t})\Bigr\}.$
+The static benchmark is
+$C_{\mathsf{OPT}}(\sigma)=\min_{x\in\{0,1\}}\max\Bigl\{\sum_{t=1}^{T}\mathbf{1}\{x\neq b_{t}^{1}\},\;\sum_{t=1}^{T}\mathbf{1}\{x\neq b_{t}^{2}\}\Bigr\}.$
+K.1 Adversarial construction
+Assume $T$ is even. The adversary constructs two inputs $\sigma_{1}$ and $\sigma_{2}$ that are identical for the first $T/3$ rounds.
+Common prefix.
+For $t=1,\dots,T/3$,
+$b_{t}^{1}=0,\qquad b_{t}^{2}=1.$
+Let
+$\bar{x}:=\sum_{t=1}^{T/3}x_{t}$
+denote the number of times the algorithm predicts $1$ during the prefix.
+Suffix selection rule.
+After observing $\bar{x}$, the adversary completes the input as follows:
+  • If $\bar{x}>T/6$, the adversary chooses input $\sigma_{1}$.
+  • If $\bar{x}\leq T/6$, the adversary chooses input $\sigma_{2}$.
+Suffix of $\sigma_{1}$.
+For $t=T/3+1,\dots,T$:
+$b_{t}^{1}\sim\mathrm{Bernoulli}\Bigl(\tfrac{1}{2}\Bigr)\;\text{i.i.d.},\qquad b_{t}^{2}=0.$
+Suffix of $\sigma_{2}$.
+For $t=T/3+1,\dots,T$:
+$b_{t}^{1}=1,\qquad b_{t}^{2}\sim\mathrm{Bernoulli}\Bigl(\tfrac{1}{2}\Bigr)\;\text{i.i.d.}.$
+K.2 Cost of the static benchmark
+We compute $C_{\mathsf{OPT}}(\sigma_{i})$ for $i=1,2$. With $\sigma_{1}$, $\mathsf{OPT}$ predicts $x=0$ always, and
+${\mathbb{E}}\{C_{\mathsf{OPT}}(\sigma_{1})\}={\mathbb{E}}\left\{\max\left\{\sum_{t=1}^{T/3}\mathbf{1}\{0\neq b_{t}^{1}\}+\sum_{t=T/3+1}^{T}\mathbf{1}\{0\neq b_{t}^{1}\},\ \ \sum_{t=1}^{T}\mathbf{1}\{0\neq b_{t}^{2}\}\right\}\right\}\leq T/3+o(T).$
+Similarly, with $\sigma_{2}$, $\mathsf{OPT}$ predicts $x=1$ always, and
+${\mathbb{E}}\{C_{\mathsf{OPT}}(\sigma_{2})\}={\mathbb{E}}\left\{\max\left\{\sum_{t=1}^{T}\mathbf{1}\{1\neq b_{t}^{1}\},\ \ \sum_{t=1}^{T/3}\mathbf{1}\{1\neq b_{t}^{2}\}+\sum_{t=T/3+1}^{T}\mathbf{1}\{1\neq b_{t}^{2}\}\right\}\right\}\leq T/3+o(T).$
+K.3 Lower bound on any online algorithm’s cost
+We now lower bound the expected cost of $\mathcal{A}$.
+Case 1: $\bar{x}>T/6$ (input $\sigma_{1}$).
+For the common prefix,
+$\sum_{t\leq T/3}\ell_{t}^{1}(x_{t})=\bar{x},\qquad\sum_{t\leq T/3}\ell_{t}^{2}(x_{t})=\frac{T}{3}-\bar{x}.$
+In the suffix of $\sigma_{1}$, $b_{t}^{1}$ is unbiased and independent of $x_{t}$, so
+$\mathbb{E}\!\left\{\sum_{t>T/3}\ell_{t}^{1}(x_{t})\right\}=\frac{T}{3}.$
+Thus,
+$\mathbb{E}\!\left\{\sum_{t=1}^{T}\ell_{t}^{1}(x_{t})\right\}=\bar{x}+\frac{T}{3}.$ (26)
+Also,
+$\sum_{t>T/3}\ell_{t}^{2}(x_{t})=\sum_{t>T/3}x_{t}\geq 0.$
+By definition,
+${\mathbb{E}}\{C_{\mathcal{A}}(\sigma_{1})\}\;\geq\;\mathbb{E}\!\left\{\sum_{t=1}^{T}\ell_{t}^{1}(x_{t})\right\}.$
+Thus, using (26) and the condition that $\bar{x}>T/6$, we have
+${\mathbb{E}}\{C_{\mathcal{A}}(\sigma_{1})\}\;\geq\ {\bar{x}}+\frac{T}{3}=\frac{T}{2}.$
+Case 2: $\bar{x}\leq T/6$ (input $\sigma_{2}$).
+By symmetry,
+${\mathbb{E}}\{C_{\mathcal{A}}(\sigma_{2})\}\geq\mathbb{E}\!\left\{\sum_{t=1}^{T}\ell_{t}^{2}(x_{t})\right\}=\Bigl(\frac{T}{3}-\bar{x}\Bigr)+\frac{T}{3}\geq\Bigl(\frac{T}{3}-\frac{T}{6}\Bigr)+\frac{T}{3}=\frac{T}{3}+\frac{T}{6}=\frac{T}{3}+\frac{T}{6}$
+since $\bar{x}\leq T/6$.
+Thus, in both cases
+${\mathbb{E}}\{C_{\mathcal{A}}(\sigma_{i})\}\geq\frac{T}{3}+\Omega(T),$
+which together with ${\mathbb{E}}\{C_{\mathsf{OPT}}(\sigma_{1})\}={\mathbb{E}}\{C_{\mathsf{OPT}}(\sigma_{2})\}\leq\frac{T}{3}+o(T)$ imply that the static min-max regret $R^{b}_{\mathcal{A}}(T)$ for any online algorithm is $\Omega(T)$.
+Appendix L Deriving ${\mathbb{E}}\{R^{b}_{\mathcal{A}}(T)\}=O(\sqrt{T\log T})$ for the binary prediction problem using Algorithm 1
+Let an online algorithm $\mathcal{A}$ choose $x_{t}=1$ with probability $p_{t}\in[0,1]$ and $x_{t}=0$ with probability $1-p_{t}$. Define the expected losses
+$f_{t}(p):=\mathbb{E}[\mathbf{1}\{x_{t}\neq f_{t}\}]=|p-f_{t}|,\qquad g_{t}(p):=\mathbb{E}[\mathbf{1}\{x_{t}\neq g_{t}\}]=|p-g_{t}|.$
+Both $f_{t}(p)$ and $g_{t}(p)$ are convex, 1-Lipschitz functions of $p$, thus, exactly matching the setup of Section 6 with convex losses.
+Thus, if we use Algorithm 1 to find probability $p_{t}$ for choosing $x_{t}=1$, Theorem 5 implies that the min-max static regret with the true benchmark ${\mathbb{E}}\{C_{\mathsf{OPT}}\}$ is ${\mathbb{E}}\{C_{\mathcal{A}}\}-{\mathbb{E}}\{C_{\mathsf{OPT}}\}\leq O(\sqrt{T\log T}).$
+Appendix M Proof of Theorem 11
+We consider the Adaptive Stronger-Bias Double Majority (ASB-DM) algorithm (Algorithm 4) to solve the min-max prediction of binary sequences.
+The ASB-DM algorithm employs a hierarchical voting structure to resolve the online min-max objective.
+At each time step $t$, the algorithm maintains empirical means $\hat{q}_{1,t}$ and $\hat{q}_{2,t}$ for the two cost sequences.
+The decision rule is governed by two distinct layers of majority-based logic:
+  1. 1.
+    The First Majority: Marginal Consensus.
+The algorithm first evaluates the marginal preference of each sequence.
+If both sequences independently reach a consensus—meaning both $\hat{q}_{1,t},\hat{q}_{2,t}<1/2$ (favoring action 0) or both $\hat{q}_{1,t},\hat{q}_{2,t}>1/2$ (favoring action 11)—the algorithm follows this local majority.
+In such cases, the optimal action is unambiguous as both sequences indicate the same superior choice.
+  2. 2.
+    The Second Majority: Stronger-Bias Resolution.
+A fundamental challenge arises during disagreement scenarios, where the sequences provide conflicting marginal majorities.
+To achieve a double majority that resolves this conflict, the algorithm evaluates the sum of empirical means, $\Sigma_{t}=\hat{q}_{1,t}+\hat{q}_{2,t}$.
+This sum identifies which of the two conflicting marginal preferences is statistically “stronger.” The threshold $\Sigma_{t}=1$ acts as the secondary decision boundary: if $\Sigma_{t}<1$, the magnitude of the evidence favors action 0; if $\Sigma_{t}\geq 1$, the bias favors action $1$.
+The efficacy of this approach relies on the concentration of the empirical vector $(\hat{q}_{1,t},\hat{q}_{2,t})$ toward the true parameters $(q_{1},q_{2})$.
+As long as the true parameters do not lie exactly on the boundary $q_{1}+q_{2}=1$, the probability of the algorithm’s double-majority resolution failing decays exponentially.
+This ensures that the decision process stabilizes after a finite number of steps, effectively bounding the total expected mistakes by a constant and yielding the $O(1)$ regret guarantee.
+Algorithm 4 Adaptive Stronger-Bias Double Majority (ASB-DM)
+1:for $t=1,2,\dots,T$ do Compute empirical means $\hat{q}_{1,t}$ and $\hat{q}_{2,t}$.
+2:  if $\hat{q}_{1,t}<1/2$ and $\hat{q}_{2,t}<1/2$ then $x_{t}\leftarrow 0$ $\triangleright$ Case A: Agreement on 0
+3:  else if $\hat{q}_{1,t}\geq 1/2$ and $\hat{q}_{2,t}\geq 1/2$ then $x_{t}\leftarrow 1$ $\triangleright$ Case B: Agreement on 1
+[…]
+5:   if $\Sigma_{t}<1$ then $x_{t}\leftarrow 0$ $\triangleright$ Bias favors 0
+6:   else $x_{t}\leftarrow 1$ $\triangleright$ Bias favors 1
+7:   end if
+8:  end ifPlay $x_{t}$.
+9:end for
+Proof of Theorem 11.
+Let the cumulative loss $L_{k}(x)$ for a fixed static action $x\in\{0,1\}$ on sequence $b_{t}^{k},k=1,2$ be
+$L_{k}(x)=\sum_{t=1}^{T}\ell_{t,k}(x)=\sum_{t=1}^{T}\mathbf{1}\{x\neq b_{t}^{k}\}.$
+Let $L_{k}(x_{1:T})$ denote the cumulative loss incurred by an algorithm $\mathcal{A}$ actions $x_{t}$ on sequence $b_{t}^{k},k=1,2$:
+$L_{k}(x_{1:T})=\sum_{t=1}^{T}\ell_{t,k}(x_{t})=\sum_{t=1}^{T}\mathbf{1}\{x_{t}\neq b_{t}^{k}\}.$
+Let $k^{*}$ be the index of the sequence that maximizes this cumulative loss:
+$k^{*}=\operatorname*{arg\,max}_{k\in\{1,2\}}L_{k}(x_{1:T}).$
+For further exposition, we distinguish between the true path-dependent optimal action and the optimal static action based on expected losses.
+  1. 1.
+    Let $x^{\star}$ be the random variable representing the best action in hindsight for the specific realization of losses:
+    $x^{\star}=\arg\min_{x\in\{0,1\}}\max_{k\in\{1,2\}}L_{k}(x).$
+    The true regret is defined as the difference between the algorithm’s worst-case loss and the benchmark’s worst-case loss:
+    $R_{\mathcal{A}}^{b}(T)=L_{k^{*}}(x_{1:T})-\max_{k\in\{1,2\}}L_{k}(x^{\star}).$
+    Since the maximum loss of the benchmark is necessarily greater than or equal to its loss on the specific sequence $k^{*}$ (i.e., $\max_{k}L_{k}(x^{\star})\geq L_{k^{*}}(x^{\star})$), subtracting the latter yields a valid upper bound on the regret:
+    $R_{\mathcal{A}}^{b}(T)\leq L_{k^{*}}(x_{1:T})-L_{k^{*}}(x^{\star}).$ (27)
+  2. 2.
+    Let $\bar{x}^{\star}$ be the fixed static action that minimizes the worst-case expected cumulative loss:
+    $\bar{x}^{\star}=\arg\min_{x\in\{0,1\}}\max_{k\in\{1,2\}}{\mathbb{E}}\{L_{k}(x)\}=\arg\min_{x\in\{0,1\}}\max_{k\in\{1,2\}}T\mathbb{P}(b_{t}^{k}\neq x).$ (28)
+Regret Decomposition: Using (27), we decompose the regret relative to the fixed reference action $\bar{x}^{\star}$:
+$R^{b}_{\mathcal{A}}(T)$ $\leq L_{k^{*}}(x_{1:T})-L_{k^{*}}(x^{\star})$
+$=\underbrace{\left(L_{k^{*}}(x_{1:T})-L_{k^{*}}(\bar{x}^{\star})\right)}_{\text{(A) Algorithm's Mistakes}}+\underbrace{\left(L_{k^{*}}(\bar{x}^{\star})-L_{k^{*}}(x^{\star})\right)}_{\text{(B) Estimation Cost}}.$ (29)
+Bounding Term (B): We define the expected worst-case loss rate for any fixed action $x$ as:
+$\nu(x)=\max_{k\in\{1,2\}}\mathbb{P}(b_{t}^{k}\neq x).$
+Since we assume the asymmetric case ($q_{1}+q_{2}\neq 1$), the minimizer $\bar{x}^{\star}$ (28) is unique.
+Let $x_{\text{alt}}$ denote the alternative action ($x_{\text{alt}}\neq\bar{x}^{\star}$).
+There exists a strictly positive gap $\delta>0$ defined as:
+$\delta=\nu(x_{\text{alt}})-\nu(\bar{x}^{\star})>0.$
+This gap $\delta$ is strictly positive because $\delta=0$ implies $\nu(0)=\nu(1)$, which can only occur if $q_{k}=1/2$ or $q_{1}+q_{2}=1$.
+Since our theorem assumes $q_{1},q_{2}\neq 1/2$ and $q_{1}+q_{2}\neq 1$, the costs are distinct, ensuring a unique minimizer.
+Note that $\delta$ depends only on the fixed distribution parameters $q_{1},q_{2}$ and is independent of the horizon $T$.
+We now bound Term (B), which represents the estimation cost:
+$\text{(B)}=V(\bar{x}^{\star})-\min_{x\in\{0,1\}}V(x),$
+where $V(x)=\max_{k}L_{k}(x)$ is the realized worst-case cumulative loss.
+Term (B) is positive only if the alternative action becomes empirically optimal (i.e., $V(x_{\text{alt}})<V(\bar{x}^{\star})$, meaning $x^{\star}\neq\bar{x}^{\star}$).
+Since the difference in losses is at most $T$, we bound the expectation as:
+${\mathbb{E}}\{\text{(B)}\}\leq T\cdot\mathbb{P}(V(x_{\text{alt}})<V(\bar{x}^{\star})).$
+The event $\{V(x_{\text{alt}})<V(\bar{x}^{\star})\}$ implies a significant deviation from the expected rates.
+Specifically, if for all $k$ and $x$, the empirical losses satisfy $|L_{k}(x)/T-{\mathbb{E}}\{L_{k}(x)/T\}|<\delta/2$, then:
+$\frac{V(x_{\text{alt}})}{T}=\max_{k}\frac{L_{k}(x_{\text{alt}})}{T}>\nu(x_{\text{alt}})-\frac{\delta}{2}=\nu(\bar{x}^{\star})+\frac{\delta}{2}>\max_{k}\frac{L_{k}(\bar{x}^{\star})}{T}=\frac{V(\bar{x}^{\star})}{T}.$
+This contradicts $V(x_{\text{alt}})<V(\bar{x}^{\star})$.
+Therefore, for the ordering to flip, i.e.
+$\{V(x_{\text{alt}})<V(\bar{x}^{\star})\}$, at least one empirical loss $L_{k}(x)$ must deviate from its mean by at least $\delta/2$, i.e.
+$|L_{k}(x)/T-{\mathbb{E}}\{L_{k}(x)/T\}|\geq\delta/2$.
+Applying the Union Bound over the 4 quantities ($k\in\{1,2\},x\in\{0,1\}$) and Hoeffding’s inequality:
+$\mathbb{P}(V(x_{\text{alt}})<V(\bar{x}^{\star}))\leq\sum_{x,k}\mathbb{P}\left(\left|\frac{L_{k}(x)}{T}-\mathbb{E}\left[\frac{L_{k}(x)}{T}\right]\right|\geq\frac{\delta}{2}\right)\leq 4\cdot 2e^{-2T(\delta/2)^{2}}.$
+Substituting this back into the expectation:
+${\mathbb{E}}\{\text{(B)}\}\leq 8Te^{-\delta^{2}T/2}=o(1).$ (30)
+Bounding Term (A): We now upper bound Term (A), which is defined as $L_{k^{*}}(x_{1:T})-L_{k^{*}}(\bar{x}^{\star})$. Writing in terms of $\ell_{t,k}(x_{t})$, we get
+$\text{(A)}=\sum_{t=1}^{T}\left(\ell_{t,k^{*}}(x_{t})-\ell_{t,k^{*}}(\bar{x}^{\star})\right).$
+Consider the term inside the sum:
+  • If $x_{t}=\bar{x}^{\star}$, then $\ell_{t,k^{*}}(x_{t})-\ell_{t,k^{*}}(\bar{x}^{\star})=0$.
+  • If $x_{t}\neq\bar{x}^{\star}$, the difference is at most $1$ (since losses are binary).
+Thus, regardless of the index $k^{*}$, the term is bounded by the indicator of a mistake:
+$\ell_{t,k^{*}}(x_{t})-\ell_{t,k^{*}}(\bar{x}^{\star})\leq\mathbf{1}\{x_{t}\neq\bar{x}^{\star}\}.$
+Summing over $t$:
+$\text{(A)}\leq\sum_{t=1}^{T}\mathbf{1}\{x_{t}\neq\bar{x}^{\star}\}.$ (31)
+This bounds Term (A) by the total number of mistakes against the fixed reference.
+Next, to bound the expected regret, we analyze the probability of error $\mathbb{P}(x_{t}\neq\bar{x}^{\star})$.
+Note that while $b_{t}^{1}$ and $b_{t}^{2}$ may be correlated at any instant $t$, the vectors $\mathbf{b}_{t}=(b_{t}^{1},b_{t}^{2})$ are i.i.d.
+across time $t$.
+Characterization of Optimal Static Action $\bar{x}^{\star}$.
+We explicitly determine $\bar{x}^{\star}$ by minimizing the worst-case expected cumulative loss derived in the definition.
+For $x=0$, the loss is $T\cdot\mathbb{P}(b_{t}^{k}=1)=Tq_{k}$.
+For $x=1$, the loss is $T\cdot\mathbb{P}(b_{t}^{k}=0)=T(1-q_{k})$.
+The min-max optimal action minimizes the maximum of these expected cumulative losses:
+$\bar{x}^{\star}=\arg\min_{x\in\{0,1\}}\max\left({\mathbb{E}}\{L_{1}(x)\},{\mathbb{E}}\{L_{2}(x)\}\right).$
+  • Case I (Agreement): If $q_{1}<1/2$ and $q_{2}<1/2$, the max cost of $x=0$ is $T\cdot\max(q_{1},q_{2})<0.5T$.
+The max cost of $x=1$ is $T\cdot\max(1-q_{1},1-q_{2})>0.5T$.
+Thus, $\bar{x}^{\star}=0$.
+By symmetry, if both >1/2>1/2, $\bar{x}^{\star}=1$.
+  • Case II (Disagreement): Assume w.l.o.g. $q_{1}<1/2<q_{2}$. The max cost of $x=0$ is determined by sequence 2 (since $q_{2}>q_{1}$):
+    $\max(Tq_{1},Tq_{2})=Tq_{2}.$
+    The max cost of $x=1$ is determined by sequence 1 (since $1-q_{1}>1-q_{2}$):
+    $\max(T(1-q_{1}),T(1-q_{2}))=T(1-q_{1}).$
+    Therefore, $\bar{x}^{\star}=0$ if and only if $Tq_{2}<T(1-q_{1})$, which simplifies to $q_{1}+q_{2}<1$.
+Case I: When either both $q_{1}<1/2$ and $q_{2}<1/2$ or $q_{1}>1/2$ and $q_{2}>1/2$.
+Assume w.l.o.g.
+that $q_{1}<1/2$ and $q_{2}<1/2$; the other case follows similarly.
+The optimal action is $\bar{x}^{\star}=0$.
+The algorithm plays $x_{t}=0$ if $\hat{q}_{1,t}<1/2$ and $\hat{q}_{2,t}<1/2$.
+Conversely, a mistake ($x_{t}=1$) implies at least one empirical mean has deviated:
+$\{x_{t}=1\}\subseteq\{\hat{q}_{1,t}\geq 1/2\}\cup\{\hat{q}_{2,t}\geq 1/2\}.$
+Let $\Delta=\min(1/2-q_{1},1/2-q_{2})>0$. We apply the union bound and Hoeffding’s inequality to the marginal probabilities:
+$\mathbb{P}(x_{t}\neq\bar{x}^{\star})\leq\mathbb{P}(\hat{q}_{1,t}\geq 1/2)+\mathbb{P}(\hat{q}_{2,t}\geq 1/2)\leq 2e^{-2(t-1)\Delta^{2}}.$
+Case II: When either $q_{1}<1/2<q_{2}$ or $q_{2}<1/2<q_{1}$ while $q_{1}+q_{2}<1$.
+Recall the definition of disagreement, denoted by $\mathcal{D}_{t}$, which holds if the empirical means imply conflicting actions:
+$\mathcal{D}_{t}=\left\{\hat{q}_{1,t}<\tfrac{1}{2}\leq\hat{q}_{2,t}\right\}\cup\left\{\hat{q}_{2,t}<\tfrac{1}{2}\leq\hat{q}_{1,t}\right\}.$
+Assume w.l.o.g. $q_{1}<1/2<q_{2}$ and $q_{1}+q_{2}<1$, for which $\bar{x}^{\star}=0$. A mistake ($x_{t}=1$) occurs if:
+  1. 1.
+    False Agreement on 1: $\hat{q}_{1,t}\geq 1/2$ and $\hat{q}_{2,t}\geq 1/2$. This implies $\hat{q}_{1,t}\geq 1/2$.
+  2. 2.
+    False Resolution: Disagreement holds, but $\Sigma_{t}=\hat{q}_{1,t}+\hat{q}_{2,t}\geq 1$.
+(Recall that per Algorithm 4, the algorithm selects $x_{t}=1$ in a disagreement scenario only if $\Sigma_{t}\geq 1$).
+This implies $\Sigma_{t}\geq 1$.
+Thus, the mistake event is a subset of $\{\hat{q}_{1,t}\geq 1/2\}\cup\{\Sigma_{t}\geq 1\}$. Let $\Delta_{1}=1/2-q_{1}>0$ and $\Delta_{\Sigma}=1-(q_{1}+q_{2})>0$.
+For the first term, we apply Hoeffding’s inequality to the marginal $\hat{q}_{1,t}$.
+For the second term, we define the sum variable $Z_{s}=b_{s}^{1}+b_{s}^{2}$.
+Since the pairs $(b_{s}^{1},b_{s}^{2})$ are i.i.d.
+across time, the variables $Z_{s}$ are i.i.d.
+across time with range $[0,2]$.
+The correlation between $b_{s}^{1}$ and $b_{s}^{2}$ affects the variance but not the independence or the range of $Z_{s}$.
+Applying Hoeffding’s inequality to the average $\Sigma_{t}=\frac{1}{t-1}\sum_{s=1}^{t-1}Z_{s}$:
+$\mathbb{P}(\Sigma_{t}\geq 1)=\mathbb{P}(\Sigma_{t}-(q_{1}+q_{2})\geq\Delta_{\Sigma})\leq e^{-\frac{2(t-1)\Delta_{\Sigma}^{2}}{(2-0)^{2}}}=e^{-(t-1)\Delta_{\Sigma}^{2}/2}.$
+Combining these, we bound the probability of error:
+$\mathbb{P}(x_{t}\neq\bar{x}^{\star})\leq e^{-2(t-1)\Delta_{1}^{2}}+e^{-(t-1)\Delta_{\Sigma}^{2}/2}.$
+For all cases, $\mathbb{P}(x_{t}\neq\bar{x}^{\star})\leq\kappa e^{-c(t-1)}$ for constants $\kappa,c>0$. Using (31) and the bound (30) for Term (B), from (29) we get
+${\mathbb{E}}\{R^{b}_{\mathcal{A}}(T)\}\leq\sum_{t=1}^{T}\mathbb{P}(x_{t}\neq\bar{x}^{\star})+8Te^{-\delta^{2}T/2}\leq\sum_{t=1}^{\infty}\left(2e^{-c_{1}t}+e^{-c_{2}t}\right)+8Te^{-\delta^{2}T/2}=O(1).$
+Appendix N Martingale Difference Inputs
+In this section, we extend our main result (Theorem 5 of Section 6) to the setting where the loss sequence satisfies a Martingale difference condition.
+Let $\{\mathcal{F}_{t}\}_{t\geq 0}$ denote the natural filtration generated by the online algorithm and the loss sequence. We assume that for every $x\in\mathcal{X}$,
+$\mathbb{E}\!\left\{\lambda_{t}(x)\mid\mathcal{F}_{t-1}\right\}=\mu(x),\qquad\lambda_{t}(x):=(f_{t}(x),g_{t}(x)),$ (32)
+where $\mu(x)$ is a deterministic vector. We further assume that $f_{t}$ and $g_{t}$ satisfy Assumptions 1 and 2 (boundedness, convexity, and Lipschitz continuity).
+Recall the benchmark
+${\mathbb{E}}\{C_{\mathsf{OPT}}\}={\mathbb{E}}\left\{\min_{x\in\mathcal{X}}\max\!\left\{\sum_{t=1}^{T}f_{t}(x),\;\sum_{t=1}^{T}g_{t}(x)\right\}\right\},$ (33)
+and the deterministic surrogate benchmark
+$\bar{C}_{\mathsf{OPT}}=T\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x),$ (34)
+and the expected min-max static regret is defined as before
+${\bar{R}}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}\{{\mathbb{E}}\{C_{\mathcal{A}}\}-{\bar{C}}_{\mathsf{OPT}}\}.$
+Lemma 14 (Benchmark Gap under Martingale Difference Inputs).
+Under the Martingale difference model (32) and Assumptions 1 and 2,
+${\mathbb{E}}\left\{\bigl|C_{\mathsf{OPT}}-\bar{C}_{\mathsf{OPT}}\bigr|\right\}\;\leq\;O(\sqrt{T\log T}).$
+Proof.
+The proof follows Appendix D verbatim with one modification.
+For a fixed $x\in\mathcal{X}$, the sequence $\{\lambda_{t}(x)-\mu(x)\}_{t=1}^{T}$ is a bounded Martingale difference sequence.
+Hence, the concentration step in Appendix D, which uses Hoeffding’s inequality for i.i.d.
+variables, is replaced by Freedman’s inequality as follows.
+Fix $x\in\mathcal{X}$ and define
+$Z_{t}(x):=\lambda_{t}(x)-\mu(x),\qquad\mathbb{E}\{Z_{t}(x)\mid\mathcal{F}_{t-1}\}=0.$
+For each coordinate $i\in\{1,2\}$, $|Z_{t,i}(x)|\leq 2B$ almost surely, and define
+$V_{T,i}(x):=\sum_{t=1}^{T}\mathbb{E}\!\left\{Z_{t,i}(x)^{2}\mid\mathcal{F}_{t-1}\right\}\leq 4B^{2}T.$
+We apply Freedman’s inequality [17], which states that for any scalar Martingale difference sequence $\{Z_{t}\}$ with $|Z_{t}|\leq b$ almost surely,
+$\Pr\!\left(\sum_{t=1}^{T}Z_{t}\geq u\right)\leq\exp\!\left(-\frac{u^{2}}{2(V_{T}+bu/3)}\right).$
+This replaces the use of Hoeffding’s inequality in Appendix D.
+We apply Freedman’s inequality coordinate-wise to $\lambda_{t}(x)$, then similar to Appendix D, take a union bound over the two coordinates and an $\varepsilon$–net of $\mathcal{X}$, and then use Lipschitz continuity to extend the bound to all $x\in\mathcal{X}$.
+This yields
+$\mathbb{E}\!\left\{\sup_{x\in\mathcal{X}}\Big\|\sum_{t=1}^{T}\lambda_{t}(x)-T\mu(x)\Big\|_{\infty}\right\}=O(\sqrt{T\log T}),$
+which implies the stated bound exactly as in Appendix D. ∎
 Theorem 15.
-Under Assumption 1, i.e., $\|\lambda_{t}(x)\|_{\infty}\leq B$, $R_{3}\leq 2B^{2}T\sum_{t=1}^{T}\eta_{\theta,t}$.
-Recall the definitions $\lambda_{t}(x)=(f_{t}(x),g_{t}(x))$,
-$R_{3}\;=\;\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\langle\theta_{t},\lambda_{t}(x)\rangle-C_{\mathsf{OPT}},\qquad\text{and}\qquad C_{\mathsf{OPT}}=\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\sum_{t=1}^{T}\langle\theta,\lambda_{t}(x)\rangle.$
-Thus, an almost natural way (or really the only way for Algorithm 1 that combines OGD+Hedge) to bound $R_{3}$ is via exploiting the stability $\sum_{t}\|\theta_{t+1}-\theta_{t}\|_{1}$ of the Hedge algorithm.
-Following this path, we next derive Lemma 11, Lemma 12 and Lemma 13 which when combined imply Theorem 10.
-However, it is not sufficient since by substituting $\eta_{\theta,t}=\frac{1}{\sqrt{T}}$ for all $t$, the usual Hedge step size, we get $R_{3}=\Omega(T^{3/2})$.
-Lemma 11.
-Define the time-averaged vector $\bar{\theta}:=\frac{1}{T}\sum_{t=1}^{T}\theta_{t}\in\Delta_{2}.$ Then,
-$R_{3}\leq B\sum_{t=1}^{T}\|\theta_{t}-\bar{\theta}\|_{1}.$
+Under Assumptions 1 and 2, for $\eta_{\theta,t}\leq\sqrt{\frac{2\ln 2}{B^{2}t}}$ and $\eta_{x,t}=\frac{D}{G\sqrt{t}}$, the expected min–max static regret (8) of Algorithm 1 under the Martingale difference model for $(f_{t},g_{t})$ satisfies
+$\bar{R}_{\mathcal{A}}(T)=O(\sqrt{T}).$
 Proof.
-Since $\bar{\theta}$ is a particular (constant) element of $\Delta_{2}$, for every $x\in\mathcal{X}$
-$\sum_{t=1}^{T}\langle\bar{\theta},\lambda_{t}(x)\rangle\leq\max_{\theta\in\Delta_{2}}\sum_{t=1}^{T}\langle\theta,\lambda_{t}(x)\rangle,$
-and therefore taking the minimum over $x$ yields
-$\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\langle\bar{\theta},\lambda_{t}(x)\rangle\leq C_{\mathsf{OPT}}.$
-Thus,
-$R_{3}$ $=\min_{x}\sum_{t=1}^{T}\langle\theta_{t},\lambda_{t}(x)\rangle-C_{\mathsf{OPT}}$
-$\leq\min_{x}\sum_{t=1}^{T}\langle\theta_{t},\lambda_{t}(x)\rangle-\min_{x}\sum_{t=1}^{T}\langle\bar{\theta},\lambda_{t}(x)\rangle$
-$\leq\max_{x}\sum_{t=1}^{T}\big\langle\theta_{t}-\bar{\theta},\lambda_{t}(x)\big\rangle,$
-where the last inequality follows as below.
-Let $x^{\star}$ satisfy $u(x^{\star})=\min_{x}u(x)$. Then
-$\min_{x}u(x)-\min_{x}v(x)=u(x^{\star})-\min_{x}v(x)\leq u(x^{\star})-v(x^{\star})\leq\max_{x}(u(x)-v(x)),$
-which proves the claim.
-This proves
-$R_{3}\leq\max_{x\in\mathcal{X}}\sum_{t=1}^{T}\langle\theta_{t}-\bar{\theta},\lambda_{t}(x)\rangle.$ (29)
-Since we are in the bounded setting, for all rounds $t$ and all $x\in\mathcal{X}$ we have $\|\lambda_{t}(x)\|_{\infty}\leq B$.
-Thus, appyling the Hölder’s inequality (using duality of $||.||_{1}$ and $||.||_{\infty}$ norms) on (29),
-$\sum_{t=1}^{T}\langle\theta_{t}-\bar{\theta},\lambda_{t}(x)\rangle\leq\sum_{t=1}^{T}\|\theta_{t}-\bar{\theta}\|_{1}\,\|\lambda_{t}(x)\|_{\infty}\leq B\sum_{t=1}^{T}\|\theta_{t}-\bar{\theta}\|_{1}.$
-Thus,
-$R_{3}\leq B\sum_{t=1}^{T}\|\theta_{t}-\bar{\theta}\|_{1}.$
-This is the desired inequality. ∎
-Next, we state an elementary inequality without proof.
-Lemma 12 (Deviation from average via consecutive differences).
-Let $\theta_{1},\dots,\theta_{T}\in\mathbb{R}^{d}$ be vectors, and define the average
-$\bar{\theta}=\frac{1}{T}\sum_{t=1}^{T}\theta_{t}.$
-Then
-$\sum_{t=1}^{T}\|\theta_{t}-\bar{\theta}\|_{1}\;\leq\;2T\sum_{t=1}^{T-1}\|\theta_{t+1}-\theta_{t}\|_{1}.$
-Lemma 13 (One-step TV bound for Hedge).
-Let $\{\theta_{t}\}$ be the Hedge distributions updated by
-$\theta_{t+1,i}\;=\;\frac{\theta_{t,i}\,e^{\eta_{\theta,t}\lambda_{t,i}}}{\sum_{j}\theta_{t,j}\,e^{\eta_{\theta,t}\lambda_{t,j}}},\qquad i=1,\dots,n,$
-with learning rate $\eta_{\theta,t}>0$. Assume the losses satisfy $\lambda_{t,i}\in[0,B]$ for all $i$. Then
-$\|\theta_{t+1}-\theta_{t}\|_{1}\leq\eta_{\theta,t}B.$
+The regret decomposition and the bounds on ${\mathbb{E}}\{R_{1}\}$ and ${\mathbb{E}}\{R_{2}\}$ are identical to those in the proof of Theorem 5. It remains to show that ${\mathbb{E}}\{R_{3}\}\leq 0$.
+Recall
+$R_{3}=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Lambda_{t}(x,\theta_{t})-\bar{C}_{\mathsf{OPT}},\qquad\Lambda_{t}(x,\theta)=\theta^{\top}\lambda_{t}(x).$
+Let $(x^{\star},\theta^{\star})$ solve
+$\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x).$
+As in the proof of Lemma 8, we have the deterministic inequality
+$R_{3}\leq\sum_{t=1}^{T}(\theta_{t}-\theta^{\star})^{\top}\lambda_{t}(x^{\star}).$ (35)
+Taking conditional expectation with respect to $\mathcal{F}_{t-1}$ and using (32),
+$\mathbb{E}\!\left\{(\theta_{t}-\theta^{\star})^{\top}\lambda_{t}(x^{\star})\mid\mathcal{F}_{t-1}\right\}=(\theta_{t}-\theta^{\star})^{\top}\mu(x^{\star}).$
+By optimality of $\theta^{\star}$, the right-hand side is non-positive almost surely. Summing over $t$ and taking expectations in (35) yields
+${\mathbb{E}}\{R_{3}\}\leq 0.$
+Combining Lemma 14 and Theorem 15, the true min-max static regret of Algorithm 1 with respect to the benchmark ${\mathbb{E}}\{C_{\mathsf{OPT}}\}$ is $O(\sqrt{T\log T})$ under the Martingale difference condition.
+Algorithm 5 Blocked Hedge+OGD (for Markovian Inputs)
+1:Input: Block width $H$, step sizes $\eta_{x,m}>0$, $\eta_{\theta,m}>0$, feasible set $\mathcal{X}$
+2:Initialize: $w_{1}=(1,1)$, $\theta_{1}=w_{1}/\|w_{1}\|_{1}$, $x_{1}\in\mathcal{X}$ arbitrary
+3:Let $M=\lceil T/H\rceil$ be the number of blocks.
+4:for $m=1,\dots,M$ do
+5:  1. Freeze: Set current action $x=x_{m}$ and weights $\theta=\theta_{m}$
+6:  Set effective block length $H_{m}=\min(H,\,T-(m-1)H)$
+7:  2. Initialize block accumulators:
+8:  $\bar{g}_{m}=0$, $\bar{\lambda}_{m}=0$
+9:  3. Play and Observe (Block Loop):
+10:  for $t=(m-1)H+1,\dots,(m-1)H+H_{m}$ do
+11:   Play action $x_{t}=x$
+12:   Observe losses $(f_{t}(x),g_{t}(x))$ and gradients $(\nabla f_{t}(x),\nabla g_{t}(x))$
+13:   Evaluate loss vector $\lambda_{t}=(f_{t}(x),g_{t}(x))$
+14:   Form surrogate gradient:
+$\nabla\tilde{\Lambda}_{t}(x,\theta):=\theta_{1}\nabla f_{t}(x)+\theta_{2}\nabla g_{t}(x)$
+15:   Accumulate:
+16:    $\bar{\lambda}_{m}\leftarrow\bar{\lambda}_{m}+\lambda_{t}$
+17:    $\bar{g}_{m}\leftarrow\bar{g}_{m}+\nabla\tilde{\Lambda}_{t}(x,\theta)$
+18:  end for
+19:  4. Block averaging:
+20:  $\bar{\lambda}_{m}\leftarrow\bar{\lambda}_{m}/H_{m}$
+21:  $\bar{g}_{m}\leftarrow\bar{g}_{m}/H_{m}$
+22:  5. Updates (End of Block):
+23:  OGD: $y_{m+1}=x_{m}-\eta_{x,m}\bar{g}_{m}$
+24:    $x_{m+1}=\mathrm{Proj}_{\mathcal{X}}(y_{m+1})$
+25:  Hedge: $w_{m+1,i}=w_{m,i}\exp(\eta_{\theta,m}\bar{\lambda}_{m,i})$, $i=1,2$
+26:  Normalize: $\theta_{m+1}=w_{m+1}/\|w_{m+1}\|_{1}$
+27:end for
+Appendix O Markovian Inputs
+In this section, we extend our main result (Theorem 5 of Section 6) to the setting where the loss sequence evolves according to a stationary Markov process.
+Unlike the Martingale difference case, temporal dependence introduces a persistent bias that requires modifying the algorithm to decouple the learning updates from the instantaneous state of the Markov chain.
+Let $\{S_{t}\}_{t\geq 1}$ be a time-homogeneous, ergodic Markov chain on a measurable state space $\mathcal{S}$ with transition kernel $P$ and stationary distribution $\pi$.
+We assume the chain is uniformly ergodic, i.e., there exist constants $C>0$ and $\tau_{\mathrm{mix}}>0$ such that for any initial distribution $\nu$,
+$\|\nu P^{t}-\pi\|_{\mathrm{TV}}\leq Ce^{-t/\tau_{\mathrm{mix}}},\qquad\forall t\geq 1.$
+Let $\lambda:\mathcal{X}\times\mathcal{S}\to\mathbb{R}^{2}$ be a measurable mapping and define
+$\lambda_{t}(x):=\lambda(x,S_{t})=(f_{t}(x),g_{t}(x)).$
+We assume that $f_{t}$ and $g_{t}$ satisfy Assumptions 1 and 2.
+Define the stationary mean
+$\mu(x):=\mathbb{E}_{\pi}[\lambda(x,S)],\qquad\forall x\in\mathcal{X},$ (36)
+which is independent of $t$ by stationarity.
+Recall the benchmark
+${\mathbb{E}}\{C_{\mathsf{OPT}}\}={\mathbb{E}}\!\left[\min_{x\in\mathcal{X}}\max\!\left\{\sum_{t=1}^{T}f_{t}(x),\;\sum_{t=1}^{T}g_{t}(x)\right\}\right],$ (37)
+and the deterministic surrogate benchmark
+$\bar{C}_{\mathsf{OPT}}=T\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x),$ (38)
+and the expected min-max static regret is defined as before
+$\bar{R}_{\mathcal{A}}(T)=\sup_{P}\{{\mathbb{E}}\{C_{\mathcal{A}}\}-\bar{C}_{\mathsf{OPT}}\}.$
+Lemma 16 (Benchmark Gap under Markovian Inputs).
+Under the Markovian model described above and Assumptions 1 and 2,
+${\mathbb{E}}\left\{\bigl|C_{\mathsf{OPT}}-\bar{C}_{\mathsf{OPT}}\bigr|\right\}\;\leq\;O(\sqrt{T\log T})+O(\tau_{\mathrm{mix}}).$
 Proof.
-Let
-$Z_{t}:=\sum_{j}\theta_{t,j}e^{\eta_{\theta,t}\lambda_{t,j}}\;=\;\mathbb{E}_{i\sim\theta_{t}}\!\big[e^{\eta_{\theta,t}\lambda_{t,i}}\big].$
-By the definition of Kullback–Leibler divergence,
-$D_{\mathrm{KL}}(\theta_{t}\|\theta_{t+1})$ $=\sum_{i}\theta_{t,i}\ln\frac{\theta_{t,i}}{\theta_{t+1,i}}=\sum_{i}\theta_{t,i}\ln\frac{\theta_{t,i}}{\theta_{t,i}e^{\eta_{\theta,t}\lambda_{t,i}}/Z_{t}}$
-$=\sum_{i}\theta_{t,i}(\ln Z_{t}-\eta_{\theta,t}\lambda_{t,i})=\ln Z_{t}-\eta_{\theta,t}\,\mathbb{E}_{i\sim\theta_{t}}[\lambda_{t,i}].$
-Let $\mu:=\mathbb{E}_{i\sim\theta_{t}}[\lambda_{t,i}]$.
-Then $\ln Z_{t}=\ln\mathbb{E}_{\theta_{t}}\big[e^{\eta_{\theta,t}\lambda_{t,i}}\big]=\eta_{\theta,t}\mu+\ln\mathbb{E}_{\theta_{t}}\big[e^{\eta_{\theta,t}(\lambda_{t,i}-\mu)}\big]$, so (1) is equivalently
-$D_{\mathrm{KL}}(\theta_{t}\|\theta_{t+1})=\ln\mathbb{E}_{\theta_{t}}\big[e^{\eta_{\theta,t}(\lambda_{t,i}-\mu)}\big].$ (30)
-Thus to upper bound the one-step KL it suffices to bound the centered MGF $\ln\mathbb{E}[e^{\eta_{\theta,t}(\lambda_{t,i}-\mu)}]$.
-Proposition 14 (Hoeffding’s MGF bound).
-For a random variable $X$ that satisfies $X\in[a,b]$ almost surely and $\mathbb{E}[X]=\mu$, then for any $s\in\mathbb{R}$
-$\ln\mathbb{E}[e^{s(X-\mu)}]\leq\frac{s^{2}(b-a)^{2}}{8}.$ (31)
-Since $\lambda_{t,i}\in[0,B]$, using (31) in (30) with $X=\lambda_{t,i}$, [a,b]=[B,B][a,b]=[-B,B], $s=\eta_{\theta,t}$, and mean $\mu$, we get
-$D_{\mathrm{KL}}(\theta_{t}\|\theta_{t+1})\leq\frac{\eta_{\theta,t}^{2}4B^{2}}{8}.$ (32)
-Proposition 15 (Pinsker’s inequality [12]).
-For any two probability vectors $p,q$,
-$\|p-q\|_{1}\leq\sqrt{2\,D_{\mathrm{KL}}(p\|q)}.$ (33)
-Since both $\theta_{t}$ and $\theta_{t+1}$ are distributions, applying (33) on (32), we get
-$\|\theta_{t+1}-\theta_{t}\|_{1}\leq\sqrt{2\cdot\frac{\eta_{\theta,t}^{2}B^{2}}{2}}=\eta_{\theta,t}B,$
-the claimed bound. ∎
+The proof follows Appendix D with the sole modification that concentration of $\sum_{t=1}^{T}\lambda_{t}(x)$ around $T\mu(x)$ uses Markov-chain concentration instead of independence.
+Specifically, using uniform ergodicity and standard additive functional concentration (e.g., [36]), for any fixed $x\in\mathcal{X}$:
+$\mathbb{E}\!\left\{\Big\|\sum_{t=1}^{T}\lambda_{t}(x)-T\mu(x)\Big\|_{\infty}\right\}=O(\sqrt{T\log T})+O(\tau_{\mathrm{mix}}).$
+All minimax and $\varepsilon$-net arguments of Appendix D remain unchanged, yielding the stated bound. ∎
+Standard application of Algorithm 1 fails in the Markovian setting because the algorithm’s updates $\theta_{t}$ become highly correlated with the instantaneous state $S_{t}$, leading to an $O(T)$ regret.
+To address this, we propose a Blocked Hedge+OGD strategy in Algorithm 5.
+The time horizon $T$ is divided into $M=\lceil T/H\rceil$ blocks of length $H$.
+Within block $m$, the algorithm freezes its action $x_{m}$ and weights $\theta_{m}$, accumulates losses, and performs a single update for both the Hedge and the OGD at the end of the block.
+Theorem 17.
+Under Assumptions 1 and 2, running the Blocked Hedge+OGD algorithm (Algorithm 5) with block size $H=\Theta(T^{1/3}\tau_{\mathrm{mix}}^{2/3})$ yields an expected min-max static regret of
+$\bar{R}_{\mathcal{A}}(T)=O(T^{2/3}\tau_{\mathrm{mix}}^{1/3}).$
+Proof.
+Let $M=T/H$ be the number of blocks.
+We decompose the regret over the blocks similar to (10).
+The expected min-max static regret ${\bar{R}}_{\mathcal{A}}(T)$ is decomposed into three terms by adding and subtracting $\sum_{m=1}^{M}\Lambda^{(m)}(x_{m},\theta_{m})$ and $\min_{x\in\mathcal{X}}\sum_{m=1}^{M}\Lambda^{(m)}(x,\theta_{m})$:
+${\bar{R}}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}\{\mathbb{E}\{R_{1}\}+\mathbb{E}\{R_{2}\}+\mathbb{E}\{R_{3}\}\}$
+where:
+$R_{1}=\max_{\theta\in\Delta_{K}}\sum_{m=1}^{M}\theta^{\top}\Lambda^{(m)}(x_{m})-\sum_{m=1}^{M}\theta_{m}^{\top}\Lambda^{(m)}(x_{m}),$
+$R_{2}=\sum_{m=1}^{M}\theta_{m}^{\top}\Lambda^{(m)}(x_{m})-\min_{x\in\mathcal{X}}\sum_{m=1}^{M}\theta_{m}^{\top}\Lambda^{(m)}(x),$
+and
+$R_{3}=\min_{x\in\mathcal{X}}\sum_{m=1}^{M}\theta_{m}^{\top}\Lambda^{(m)}(x)-{\bar{C}}_{\mathsf{OPT}}.$
+In Appendix O.1, we prove the following bounds
+$\mathbb{E}\{R_{1}+R_{2}\}\leq O(H\sqrt{M})=O(H\sqrt{T/H})=O(\sqrt{TH}).$ (39)
+$\mathbb{E}\{R_{3}\}\leq\sum_{m=1}^{M}O(\tau_{\mathrm{mix}})=\frac{T}{H}O(\tau_{\mathrm{mix}}).$ (40)
+Combining these, the expected min-max static regret is bounded by:
+$\bar{R}_{\mathcal{A}}(T)\leq O(\sqrt{TH})+O\left(\frac{T\tau_{\mathrm{mix}}}{H}\right).$
+Minimizing this bound with respect to $H$ requires balancing $\sqrt{TH}$ with $T\tau_{\mathrm{mix}}/H$, which yields the optimal block length $H=\Theta(T\tau_{\mathrm{mix}}^{2})^{1/3}$.
+Substituting this back gives the final bound $O(T^{2/3}\tau_{\mathrm{mix}}^{1/3})$.
+∎
+Combining Lemma 16 and Theorem 17, the true min-max static regret of the Blocked Hedge+OGD Algorithm 5 with respect to the true benchmark ${\mathbb{E}}\{C_{\mathsf{OPT}}\}$ is
+$O(T^{2/3}\tau_{\mathrm{mix}}).$
+O.1 Proofs of (39) and (40)
+We provide the complete derivation for the upper bounds on the expected regret terms $\mathbb{E}\{R_{1}\}$, $\mathbb{E}\{R_{2}\}$, and $\mathbb{E}\{R_{3}\}$ for the Blocked Hedge+OGD algorithm with block size $H$.
+Let $M=\lceil T/H\rceil$ be the total number of blocks.
+We denote the blocks by $\mathcal{B}_{1},\dots,\mathcal{B}_{M}$.
+O.1.1 Upper Bound for $\mathbb{E}\{R_{1}\}$
+The term $R_{1}$ captures the regret of the weight updates $\theta_{m}$ against the best fixed distribution $\theta^{\star}$ over the sequence of blocks:
+$R_{1}=\sum_{m=1}^{M}(\theta_{m}-\theta^{\star})^{\top}\Lambda^{(m)}(x_{m}),$
+where $\Lambda^{(m)}(x)=\sum_{t\in\mathcal{B}_{m}}\lambda_{t}(x)$ is the cumulative loss vector for block $m$.
+Proof.
+  1. 1.
+    The loss vector for each step $m$ is the block sum $\Lambda^{(m)}(x_{m})$. Given that $\|\lambda_{t}\|_{\infty}\leq B$ (Assumption 2), the magnitude of the loss vector is bounded by:
+    $\|\Lambda^{(m)}(x_{m})\|_{\infty}\leq\sum_{t\in\mathcal{B}_{m}}\|\lambda_{t}\|_{\infty}\leq H\cdot B\text{.}$
+  2. 2.
+    Although Algorithm 5 updates weights using the block-average loss $\bar{\lambda}_{m}=\Lambda^{(m)}/H_{m}$, this is mathematically equivalent to the gains version of Hedge running on the meta-sequence of blocks with an effective step size $\tilde{\eta}_{\theta,m}=\eta_{\theta,m}/H$.
+  3. 3.
+    For a horizon of $M$ steps with losses bounded by $L_{\max}=HB$ and $K=2$ experts, the Hedge algorithm yields:
+    $R_{1}\leq HB\sqrt{2M\ln 2}\text{.}$
+  4. 4.
+    Using $M\approx T/H$
+    $R_{1}\leq HB\sqrt{\frac{2T\ln 2}{H}}=B\sqrt{2TH\ln 2}\text{.}$
+Thus, $\mathbb{E}\{R_{1}\}=O(\sqrt{TH})$. ∎
+O.1.2 Upper Bound for $\mathbb{E}\{R_{2}\}$
+The term $R_{2}$ captures the regret of the action updates $x_{m}$ against the best fixed action $x^{\star}$:
+$R_{2}=\sum_{m=1}^{M}\left(\sum_{t\in\mathcal{B}_{m}}\theta_{m}^{\top}\lambda_{t}(x_{m})-\sum_{t\in\mathcal{B}_{m}}\theta_{m}^{\top}\lambda_{t}(x^{\star})\right)\text{.}$
+Proof.
+  1. 1.
+    The effective gradient for block $m$ is the sum of gradients: $G^{(m)}=\sum_{t\in\mathcal{B}_{m}}\nabla(\theta_{m}^{\top}\lambda_{t}(x_{m}))$.
+Given $\|\nabla\lambda_{t}\|\leq G$ (Assumption 3), the gradient norm is bounded by:
+    $\|G^{(m)}\|\leq\sum_{t\in\mathcal{B}_{m}}\|\nabla\lambda_{t}\|\leq H\cdot G\text{.}$
+  2. 2.
+    Algorithm 5 performs OGD using the block-average gradient $\bar{g}_{m}=G^{(m)}/H_{m}$.
+In the regret analysis, this is equivalent to OGD using the cumulative gradient $G^{(m)}$ with an effective step size $\tilde{\eta}_{x,m}=\eta_{x,m}/H$.
+  3. 3.
+    For $M$ steps with gradient norms bounded by $\mathcal{G}_{\max}=HG$ and domain diameter $D$, OGD yields:
+    $R_{2}\leq D(HG)\sqrt{M}=DHG\sqrt{\frac{T}{H}}=DG\sqrt{TH}\text{.}$
+Thus, $\mathbb{E}\{R_{2}\}=O(\sqrt{TH})$. ∎
+O.1.3 Upper Bound for $\mathbb{E}\{R_{3}\}$
+$\begin{aligned} R_{3}&=\min_{x\in\mathcal{X}}\sum_{m=1}^{M}\theta_{m}^{\top}\Lambda^{(m)}(x)-{\bar{C}}_{\mathsf{OPT}},\\ &\stackrel{{\scriptstyle(a)}}{{\leq}}\sum_{m=1}^{M}\theta_{m}^{\top}\Lambda^{(m)}(x^{\star})-\sum_{m=1}^{M}H(\theta^{\star})^{\top}\mu(x^{\star}),\\ &\stackrel{{\scriptstyle(b)}}{{=}}\sum_{m=1}^{M}\theta_{m}^{\top}(\Lambda^{(m)}(x^{\star})-H\mu(x^{\star}))+\sum_{m=1}^{M}(\theta_{m}-\theta^{\star})^{\top}(H\mu(x^{\star})),\\ &\stackrel{{\scriptstyle(c)}}{{=}}\sum_{m=1}^{M}(\theta_{m}-\theta^{\star})^{\top}(\Lambda^{(m)}(x^{\star})-H\mu(x^{\star}))+\sum_{m=1}^{M}(\theta_{m}-\theta^{\star})^{\top}(H\mu(x^{\star}))\\ &\quad+\sum_{m=1}^{M}(\theta^{\star})^{\top}(\Lambda^{(m)}(x^{\star})-H\mu(x^{\star})),\\ &\stackrel{{\scriptstyle(d)}}{{\leq}}\sum_{m=1}^{M}\|\theta_{m}-\theta^{\star}\|_{1}\|\Lambda^{(m)}(x^{\star})-H\mu(x^{\star})\|_{\infty}+\sum_{m=1}^{M}(\theta_{m}-\theta^{\star})^{\top}(H\mu(x^{\star}))\\ &\quad+\sum_{m=1}^{M}\|\theta^{\star}\|_{1}\|\Lambda^{(m)}(x^{\star})-H\mu(x^{\star})\|_{\infty},\\ &\stackrel{{\scriptstyle(e)}}{{\leq}}3\sum_{m=1}^{M}\|\Lambda^{(m)}(x^{\star})-H\mu(x^{\star})\|_{\infty},\end{aligned},$
+where $(a)$ follows from the suboptimality of the stationary optimal action $x^{\star}\in\mathcal{X}$, $(b)$ is obtained by adding and subtracting the mixed term $\sum_{m=1}^{M}\theta_{m}^{\top}(H\mu(x^{\star}))$, $(c)$ uses the algebraic identity $\theta_{m}^{\top}V=(\theta_{m}-\theta^{\star})^{\top}V+(\theta^{\star})^{\top}V$, where $V=\Lambda^{(m)}(x^{\star})-H\mu(x^{\star})$, to center the algorithm’s weights around the stationary optimal weights $\theta^{\star}$, $(d)$ follows from Holder’s inequality, where $u^{\top}v\leq\|u\|_{1}\|v\|_{\infty}$, and finally $(e)$ utilizes three key properties:
+  1. 1.
+    since $\theta_{m},\theta^{\star}\in\Delta_{K}$, we have $\|\theta_{m}-\theta^{\star}\|_{1}\leq\|\theta_{m}\|_{1}+\|\theta^{\star}\|_{1}=2$,
+  2. 2.
+    stationary optimality of $\theta^{\star}$ ensures $(\theta_{m}-\theta^{\star})^{\top}\mu(x^{\star})\leq 0$ for all $\theta_{m}\in\Delta_{K}$, so $\sum_{m=1}^{M}(\theta_{m}-\theta^{\star})^{\top}(H\mu(x^{\star}))$ is bounded by zero, and
+  3. 3.
+    $\|\theta^{\star}\|_{1}=1$ because it is a probability distribution on the simplex.
+To complete the proof, we now bound the expected deviation term $\|\Lambda^{(m)}(x^{\star})-H\mu(x^{\star})\|_{\infty}$ using the mixing properties of the chain.
+By the law of iterated expectations, we have $\mathbb{E}[\|\Lambda^{(m)}(x^{\star})-H\mu(x^{\star})\|_{\infty}]=\mathbb{E}[\mathbb{E}[\|\sum_{t\in\mathcal{B}_{m}}(\lambda_{t}(x^{\star})-\mu(x^{\star}))\|_{\infty}\mid\mathcal{F}_{\text{start}(m)}]]$, where $\mathcal{F}_{\text{start}(m)}$ is the filtration at the start of block $m$.
+  1. 1.
+    For a uniformly ergodic Markov chain, the expected deviation of the state from stationarity decays exponentially with time.
+Let $S_{\text{start}(m)}$ be the state at the start of block $m$.
+For the $k$-th step within the block ($1\leq k\leq H$):
+    $\left\|\mathbb{E}[\lambda_{(m-1)H+k}(x^{\star})\mid S_{\text{start}(m)}]-\mu(x^{\star})\right\|_{\infty}\leq Ce^{-k/\tau_{\mathrm{mix}}}$
+  2. 2.
+    We sum this deviation over the $H$ steps in the block:
+    $\left\|\mathbb{E}\left[\sum_{t\in\mathcal{B}_{m}}(\lambda_{t}(x^{\star})-\mu(x^{\star}))\;\Bigg|\;\mathcal{F}_{\text{start}(m)}\right]\right\|_{\infty}\leq\sum_{k=1}^{H}Ce^{-k/\tau_{\mathrm{mix}}}.$
+    This geometric series is bounded by a constant proportional to the mixing time
+    $\sum_{k=1}^{\infty}Ce^{-k/\tau_{\mathrm{mix}}}=O(\tau_{\mathrm{mix}})$
+    Thus, the total bias accumulated per block is $O(\tau_{\mathrm{mix}})$.
+  3. 3.
+    Summing this constant error over the $M$ blocks
+    $\mathbb{E}\{R_{3}\}\leq 3\sum_{m=1}^{M}O(\tau_{\mathrm{mix}})=M\cdot O(\tau_{\mathrm{mix}})=\frac{T}{H}O(\tau_{\mathrm{mix}})$
+Thus, we get
+$\mathbb{E}\{R_{3}\}=O\left(\frac{T\tau_{\mathrm{mix}}}{H}\right).$